A Complete Bibliography of The American Statistician: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.03

Title word cross-reference

2 × 2 [Dav86, Kra88]. 2 × k [Cox82]. [−1, 1] [Koc85]. _{2} [Sea88a]. X̄ [Sti84]. c [Hod84a, Hod84b]. χ^{2} [SP85a]. ED_{100p} [CCW^{+}86]. F [FK86]. f′/f [JH83]. m [PKG86]. μ [Ray85]. N [Cam88, PKG86, Kah87]. O [KD87]. QR [Ans85]. r [Hod84a, Hod84b]. R^{2} [WS88]. R^{2} > r^{2}yx_{1} + r^{2}yx_{2} [Ham87b]. S^{2} [Dav85a, Sti84, Duf86]. Σ [Ray85]. t [Ber81, Box81, GK84, Lac84]. U [Len83, Haw89]. X [Spi86, Ano87i, BA87, Spi87]. X^{2} [Cox82]. Y [Spi86, Ano87i, Spi87]. Z [Haw89, Shi88b].

-by- [Hod84a, Hod84b]. -statistics [KD87]. -Way [Cam88].

/MS [LAH89].

02 [KF89].

1 [Sug86]. 1.1 [Ber87b]. 1/2 [Man83]. 1/2-Tail [Man83]. 11 [DK88].

16-Year-Olds [Pos81]. 1910-1981 [Wei82b]. 1912-1980 [Smi81]. 1926-1980 [MMM82]. 1979 [DS81, GH80]. 1980 [BMM83, CS86, Dav81, DM81]. 1981 [Bec82, Kis82, PS82, Rya82]. 1983 [Ano84h, Ord84]. 1985 [AS86, Ham87a]. 1986 [FK87]. 1988 [TG89].

2 [Lev87a]. 2-Tail [Man83]. 2SLS [Ste88a].

3.77 [NN86]. 350 [Sau84b]. 38 [SYC88, Zel85]. 39 [CH86, Kar86, Kar87].

40 [Spi87]. 43 [Ang90, GY90]. 4CaST [Lev87a]. 4CaST/2 [Lev87a].

5-T [Nea82, NP82, Par82]. 5-T/Dioxin [Nea82].

62004a [Zel85].

77 [Chu88]. 78 [Wai80].

85b [SYC88]. 87b [Kar86, Kar87]. 87c [Zel85].

Ability [AG88]. absolute [Bly80]. Absorption [MGS85]. Academic [MO84]. accelerated [MS87]. accelerated-life-testing [MS87]. Accessible [RD88]. ACED [Wel85]. ACFPLOT [Pol80]. Acid [FBC^{+}83, Rob83b]. ACO [Sea88a]. Activities [LH81]. Activity [Pos82]. Ad [Sha86b, Bar83a, Boa83, Bro83b, Fre83a, Geh83, Gre83b, IHD^{+}83, Kem83, Kis83, Mar83d, Mos83, Rem83, Ric83, Rob83a, Sol83]. Adaptive [Joh81]. Add [Dal88a]. Add-Ons [Dal88a]. Addelman [DS81]. Addendum [SP85b]. Adequacy [Jen87]. adjustment [Pie80]. Administrative [Joh82b]. Advanced [Hin88]. Adverse [CG82]. Affairs [WBH^{+}88]. after [Lou81]. Age [Mar83b, Mar83c, Yel83a]. Agencies [Bis84, Bro84, PHA84]. Agency [Joh82b]. Agents [CCW^{+}86]. Agglomerative [Gil86]. Aggregates [Bar84]. Aging [Sto88]. Aid [PT80]. Aided [Hah89]. Air [BH83, Dow85, WL85]. Alfonso [FK87]. Algebraic [MGS85]. Algorithm [Ans85, Far81, Lin80, Pau85, Pos89]. Algorithmic [Cob84, Whe85]. Algorithms [CGL83, Gle88, Wel85]. ALLOC.P [Mer88]. Allocation [Gon80, Mer88, Mer89, WC87]. Allocations [Ban88]. Almanac [BDMW87]. Alpha [Smi84]. ALSCAL [YTL80]. alternate [Far89]. Alternative [Cox82, Ham83a, KM85, KM88, Nel85, SSM80, Bro83a, GL83, Ham83b, Joh83, Naj83]. Alternatives [Man83]. Always [Ham87b]. Amer [Ang90, CH86, GY90, SYC88, Spi87, Zel85]. Amer. [Kar86, Kar87]. American [BFM^{+}88, Boa85, God85, Hah85c, Joi85a, Joi85b, LH81, Wac85]. Among [Lee86]. AMOS [Arb89]. Analyses [BK80, Ela84, Gam81, GNS82, Gil84, MBM84a, MBM84b, VPO^{+}85, SSH81]. Analysis [ABS81, Alb89, Ami87, AC82, Ano85i, Ano85g, Ano86e, Ano89g, Arb89, AG84, Bak84, BR84, BDMW87, BW84, Cam88, Car81a, Car81b, Cas85, CGL83, CC89, Coo83, Dal87a, Daw89, Del86, DL82, DK82, Don82, Edm85, Ell88, FS87, FC81, Fer88, FC87, FR86, Flu89, FHI^{+}88, FTL84, Fro81, GG80a, GH88, Gil86, Gre86a, Gre86b, GD81, Har86b, HN86, Hoe85, IMS87, Joh81, KG80a, KGHS84, Lam80, Lev87a, LSWB89, Lew80, LNC^{+}81, Lit88, LM86, LS89, Lun80, Lun84, MP80, MEP^{+}86, MRP^{+}89, Mas89, MLB80, McC80, MD81, MM84, NZ88, Nor87, Per84, PH86, Pin88, PC83, Rai82, Rei84, Rei87, Rel85, Rob87, Roh87, RM80, RM81, Rom87, Sta86, Sau84a, Sav82, Saw84, Sch88a, SHF83, Sha84, Sta88, ST83, Ste82, Ste87, Sti80, Str80, Stu82]. Analysis [TH81, Tia83, Van80, Wai80, Wal87, War85, Wat82, Wat81, Web87, Wei82a, WS88, Wil82, Ber87c, Man82, Rav85, Sch85b, Geb89]. Analysis-of-Variance [Cam88]. Analytic [Ano87g, DW86]. Analyze [Ant80]. Analyzing [Hod84a, Whi85]. Animal [MGS85]. Annotated [Sea81, SHF83, Sea88a]. Announcement [Ken86]. ANOVA [Ano88g, Arm81, Cob84, Fir82, Fra81b, Hei81a, Hei81b, Her86, NW81, Pay81, Rea85, Rog84, SD87, Sto81]. ANOVR [Gam81]. Answers [VR83]. APCLUST [Gil86]. APL [Hun89, Lew80, Pol80, Pol82, Sav82]. APLECON [Lew80]. APLECON-Econometric [Lew80]. Apple [Dus86, LMM84]. Applicable [Man81, TW86]. Application [HC86, KC84, Low83, Saw84, Sch84a]. Applications [AL86, BH86, Lev80, Lun88, Moo81, RC88]. Applied [Bas81, KS88]. Approach [Ano89f, Cha82, Cob84, Dun89, Hu88, LV89, MBM^{+}85, RBL86, Rup87, Sha82, Stu82, Wat81, GP89, MS87]. approaches [Gol86]. Appropriate [KH89]. Appropriateness [DCB88]. Approximate [Dri80]. Approximating [Sch88b]. Approximation [Ano87g, DW86, Nor89, SS89, Sha85, She84]. Arbitrarily [SZM87]. Arbitrary [Man81]. Areas [Ban88, Nor89, Sha85]. Arnold [Gei84b]. Art [PS83]. Arts [Ive85, MR89]. ASA [Ano80e, Ano85h, Bis84, BHH^{+}80, Box80a, Bra82, Bro84, DL86, EWW^{+}82, Gup82, Hig82, Joh82a, Kna82, Leo82, MFG^{+}84, Min80, Mos82, Owe82, Par80, PHA84, Sne82, Swi82, Tan82a, Tan82b, Wat80, Zah82]. ASCII [Dal87b, KF89]. Aspects [Bol85, Lov83, Mul80, NP82, SC86]. Assays [SD89]. Assess [Gup82, Hig82, Joh82a, Kna82, Leo82, Owe82, Swi82, Tan82a, Tan82b]. Assessing [BS85]. Assessment [CA85, Die83b, Gun84, Joh82b, Kra88, Par82, PS83, Str82b, Swa82, Wei82a, MS87]. Assist [Gup82, Hig82, Joh82a, Kna82, Leo82, Owe82, Swi82, Tan82a, Tan82b]. Assisted [LNC^{+}81]. Associated [Chu88]. Association [Cox82, Rav86, BFM^{+}88, LH81]. Associations [Ern83, Kra88, Lov83]. Assumption [BA87, KMM87]. Assumptions [Ano88g, Gam81, Mag82, SD87]. Assurance [DP85]. Asymmetric [HPB82]. asymmetry [CW88]. Asymptotic [CCW^{+}86, Haw89, KB84a, Thu89, Har80]. Atari [KC84]. Audio [SP84, SP85b]. Audio-Visual [SP84, SP85b]. author [Efr86b, Gei84b, Zel88b]. authors [PS89a]. AUTOBJ [Rei84, Rei87].

AUTOBOX [Rei84, Rei87]. Autocovariance [MJ84]. Automatic [Sto80]. Automation [Del86]. Autoregressive [Lie80, ME84]. Available [JSL89]. Averages [Tor86, AT86, SP86]. Averaging [Ano87e, Kru87]. Avoid [Bis84, Bro84, PHA84]. Awareness [RM88].

B [KF89]. B-02 [KF89]. Back [Ano80a, Ano80b, Ano80c, Ano80d, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d]. Background [Atk82]. Bad [Loy83]. Badly [Wai84]. Balanced [Cam88, Gle88, Gun84]. Balkrishna [Ful80]. Ball [HC86]. Balloon [Cha84]. Balls [Pen88]. Ban [ZR82]. Bancroft [FK87]. Bands [Ano84f, Hau83]. Bank [Ano88h, Bal87]. BAR [Ano87p]. Baseball [BF83]. Based [Bel87, KMM87, AO89, Aso80]. Baseline [Sei89]. Basic [Koc85, Pos83]. BAYCAT [AG84]. Bayes [AG84, Ber88, BM88b, Cas85, Hil88, Jay88, Kul88, Pau81, RR84, Zel88b, Zel88a]. Bayesian [Che86, CH86, Lin86b, Mor86, Pre86, Smi86, ABS81, Alb89, Ano86f, CH85, Edw86, Efr86b, Efr86a, FHI^{+}88, Kah87, Nov84, Zel89]. Be [Ame85, BA87, Ehr84, Sei89, Shv84, TVV85, BvE88, MA84]. Bear [DS89]. Becoming [BZ89]. Behavior [AL86]. Behavioral [Lun88]. Benzene [Lav82, Swa82, Tho82]. Bernard [AS86]. Bernardo [Gei84b]. Bernoulli [NW86]. Best [HC86, Kem89b, PS89b, Sea89b, BvE88, Kot86, MA84, PS89a, SS84]. Best-Ball [HC86]. Better [GH82, KL83]. Between [Ame85, BH83, CI81a, CI81b, Ern83, Fli81, Lov83, MLB80, Noe81, Rob83b, Sne84, Aso80, FK86, VM89, Wil82]. Between-Group [Ame85]. Bias [Dav85a, Duf86, Mil84]. Biased [Fox89]. Bibliography [MSGG87]. Bill [Col81]. Binary [Ano85f, Ber84, Gei84a, HTM89, Zel84, Gei84b, Zel85]. Binomial [Ano85e, DCB88, HA86, Lit89, Lou81, Loy83, MPH88, Pau85, SS89, Shu88, AS84, Bly80, Kah87, She84]. biographical [Bol84]. Biological [Lov83]. BIOM [Roh84]. Biostatistics [Bor87, Car87, DSAGR87]. Birch [Cox84]. Birthday [Sch88b]. Bivariate [Low83, Tru81, GM86]. BKWPAC [Wil81]. Block [Mur88, Mur86]. Blocking [LAH89]. Blocks [LAH89]. Blood [BH83]. BMD [MBD84]. BMDP [Fra81a, MBD84]. Bonferroni [Sch84c]. Book [Bar86, Bel80, Ber89, Bes87, Boa82a, Bur87, BG80, CJ81, CT89, Chr81, Con89, Coo87, Dav85b, Dav89, Dix88, Ell87, Gil80, Gol88, Hal89, HB86, Hil85, KM89a, KP87, KG80b, KSB^{+}81, Koe87, Koe86, Lac85, LeP87, Lev87b, Lig85, Lig88, Lin85b, Lin87, Lin88, Man88, Mar85, McB80, Mil89, Mor87a, Mor87b, Nas86a, Nas86b, O’B88, Par89, Pla86, Pos80, Pos82, Pos83, PF89, Reg86, Sch85a, Shr87, Shu86, Ste88b, Ste86, Tan89, Tur87a, Vin89, Wei88, WM85, Woo87]. Books [Pos82]. Boolean [Hel88]. Bootstrap [Gle88, Lun88, Yan88, EG83].

Borel [Tom84]. Both [Ano87i, Tuk80, Spi86, Spi87]. Bounds [FP81]. Box [BG87, Ben88, Pen88, Lie80, Rei80, RC88]. Boxplot [Ben88, FHI89]. BOXX [Rei84, Rei87]. Breaks [Cle84a]. Breeding [MGS85]. Bridge [Chu88, CI81a, CI81b, Fli81, Noe81]. Brier [Ano85i, BL85]. Broadening [Mos88]. Brownian [Chu88]. Buffered [WMB82]. Bureau [Bai84]. Business [Ano80e, Mar83a].

CADA [LNC^{+}81]. Calculate [Hun89, MPH88]. Calculating [Cox82, Dal89d, PG83, PT80]. Calculation [Ena88, FS89, Man81]. Calculator [Dus89, Nor89, Tur87b, dJ84a]. Calculators [Boa82a]. Calculus [Nor84]. Calibration [Ano85i, BL85]. Campus [BS89]. Can [BA87, Bra83a, Bra83b, Car83, Hil83, Hun83b, Leo83, McD83, Ott83]. Canal [SC86]. Cancer [Ern83, Lov83]. Canonical [HPB82, Mul82, Wil82]. Carcinogenesis [KGHS84, PS83]. Care [Coh82]. Careers [BHH^{+}80, Box80a, Bra82, EWW^{+}82, Min80, Mos82, Par80, Sne82, Wat80, Zah82]. Carl [Bec82]. Carlo [Gle86a]. CASCNTRL [GD81]. Case [Ano88h, Bal87, BMM83, Fre80, GH82, GD81, JW82, RD88]. Case-Control [GD81]. Cassette [GW80, Koe81]. Categorical [AG84, Saw84]. Category [Gas89, Kai89b, Kai89a, Sch89b]. Causal [LM86]. Caused [CM83]. Caution [Sea89a]. Cautionary [WS88, Kah87]. CDM [DP85]. Cell [Fre80, Hod84a, JW82, Mur88, Sch84a]. Cells [Web87, Mur86]. CENSOR [MD81]. Censorship [Man81]. Census [Ano85h, MFG^{+}84, Bai84, BMM83, CS86]. Center [Boe82, Hac87]. Centering [Bel84a, Bel84b, Coo84, Woo84]. Central [Gus86, PG82, Tar81, Yan88]. Certain [KHK89, WB81, MS87]. Change [MO84]. Changes [BMP^{+}85, TVV85]. Characteristic [Col83]. Characteristics [Gon80]. Characterization [DL86, Nai87, GY89, GY90]. Characterizing [BCV88]. Charges [DS89]. Chart [Bol85, Sch86]. Charts [Cle84a, Dun87, Dun88b]. Chebyshev [SYC88, Ano88e, SYM84]. Checking [GH84]. chemistry [RC88]. Chernoff [TT80]. Chernoff-Type [TT80]. Cherry [Fai86]. Chi [Coh88, Gil81, Han83, HW86, Hjo88, PT80, Pos89, WB81]. Chi-Square [Coh88]. Chi-Squared [Gil81, Han83, HW86, PT80, Pos89, WB81]. Chinese [Cho86, Hsi81]. Chlorofluorocarbon [Wei82a]. Chlorofluorocarbons [Atk82]. Choosing [Ano89g, Sta88]. Choropleth [GH82]. Chronic [BS83, Die83b, KGHS84]. Class [AG88, DM87, Mee86]. Classic [Fle82]. Classification [CS86, KG80a, Poo80, Wil82]. Classifier [Joh81]. Classroom [FHI^{+}88, Mos80, SI82]. Clinical [DP85, Sei89, Ber87c, RC88]. cLOGISTIC [Dal89a]. Closer [Has87]. Clumped [Pau85]. CLUSTAR [RM80, Rom87]. CLUSTAR-PC [Rom87]. Cluster [Edm85, RM80, Rom87, SR83, Edm85]. Clustering [Gil86, KL83]. CLUSTID [RM80]. Cochran [Col81, DM81]. Code [Smi84]. Coefficient [Cha84, Kin89, Koc85, LG85, RT81, RN88, ST83, WT80, GP89]. Coefficients [Gus86, Han89, KMM87]. Cohort [FM89, MP80, MEP^{+}86, Whi85]. Collaborators [Ano80g, Ano82f, Ano84i, Ano86i, Ano87j, Ano88i, Ano89i]. Collapsing [Sha82]. Collected [Sei89]. Collection [Coh82, Dal89b, Hof84, ONC85]. College [Car87]. Colleges [GF80, MR89]. Collinearity [Gun84, SM84, Woo84]. Color [Bea83, CM83, Dun89, WF80]. Color-Caused [CM83]. Coloring [Tru81]. Column [BZ89, Gat89, Kem89a, Sea89a]. Combination [CCW^{+}86]. combinations [Ham84]. Combined [Har86c]. Combining [Han89]. Come [Hun81]. Command [Fir82]. Comment [Arm81, Bai84, Bar83a, Ber84, Ber88, Boa83, Boa85, Bon81, Bon88, Box80a, Bra82, Bro83a, Bro83b, Car83, Che86, Coo84, Dat82, DeM83, Dun81, Dun88a, Ela84, Fie83, Fli81, Fra81b, Fre83a, GL83, Gas89, Geh83, Gen85, Gil84, God85, Gre83b, Gun84, Gup82, Hah85c, Hig82, Hil83, Hil88, Hun83b, Jay88, Joh82a, Joh83, Kas82, Kem83, Kem89b, Kis83, Kna82, Kul88, Leo82, Leo83, Lin86b, Mar81b, Mar83d, McD83, Min80, Mor86, Mos81, Mos82, Mos83, Naj83, Noe81, NW81, Nov84, Ott83, Owe82, Par80, Pay81, Pre86, Pri87, Rem83, Rey88, Ric83, Rob83a, Ros87, Sch89b, Sea89b, Sla83, Smi86, Sne82, SM84, Sol83, Swi82, Tae81, Tuk85c, VW82, Vel85, Wac85, Wal82, Wat80, Woo84, Zah82, Zel82b, Zel84]. Comments [Ros81b, Tuk85a, Gei84b, PS89a, Zel88b]. Committee [Ano80e, Bar83a, BHH^{+}80, Boa83, Box80a, Bra82, Bro83b, EWW^{+}82, Fre83a, Geh83, Gre83b, IHD^{+}83, Kem83, Kis83, Mar83d, Min80, Mos82, Mos83, Par80, Rem83, Ric83, Rob83a, Sne82, Sol83, Wat80, Zah82, MGS82]. Common [Bry84, DCB88, Lee86]. Commonly [Ano84g, Don82]. Community [VPO^{+}85]. Compact [Per84]. Comparative [Ano80f, BK80]. Comparing [BB89]. Comparison [DGS^{+}89, HA86, Pav88, REZ88, SC81, Aso80, BvE88, SI82]. Comparisons [FTL84, NZ88, Sto81, DK87, HT83]. Compatible [Ell88]. Complete [Del86, Hin82, LR83, SR83]. Complete-Data [LR83]. Complex [BMM83, CBJ86]. Components [Ano85i, BL85, Mul82, Sea81, Stu82]. Composition [Sto80]. Comprehensive [Geb81, MD85]. COMPSTAT [Geb81]. Computation [FR85, Pec80]. Computational [EJKS87, MP88, SSH81]. Computations [SBA85]. Compute [Har86c, Sav82]. Computer [BR84, BS85, BW84, CC89, Dem86, FS87, Fox89, Gal84, GG80a, GNS82, Ham81, JM85, Jer88, Lam80, Lev87c, LNC^{+}81, Lie80, MD81, NN82b, NN82a, Rei80, Roh84, RM80, RM81, Saw84, SD86, SF80, Sea81, SHF83, Sea88a, Sha86a, SD89, VW82, Web87, WM87, Zar80, PK82]. Computer-Assisted [LNC^{+}81]. Computer-Generated [JM85]. Computerized [PT80]. Computers [Boa82b, DK88, Ell88, Gol89, LMM84]. Computing [Ano89e, Ber87b, Boa82b, Cha80, Chu88, Con87, Lam83, Lin86a, Mul80, Sea89a, SCR88, VW81, Wai80, CGL83]. Concepts [BAP89, KM89b]. Conceptual [Rue85]. Concerning [Mas89, WF80]. Concomitant [DM87]. Concomitant-Variable [DM87]. Conditional [Dal89a, Kor84, Sch89a]. Conditioning [Bel84a, Bel84b, Coo84]. conditions [MA84]. Cones [Fai86].

CONF [Shu88]. Confidence [Ano84f, Ano85e, Car86, CCW^{+}86, FBC^{+}83, Hau83, HA86, Jen87, Lou81, MD83, Sch81a, Ang89, Ang90, AS84, Ber81, Buo87]. Confidence-Interval [Jen87]. Confidentiality [Ano80e]. Confirmatory [Tuk80]. connection [VM89]. Consequence [KK84]. Consistent [DKY80]. Constant [Woo84]. Constrained [SM83]. Constructing [RR85]. Construction [Lin85a, Wel85, Hid86]. Consultants [Mar81a, MBM^{+}85]. Consultation [Bro80a]. Consulting [Bas81, Boe82, CSM86, ED84, GF80]. Contaminated [Riv81]. Content [Bis84, Bro84, Kem80, PHA84]. Contingency [Cox82, Dal87a, Dav86, Gil81, Hod84a, Hod84b, Pai85, Sha82]. continuous [Mul88]. Contrasting [MBM84a, Ela84, Gil84, MBM84b]. Contrasts [Har86c, Sch85b]. Contributions [Col81]. Control [GD81, RR85, Sal85b, VR83]. Controversies [Ste82]. Controversy [FBC^{+}83]. Conversational [NW80]. Convolutions [Dri80]. Cooperation [Sne84]. Coordination [Bon83a, Bon83b, DeM83, Fie83, Sla83, Wal83]. copulas [GM86]. Cornfield [GH80]. Correction [Ang90, Ano85e, Ano86e, Ano87e, Ano89g, Ano89h, Ano89e, Ano89f, CH86, GY90, Kar86, Kar87, SYC88, Spi87, Zel85]. Corrections [Ano85f, Ano85h, Ano85i, Ano85g, Ano86f, Ano86g, Ano87g, Ano87f, Ano87h, Ano87i, Ano88e, Ano88g, Ano88f, Ano88h]. Correctly [LAH89]. CORREL [DK82]. Correlated [Ham87b]. Correlation [Ano85g, Bro86, Cha84, Han89, Har86a, KMM87, Kin89, Koc85, LSWB89, LG85, Mul82, Pai85, Pol80, RN88, SZM87, ST83, WT80, FK86, Rav85, SP86, Sti86]. Correlations [Kor84]. Correspondence [Gre86a, MLB80]. Corridenda [Ano84e]. Corrigenda [Ano84f, Ano84g]. Corrigendum [Ano82e, Ano84h]. Cost [Ell88, Lev87c]. Cotton [BS83, Die83b]. Course [Die89, Fie84, Koe81, Mar83a, Nel81, Sal85b, VD84, WMB82, Wat81, Wat83, WDS84]. Courses [BS89, BK86, Noe80]. Court [Tho82]. Covariance [Ame85, Sch88a, SHF83, ST83]. Covariances [Tam84]. Covariate [SD86]. Cox [Lie80, Lin86a, MM80, RC88]. Craig [DG86, RD88]. CRAWTRAN [Ano86h]. CRAWTRAN-PC [Ano86h]. Create [Hin80, Sto80]. Criteria [Mar81a, BvE88]. Criterion [KM85]. Critical [BM88a, MCJ87, Har80]. Cross [Die83a, Pol80, Wai80, EG83]. Cross-Correlation [Pol80]. Cross-Sectional [Die83a]. Cross-Tabulation [Wai80]. cross-validation [EG83]. CRUNCHER [Hin82]. Curiosities [Kem89a]. Current [Die83a, Ste82, Sto82, Gol86]. Curriculum [Hof84]. Curve [Ano84f, DS89, Hau83, Mil84, Nor89, Sha85, SK85, Woo80]. Curve-Fitting [Woo80]. Curves [Ami87, DGS^{+}89, IMS87, NZ88, REZ88]. Cyclone [Neu85]. Cyclones [Pik85].

Daily [Est85, KS83]. Damage [Dow85, Lod85, Rue85, WL85, MS87]. damage-assessment [MS87]. Darlington [Moo86]. dasy [Rel85]. Data [Ami87, Ano85g, AG84, BDMW87, BB89, Cas85, Cle84a, Coh82, CBJ86, Die83a, DK82, Dri86, DP85, Fox83, FTL84, FP87, Geb89, Gol82, Gon80, Gre86b, Has81, Hin80, HTM89, Hod87, HE84, KF89, Lev87a, LNC^{+}81, LV89, Lit82, LR83, Lod85, Lov84, Man81, MD81, ONC85, Pav88, Per84, PH86, Pik85, Rei84, Rel85, Rob87, Ros81a, RR84, Ros89, Saw84, SF80, Sei89, Sha84, Shu88, SM83, SM84, Ste87, Sti80, Str80, SA87, Swa84, Tia83, Van80, Wai84, Wat82, Web87, Whi85, KD87, Rav85, SSH81]. Data-Dependent [RR84]. Data-Fitting [LV89]. Davis [Dal89d]. Dealing [Ano84g, Don82, Sto88]. Death [Kem89a]. DEC [Sau84b]. Decision [GW80, Str82b, Swa82]. Decisions [BR84, Yel83b]. Decomposition [Nel85, EW85, Man82]. defined [BG89]. Definite [Ame85]. Definiteness [MJ84]. Degree [DJ84b, SZM87]. Degrees [CC89, Hod84a]. delta [BG89, Cox84]. Demeaning [Bel84a, Bel84b, Coo84]. Democracy [Mar83b, Mar83c, Yel83a]. Demography [Coh86, Fer88]. Demonstrate [Jer88]. Demonstration [Koc85, Wil82]. Dennis [Ber87b]. denominator [FK86]. Densities [GSW82]. Density [LW82]. Departments [CSM86]. Depend [SM84]. Dependence [Dav85a, FK86]. Dependent [Ano86h, Gre83a, Pav88, RR84, Sei89, Duf86]. depending [GD89]. Depreciation [DS89]. Derivation [Jac84, KK84, WS87]. Derive [Haw89]. Derived [Lin81]. Deriving [Ehr83a, GSW82, LW82, dCS86]. DESCRIP [Gus86]. DESCRIP-STAT [Gus86]. description [Sch85b]. Descriptions [Ste85]. Design [BPT82, Bol85, Koe81, Mur88, SM83, SD89, Whe85, Box80b, Hid86]. Designed [CC89]. Designing [Wat81]. Designs [Arm81, BK80, Car81b, Fra81b, Gam81, Hei81b, NW81, Pay81, Pin88, SC81, Sto81, Wel85, GS86, Hei81a, Her86, Mur86, WC87]. Desktop [Boa82b]. Detection [Kra88]. Determination [Pen81, RT81]. Determinations [Dal86a]. Determine [Fu84]. Determining [AG88, Ban88, Smi84]. Development [DG86, RC88]. Developments [Hal88, Pie80]. Deviance [Has87]. Diagnostic [HG83, MC87, Sti86]. Diagnostics [Ano87i, Bel84a, Bel84b, Coo84, Gun84, SM84, VW81, Wil81, Spi86, Spi87]. Diagonal [Col83]. Diagram [TH81]. DIALOG [Rah81]. Diet [Juk82]. Difference [HA86, Lin81, MBD84, Pen81, RJ82, Wei83]. Differences [FP81, YTL80]. Differencing [Che88]. Differentiation [Saw84]. Digidot [Hun88]. Dilution [SD89, Web87]. Dimensional [Ano87f, DHPR88, SW86]. Dioxin [HHF82, Nea82]. Direct [Lit82]. Directions [Gen85, Hah85a, Hah85b, Mar84, Tuk85c, Vel85, Tuk85a]. Directories [Ano80e]. Directory [Ste85]. Disaster [Bon83a, Bon83b, DeM83, Fie83, Sla83]. Discrete [Ano87p, FM84, Has81, Lin81, RJ82, SA87, BL87, EG87, GD89, Mee86]. Discriminant [FR85, Flu89]. Discrimination [McC80]. Discussion [Tuk85b, Efr86b]. Disease [BS83, Die83b]. DISFIT [Has81]. Disgrace [Bon83a, Bon83b, DeM83, Fie83, Sla83]. Dispersed [ED84]. Display [ASS80, Pol80, Ros81a, Wai84]. Displays [KL83]. DIST [Ano87p, Edg84, MPH88]. Distance [FR86]. distinct [Aso80]. Distribution [Ano86f, Ano87g, BB84, Bur86, DW86, DGS^{+}89, FM84, Fu84, GS86, Han83, Han85, Haw89, Hef88, HC86, HA89, Kin89, Kon87, MN88, MT82, Nor84, Pos89, Ray85, SS89, Smi85, Sti84, Box81, CH85, CH86, Dav80, Gle89, She84, WS87, dCS86]. Distribution-Free [DGS^{+}89, Kon87, GS86]. Distributional [FS87, Mag82]. Distributions [Ano85f, Ano87p, Ber84, Edg84, Gei84a, Has81, Ima82, Lee86, LS88, Lin81, Mac86, RJ82, Zel84, BL87, Gei84b, GM86, GD89, Gle89, Mee86, Mul88, Zel85]. Divisia [Bar84]. Dixon [Che82]. Do [Bra83a, Bra83b, Bro80b, Car83, Cov89, Fu84, Hil83, Hun83b, Jen87, Leo83, McD83, Mul80, Ott83, Sch89a]. Domain [BCV88, SM84]. Done [Ame85]. Dose [Ami87, IMS87, MS87]. Dose-Response [Ami87, IMS87, MS87]. Dot [Cle84a]. Double [AJ88, Nor84]. Dunnett [Lev80]. Durbin [FS87]. during [CG82]. Dust [Die83b]. Dynamic [Dun89, Bol86]. dynamics [VY85].

Early [Sch81b, Noe84]. Easy [Ena88, Sch84b]. EBA [Fow88]. Eccentric [Hjo88]. Ecological [Fer88, VPO^{+}85]. Econometric [Bak84, Gre85, Lew80, Pec80, WM87]. Econometrics [Bar85, BA87]. Economic [AL86]. Economics [Dow85]. ECPSES [MCJ87]. EDA [GSW82, Rah81]. Edition [Sea88a]. Editor [Gen86, Ken86, Ano84e, Ano84j, Ano88f, ADS81, BSB^{+}86, BGF^{+}81, BBC^{+}89, BZ89, BDR^{+}88, CPE^{+}84, CHP^{+}89, ENH^{+}80, EFJ^{+}89, FSW^{+}87, FSJ^{+}89, FSM^{+}80, Gat89, GBL^{+}87, GFL^{+}85, GN85, GLM^{+}81, GMB^{+}84, HLT^{+}87, HRP^{+}86, JEE^{+}83, JKL^{+}81, KKN^{+}80, Kem89a, KHH^{+}86, KLC^{+}88, LBW^{+}88, MHP^{+}87, MKF^{+}86, NMF^{+}80, NBA^{+}82, PSD^{+}83, RSI^{+}84, SCC^{+}84, Sea89a, SB83, SBE^{+}82, WSA^{+}85, vdLFS^{+}88]. Editorial [Ano80g, Ano82f, Ano84i, Ano86i, Ano87j, Ano88i, Ano89i, Bis84, Bro84, PHA84]. Educating [MBM^{+}85]. Education [BHH^{+}80, Bor82a, Bor82b, Box80a, Bra82, Dat82, EJKS87, EWW^{+}82, Gup82, Hig82, Hog85, Ive85, Joh82a, Kna82, Leo82, Min80, Mos82, Mos88, Owe82, Par80, Sne82, SP84, SP85b, Swi82, Tan82a, Tan82b, Wal82, Wat80, Zah82]. Educational [Mar84]. Effect [Coh82, Roy86, Rue85, Woo84]. Effective [Ber87a, Pos89, SI82]. Effectiveness [Ano84g, Don82]. Effects [AG88, Ano88g, Bro83a, CG82, GL83, Ham83a, Ham83b, Joh83, Mur88, Naj83, SP86, GS86, SD87, VY85]. Efficiencies [CBJ86]. Efficiency [BN88, LAH89]. Efficient [BB84, Leh80, VW81, JH83]. Egon [Dav81]. Election [Har88]. Elementary [Cob84, Fie84, Hof84, Pos82, Cox84, GY89, GY90]. Elements [BH86]. Elimination [MGS85]. EMEPS [LK86]. Emory [Bor87]. Empirical [Cas85, LV89, WF80]. End [KLV89]. Enforcement [VPO^{+}85]. Engineering [Nel81, WDS84]. Engineers [Hog85]. Enhanced [Sav82]. Enhancements [GG80b]. Entropy [Shi88a]. Entry [DP85]. Environment [BMM83, Hun83a]. Environmental [Ern83, Lod85, Lov83, Mil87, MCJ87, SC86, Str82b, Tia83, Yel83b, Pri87, Ros87]. Epidemiologic [CG82, FS89]. Epilogue [Wal83]. Equality [DCB88, Kem89b, Lit89, PS89b, Sea89b, PS89a]. Equation [Ehr83a, Rei80]. Equations [CA87, Fre83b, FP89, Wom85]. Equivalence [MGS85, Poo80, SK85, WB81]. equivalences [SSH81]. EQUORM [Har86a]. Error [Ano87i, CC89, Fle82, Gus86, LAH89, Pav88, Wai83, Bly80, Spi86, Spi87]. Errors [Whi89, Duf86]. Errors-in-Variables [Whi89]. Establishing [SP85a, WBH^{+}88]. estimability [EW85]. Estimate [Cha84]. Estimated [Ame85, Ano88e, Coh82, SYM84, SYC88]. Estimates [Ano89e, BM88b, Har86c, Mur88, SK85, SCR88, Whi89]. Estimating [Ano82e, Gre83a, LR83, MCJ87, Pen81, Sti86, Tho87, Dah81]. Estimation [AJ88, Ano86f, Bel87, CBJ86, FS89, HTM89, KM85, KM88, LK86, Loy83, Mee87, Mur86, Pau85, Ray85, SZM87, SM83, ST83, Sto81, WT80, CJT87, CH85, CH86, Cox84, GD89, Gol86, JH83, Kah87, SR83, VM89]. Estimator [Kem89b, MN87, Nor84, PS89b, Sea89b, Bly80, Kot86, PS89a]. Estimators [Ano86g, Ano87h, Ano89h, DJD85, Did85, JN89, Leh80, TW86, BvE88, Kar85, Kar86, Kar87, MA84, SI82]. Estrogens [Juk82]. Ethical [Ell83, IHD^{+}83, Ad 83, Bar83a, Boa83, Bro83b, Fre83a, Geh83, Gre83b, Kem83, Kis83, Mar83d, Mos83, Rem83, Ric83, Rob83a, Sol83]. Ethics [Bar83a, Boa83, Bro83b, Fre83a, Geh83, Gre83b, IHD^{+}83, Kem83, Kis83, Mar83d, Mos83, Rem83, Ric83, Rob83a, Sol83]. Evaluate [Ham81]. Evaluating [FM89, KHK89, Mar81a, NN82b, NN82a, VW82]. Evaluation [BF83, GW81, Juk82, Nag81, SD89, TW86]. Evaluations [Ste85]. Event [BT86]. Events [HC86, Jac84, KB84b, PKG86, EG87]. Everyday [Sch84a]. Everyone [Che86, Efr86a, Lin86b, Mor86, Pre86, Smi86, Efr86b]. Evidence [BS83, CG82, Die83b]. EXACHI [Hod84a]. Exact [Dav86, FS89, Lac84, Shu88]. EXACTB [Shu88]. Example [DP87]. Examples [DJD85, Joi81, KB84a, MP88]. Execute [NZ88]. Exercises [SF80]. Exhaustive [KB84b]. Existence [PC85]. Expanded [MBM^{+}85]. Expansion [Col83]. Expectations [Sch89a]. Expected [Bly80, Bol85, EL88, Hun89, Sav82]. Expected-Net-Gain [Bol85]. Expenditure [Coh82]. Experience [Bar83b]. Experiment [Dun88b, SPC87]. Experimental [Arm81, BPT82, Bor82a, Dem86, Fra81b, Hei81b, Hei86, NW81, Pay81, SD89, Wel85, Hei81a, Bor82b, Dat82, Wal82]. Experiments [Car81a, CC89, Del86, Ela84, Gil84, Koe81, MBM84a, MBM84b, Web87, Whe85, Box80b]. Expert [Bar83b, FSL82, Gen85, Hah85a, Hah85b, Tuk85c, Vel85, Tuk85a]. Exploration [Lin80]. Exploratory [Alb89, FC81, FTL84, LS89, Ros81a, Ros89, Sti80, Tuk80]. Exponential [BS87, GSW82, Ima82, Man87, Nor84, REZ88, Sha86a, Gle89, McC88]. Exponentiality [Swa84]. Exposition [BL85]. EXPOSMOO [Sha86a]. Exposure [CG82, Die83b, ONC85]. Exposures [Ern83, Lov83]. Extended [MRP^{+}89]. Extension [Lev80]. extensions [Far89]. extrapolations [MS87]. Extreme [Bur86, Kin89]. Extreme-Value [Kin89]. Eye [MSTY81].

F [Han85, Job82, Lac84]. Fable [Ram86]. Fables [FSL82]. Faces [TT80, TT80]. Facility [Hei86]. Fact [Mar83b, Mar83c, Yel83a, KLV89].

Factor [Coo83, Har86b]. Factorial [Car81a, Del86, Her86]. Factorization [Hin80]. Factors [AG84, HPB82]. Failure [Gre86b]. Families [BS87, Man87, Gue87, McC88]. Family [DK82, DK88]. Fast [Gle86a]. Fate [Kem89a]. Favor [Cov89]. FDA [ZR82]. FDSM [Che88]. Federal [BMM83, Bis84, BCF^{+}81, Bon81, Bon83a, Bra82, Bro84, Dun81, EWW^{+}82, Gon80, Mar81b, MO84, Mos81, Mos82, PHA84, Ros81b, Sne82, Tae81, Wal83, Zah82, Bon83b, DeM83, Fie83, Sla83]. Ferber [PS82]. Fertility [Lit88]. Field [BB85]. Fields [Hod87]. File [Smi84, Wai80]. Files [Dal87b, KF89]. Film [Boe81]. Filmstrip [BF81]. Filter [Did85, MS83, MS87]. Find [Nor84]. Findings [Ano80e]. Finite [Bel87, Lin81, RJ82, Buo87, PG82]. First [PC85, Rog84, Sal85b, Saw84, Her86, ME84]. first-order [ME84]. Fisher [Box80b, Box81, Haw89]. Fisheries [BMP^{+}85, DH85]. Fit [CA87, FM89, GH84, JSL89, Kin89, TC88, Smi85]. Fits [BAP89]. Fitting [Has81, Haw80, LV89, Mac86, Mil84, MSTY81, Smi85, Str80, Woo80]. Fixed [BA87]. Fluorocarbons [Sto82]. Foibles [FSL82]. Following [MD83]. Force [Hog85]. Forecasting [DH85, Lev87a, Mak81, Pit80, Rei87, Sch84b, WL85, Bol86]. Forestry [DH85]. Forests [BMP^{+}85]. Form [Kem80]. Formal [Sav82]. Formula [Ano89f, Hef88, Hu88, PKG86]. Formulas [Gon80, Gue81, KH89, Har80]. Formulation [VPO^{+}85]. FORTRAN [Ant80, Chu88, Dal89d, HPB82, TT80]. Foulkes [Dal89d]. Four [Fol82, KB84a]. Fractional [Che88]. Fractional-Differencing [Che88]. Framed [Dun87, Dun88b]. Francis [Wei83]. Free [DGS^{+}89, Kon87, GS86]. Freedom [CC89, DJ84b, Hod84a]. Frequencies [Hod84a]. Frequency [FC87]. Friendly [Nor87, Wal87]. Front [Ano80h, Ano80i, Ano80j, Ano80k, Ano81e, Ano81f, Ano81g, Ano81h, Ano82g, Ano82h, Ano82i, Ano82j, Ano83e, Ano83f, Ano83g, Ano83h, Ano84k, Ano84l, Ano84m, Ano84n, Ano85j, Ano85k, Ano85l, Ano85m, Ano86j, Ano86k, Ano86l, Ano86m, Ano87k, Ano87l, Ano87m, Ano87n, Ano88j, Ano88k, Ano88l, Ano88m, Ano89j, Ano89k, Ano89l, Ano89m, KLV89]. Full [Cle84a]. Function [Fle82, FR85, KE83, Lin86a, Lin80, MJ84, Pol80, Pos89, Rup87, CDFP81, McC88]. Functions [JM85, BL87, Hel88]. Funds [Gon80]. Further [HA89]. Future [Boa82b, Gen85, Hah85a, Hah85b, Hah89, Tuk85a, Tuk85c, Vel85].

Gadfly [SB85]. Gain [Bol85]. Galton [Wei83]. Game [KC84]. Games [Cov89]. gamma [Gle89]. GANOVA [BW84]. Gauss [dJ84a, KK84]. Gaussian [Dav80, WS87]. Geisser [Zel85]. Gemmell [DM81]. General [BW84, Bur82, CA87, Mul82, Spi85, Str80, WM87, Wil80, CJT87, Her80]. Generalization [AL86, Kra88, Man81]. Generalized [MLB80, Mer89, Rei80, Sea84, Som86, SA87]. Generate [IS85]. Generated [JM85]. Generating [BB84, KE83, TT80, CDFP81]. Generation [DR83, DK88, NP82, SF80, PK82]. GENERATOR [ES86]. GENERIC [Dal88a]. Geographically [ED84]. Geometric [Sha82, Stu82, Sch85b].

Geometry [Ano87f, Bry84, SW86, Swi81, Her80]. George [AS86, Ano84h, Ord84]. Geostatistics [Cre89]. Gerard [Ber87b]. Gertrude [MM80]. Getting [Hoo80]. Gilbert [MMM82]. GLIM [NN86]. glimpse [Bol84]. GMDH [Far81]. Goal [BB85]. Goldberger [Did85]. Golf [HC86, KC84]. Good [Loy83]. Goodness [CA87, FM89, Kin89]. Goodness-of-Fit [CA87, Kin89]. Gosset [Bol84, Box81]. Government [Bra82, EWW^{+}82, Mos82, Pit80, Sne82, Zah82, Bra82, EWW^{+}82, Mos82, Sne82, Zah82]. Governments [Bon88, Dun88a, Leh88, Rey88]. Graduate [BS89, Bro80a, BK86, EJKS87, VD84]. Graduate-Level [VD84]. Graph [CM83]. Graphic [Pai85]. Graphical [Alb89, ASS80, Ano85i, BL85, Cle84a, Dun88b, FTL84, HH87, Mag82, Pol80, TW86, KD87]. Graphics [Bea83, DP87, GH82, Hin88, Pol82]. Graphs [Cle84b, Dri80, Ima82]. Greenberg [AS86]. Group [Ame85, BR84, FR85, Loy83, RR85]. Groups [LS89]. Growth [Ami87, DGS^{+}89, MO84, REZ88, SK85, DBS^{+}85]. Guidelines [Ad 83, Bar83a, Boa83, Bro83b, Ell83, Fre83a, Geh83, Gre83b, IHD^{+}83, Kem83, Kis83, Mar83d, Mos83, Rem83, Ric83, Rob83a, Sol83]. Gurland [AL86].

H [TG89]. Haenszel [Som86]. Handling [BT86]. Happened [Sch86]. Harrison [Bol86]. Hartley [Smi81, Edw86]. Harvey [SPC87]. Hausdorff [KS88]. Hawaiian [Sch81b]. Hazard [Lin86a, SB85]. Health [Bro83a, Col81, GL83, Hac87, Ham83a, Ham83b, Joh83, Lev87c, Naj83]. Heisey [Ber87b]. Helpers [Wai83]. Herman [Smi81]. Heterogeneity [VY85]. Heterogeneous [Ros81a]. Heteroscedastic [Lie80]. heuristic [WS87]. Hierarchical [Gil86, Har86b, KL83, Pei87, RM80, Rom87]. HIFACT [Har86b]. High [Pos82, RM88]. Highlights [Van80]. Histograms [Dic81]. Historical [Ell83]. Histories [BT86, PBM84]. History [Cha80, DG86, Hut82, Her80, Her86, WS80]. Hoc [Bar83a, Boa83, Bro83b, Fre83a, Geh83, Gre83b, IHD^{+}83, Kem83, Kis83, Mar83d, Mos83, Rem83, Ric83, Rob83a, Sha86b, Sol83]. Holtzman [Bor82a, Bor82b, Dat82, Wal82]. Homework [SF80]. Homogeneity [Ami87, Har86a, HL87]. HOMPRO [HL87]. Houston [Hac87]. Human [JM85, Tro86, WF80]. Hyperbolic [MN88]. Hyperbolic-Secant [MN88]. Hypotheses [SM83, Hel88]. Hypothesis [JW82, Jer88, KMM87, MD83, Wil80, Nai87]. Hypothesis-Testing [Jer88].

IBM [Con87, Dal88b, Dal89a, Ell88, Gre85, MEP^{+}86, Sau84a]. IBM-PC [Con87]. IBM-PC-Compatible [Ell88]. Icicle [KL83]. Icon [TW86]. Idea [Hun81]. Identification [Bar85, JW82, Wom85]. Identifying [Rue85]. Identities [Roc88]. If [Sei89]. II [LMM84]. IIe [LMM84]. Illnesses [Lev87c]. Illusion [CM83]. Illustrate [Bol85]. Illustrated [Swi81]. Illustrative [MP88]. Imaging [HE84]. IML [NZ88, Sal88]. Immune [Web87]. Impact [CS86]. Imperfect [KM87]. Implementation [VPO^{+}85]. Implementations [FHI89]. Implications [DBM86, Mar82]. implicitly [BG89]. Importance [Ano87e, Kru87, KM89b]. impressions [Noe84]. Improved [AS84, Ano82e, Ano85e, KLV89, Dah81]. Improvement [Mil87, Pri87, Ros87]. Improving [BCF^{+}81, Ros81b, Bon81, Dun81, Mar81b, Mos81, Tae81]. IMSL [GG80b]. Including [Del86, Juk82]. Incomplete [Cha82, Hin80]. Incorporate [RR85]. Increasing [RM88]. INDEP [HG83]. INDEP-SELECT [HG83]. Independence [Ano88g, Bro86, Cox82, Flu86, Gil81, Kor84, SD87]. Independent [DCB88, EG87, KMM87, KB84b, Lit89, RNT84]. Index [Dal89d]. Indifference [Nov84]. Individual [Lin86a, PN87, YTL80]. induction [SP85a]. inductive [FF80]. Industry [BHH^{+}80, Boa85, Box80a, Bra83a, Bra83b, Car83, God85, Hah85c, Hil83, Hun83b, Joi85a, Joi85b, Leo83, Mar81a, McD83, Min80, Ott83, Par80, Sne84, Wac85, Wat80, BHH^{+}80, Box80a, Min80, Par80, Wat80]. Inequality [Ano88e, Hil84, Rag86, Spi85, Far89, NW86, SYM84, SYC88, WS80]. Inference [Kap87, RR84, RBL86, Edw86, GP89]. Inferences [Sch89a, Gue87]. Inferential [MRP^{+}89]. Infinite [Tor86, AT86]. Influence [Bis84, Bro84, PHA84, Rup87]. Information [Ano80l, Ano81i, Ano82k, Ano84o, Ano86n, Ano87o, Ano88n, Ano89n, Ber88, Hil88, Jay88, KM87, Kul88, McC88, Sau84a, TC88, Zel88a, AO89, Zel88b, dCS86]. Information-theoretic [TC88]. Initial [Hog85]. Input [IS85]. Inquiry [WF80]. Inspection [Gol82, Var86, Sti86]. Instruction [Car87, DSAGR87]. Instructional [Dri80, SJ86]. Integer [Ano82e, CDFP81, Dah81]. Integer-Parameter [Ano82e, Dah81]. Integral [Bas81, Low83, Roc88]. Integrals [Cha82]. Integration [Rai82]. Intelligent [Gen85, Hah85a, Hah85b, Tuk85c, Vel85, Tuk85a]. Intelligibly [Bro80b]. INTELREG [GNS82]. Interaction [Del86]. Interactive [BH81, Cor87, Dem86, DP85, Ell88, Geb89, GNS82, Ham85, HN86, KG80a, Lun80, Lun84, Mac86, MD85, Orr80, PBM84, Pol82, PC83, Rah81, Rom87, Sal88, Sha86b, Tur87b, Wai80]. Interbattery [HPB82]. Interblock [Har86c]. Intercept [Haw80, Job82]. Interesting [MN87, ME84]. Interim [Ber87c]. International [LH81]. Interpretation [HO85, McC80, Nel85, Woo84]. Interpretations [Bro83a, GL83, Ham83a, Ham83b, Joh83, Naj83]. Interval [HA86, Jen87, WT80]. Intervals [Kon87, Lou81, Sch81a, Ang89, Ang90, Ber81, Buo87]. Intra [Har86c]. Intra-Interblock [Har86c]. Intraclass [SZM87, ST83, SP86]. Introduction [Cas85, Low82, Cox84]. Introductory [Nea82, Noe80, WMB82]. Invalid [Jac84]. Inverse [Ano85g, GK84, MLB80, Dav80, Rav85, WS87]. Inverses [Sea84]. Inversion [MGS85]. Inverted [LS88]. Investment [Boo84]. Investment-Oriented [Boo84]. Invited [BZ89, Gat89, Kem89a, Sea89a]. Iowa [DS89]. Iowa-Curve [DS89]. Isn’t [Che86, Efr86a, Lin86b, Mor86, Pre86, Smi86, Efr86b]. ISP [Cor87]. ISPAHAN [KG80a]. Issue [Coh86]. Issues [Jer88, Lit88, Wei82a]. Iteratively [Ano89e, SCR88]. IV [Van80]. Ivakhnenko [Far81].

Jack [Bec82]. jackknife [EG83]. Jacob [Wei82b]. James [KM88]. Jeffreys [Kas82, Zel82b, Zel82a]. Jenkins [Rei80]. Jensen [Hil84, NW86, Spi85]. Jerome [GH80]. Jerzy [LR82]. John [MMM82]. Joint [Sti84]. Jointly [LR83]. Jonckheere [PS81]. Jose [Gei84b]. Journals [Bis84, Bro84, PHA84, Sal85a]. joy [GM86]. Judge [Jen87]. Judgments [FSL82]. Junior [Pos82]. Justifying [Jac84].

Kalman [Did85, MS83, MS87]. Kalman-filter [MS87]. Kantorovich [Rag86]. Kempthorne [PS89a]. Kendall [Ano84h, Ord84]. Key [Boa85, God85, Hah85c, Joi85a, Joi85b, Wac85]. Kickers [BB85]. Kiefer [Bec82]. Kit [KF89, Whe85]. Knowledge [Sto82]. Kolmogorov [Har80, PG83]. Kronecker [Rog84]. Kruskal [Sti84]. Kullback [AO89]. Kurtosis [BM88a, Moo86, Rup87]. Kutner [Dri86]. Kuznets [Ham87a]. KWIKSTAT [Ell88, Ell89].

Lack [JSL89]. Lack-of-Fit [JSL89]. Language [Sal88]. Large [Fle82, HA86, Hod84a, Hid86, Tom84]. Large-Sample [Fle82, HA86]. Late [Sei89]. Latent [DM87, LK86]. Latent-Class [DM87]. Later [Ano86e, Hoe85, MGS82]. Latin [Mur86]. Latitude [Smi84]. Lattice [Car81b]. Law [Dow85, HH87, Hun83a, Mil87, Pri87, Rob83b, Ros87, VPO^{+}85, Tom84]. Lay [FSL82]. Layout [Del86]. Layouts [Car81a, Car81b]. Lead [BH83]. League [BF83]. Least [Ano89e, BM88c, Ehr83a, Har86c, HO85, Kem89b, PKG86, PS89b, Sea89b, SCR88, WS88, Wil81, Woo80, BvE88, MA84, PS89a, SSM80, Wom85]. Least-Squares [WS88]. left [Mul88]. left-truncated [Mul88]. Legal [Mar82]. Legitimate [Var86]. Leibler [AO89]. leisurely [EG83]. Leonard [Rya82]. Lessons [Lav82, Pen88]. Letter [Ano88f]. Letters [Ano84e, ADS81, BSB^{+}86, BGF^{+}81, BBC^{+}89, BDR^{+}88, CPE^{+}84, CHP^{+}89, ENH^{+}80, EFJ^{+}89, FSW^{+}87, FSJ^{+}89, FSM^{+}80, GBL^{+}87, GFL^{+}85, GN85, GLM^{+}81, GMB^{+}84, HLT^{+}87, HRP^{+}86, JEE^{+}83, JKL^{+}81, KKN^{+}80, KHH^{+}86, KLC^{+}88, LBW^{+}88, MHP^{+}87, MKF^{+}86, NMF^{+}80, NBA^{+}82, PSD^{+}83, RSI^{+}84, SCC^{+}84, SB83, SBE^{+}82, WSA^{+}85, vdLFS^{+}88]. Level [Bro83a, GL83, Ham83a, Ham83b, Joh83, Naj83, VD84]. Levels [BH83]. Leverage [Cas83]. L’Hospital [Tar81]. LHS [IS85]. Liberal [Ive85, MR89]. Library [Con87, DR83, DK88, GG80b, Gle86a]. Life [MD81, RM81, Wag82, Gol86, MS87, RM81]. Lifestyle [Lev87c]. Lifestyle-Related [Lev87c]. Light [Swa82]. Like [Dal88c]. Likelihood [Ano82e, DJD85, FS89, Leh80, LR83, Nor84, Ray85, Ber87c, Cox84, Dah81, Dav80, McC88, VM89, Hil87]. Likert [Kis82]. Lilliefors [Ano87g, DW86, Ima82]. LIMDEP [Gre83a, Gre85, Gre86b]. Limit [Tar81, Yan88, PG82]. Limited [Ano86h, Gre83a]. Limiting [Kor84, SD89, Web87]. Limits [FBC^{+}83, MD83, Whi86]. LINCHI [Hod84b]. LINCS [Sch88a]. Line [TT80]. Linear [Ano89h, Bur82, CA87, Dal87a, FR85, GH84, GNS82, Gro82, Hod84b, JN89, Kem89b, Lam80, Lit82, MLB80, Mul82, PS89b, Rag86, RT81, Sav82, Sch88a, Sea84, Sea89b, SA87, Thi87, Thu89, War85, Wil80, Woo80, BvE88, Bol86, CW88, CJT87, Duf86, Ham84, Her80, Kot86, MA84, PS89a, SSM80, TC88]. Linearly [RNT84]. Lines [MSTY81, Sea88b]. LINWOOD [Woo80]. Listing [Nef83, SP84, SP85b]. Literature [Bas81, KM89b]. Local [Bon88, Dun88a, Leh88, Rey88]. Locally [Cle81]. Location [Gus86, Smi84, Spi85, JH83]. Log [Dal87a, Sav82]. Log-Linear [Dal87a, Sav82]. Logging [Cle84a]. Logistic [Ano84f, CCW^{+}86, Dal88b, Dal89a, Hau83, dCS86, Dal88b]. LOGIT [Ste88a]. LOKATE [Smi84]. Longer [Cov89]. Longitude [Smi84]. Longitude-Latitude [Smi84]. Longitudinal [War85]. Look [BPT82, Est85, FP87, Hah89, Has87, KS83, RN88, CJT87, EG83]. Lord [Hil84]. Loses [Sch84c]. Lotteries [Est85, KS83]. Love [SC86]. Lovell [Lov84]. Low [Bro83a, Ell88, GL83, Ham83a, Ham83b, Joh83, Naj83]. Low-Cost [Ell88]. Low-Level [Bro83a, GL83, Ham83a, Ham83b, Joh83, Naj83]. LOWESS [Cle81]. LSAP [Lam80]. Lung [BS83, Die83b]. Lurking [Joi81].

M [Gei84b]. Machine [BT86]. Macintosh [Hun89]. MAD [Hil84, JD89]. Made [Sch84b]. MADMANager [KF89]. main [GS86]. Mainframe [Ami87, Gal84]. Mainframes [Bak84]. Maintenance [Wai80]. Major [BF83]. Make [Yel83b]. Making [Str82b, GW80]. Management [DP85, KF89, ONC85, Van80]. Managers [Rob87]. Managing [Mar84]. Manipulating [Hel88]. Manipulation [KF89]. Mantel [Som86]. Manufacturing [Hah89]. Map [GH82, Smi84]. Mapping [Dun87, Dun89]. Maps [Dun88b, Tru81, WF80]. marginal [SSM80]. marginals [GM86]. Marine [BMP^{+}85, DH85]. Mark [Hsi81]. Markov [KK84, MP88]. Mary [MM80]. Matched [RR85]. Material [Shu83]. Materials [Dow85, Lod85, Pos81, Rue85, SP84, SP85b, WL85]. Maternity [PBM84]. Mathematics [Shu83]. MATHIST [PBM84]. Matrices [Har86a, Pol80]. Matrix [Ame85, Ano85g, Low82, MGS85, Sal88, Tur87b, dJ84a, Rav85, dCS86]. Matter [Ano80a, Ano80b, Ano80c, Ano80d, Ano80h, Ano80i, Ano80j, Ano80k, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano82a, Ano82b, Ano82c, Ano82d, Ano82g, Ano82h, Ano82i, Ano82j, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano84a, Ano84b, Ano84c, Ano84d, Ano84k, Ano84l, Ano84m, Ano84n, Ano85a, Ano85b, Ano85c, Ano85d, Ano85j, Ano85k, Ano85l, Ano85m, Ano86a, Ano86b, Ano86c, Ano86d, Ano86j, Ano86k, Ano86l, Ano86m, Ano87a, Ano87b, Ano87c, Ano87d, Ano87k, Ano87l, Ano87m, Ano87n, Ano88a, Ano88b, Ano88c, Ano88d, Ano88j, Ano88k, Ano88l, Ano88m, Ano89a, Ano89b, Ano89c, Ano89d, Ano89j, Ano89k, Ano89l, Ano89m]. Maurice [Ano84h, Ord84]. Maximizing [LR83]. Maximum [Ano82e, DJD85, FS89, Nor84, Ray85, Shi88b, Cox84, Dah81, VM89]. Mean [Ano87p, Ano88e, EL88, Hun89, Kap87, Sch89a, Yan88, Aso80, Kot86, ME84, SYM84, SYC88, Shu82]. mean-of-ratios [Kot86]. Meaning [Moo86, Roy86]. Means [ASS80, BB89, Pen81, Sch84a, Shi88a, SSM80]. Measure [Gam81, Gus86, Hor83, Ped85]. Measured [Sei89, TVV85]. Measurement [Ano87i, Ern83, Spi86, Spi87]. Measures [BR84, Ela84, Gil84, Hef88, LS89, MBM84a, MBM84b, Rav86, TC88]. Meat [ZR82]. Median [HTM89, JD89, DHPR88]. Medical [Car87, Coh82, DSAGR87, GW81, Hac87, Sal85a, Hac87]. Medicine [Bor87, Col81]. MEDPOL [DHPR88]. Meharry [Car87]. Meinhold [HA89]. Melvin [Gei84b]. Members [DL86]. Memoriam [AS86, Ano84h, Bec82, Dav81, DM81, DS81, FK87, Ham87a, Kis82, MMM82, Ord84, PS82, Rya82, Smi81, TG89, Wei82b]. Meteorological [Pik85]. Method [Ano82e, Ano87f, Ano89f, Hu88, Kap87, Mag82, SW86, BG89, Cox84, Dah81]. Methodology [DS89, Die83a, KC84, Kra88]. Methods [Alb89, Bro80a, Cle84a, CA85, FTL84, HA86, Low82, MP88, RR85, SBA85, Tia83, WM87, PK82]. MICRO [Edm85, Fow88]. MICRO-CLUSTER [Edm85]. MICRO-EBA [Fow88]. Microcomputer [Ber87a, DM87, Edg84, Fer88, FS89, Gus86, Len87, LK86, Roh87, SJ86, Sch88a, Swa84]. Microcomputers [Edm85, Geb81, Gil86, Gle86a, HN86, Hin82, Hin88, KF89, Nef83, Orr80, Rah81, Ste85, Wal87]. Microcomputing [Dal88c]. Micros [Bak84]. MICROSCALE [Tro86]. MICROSTAT [Orr80]. Miner [Lov84]. Miniature [KC84]. Minimum [Ano86g, Ano87h, GD89, Kar85, Kar86, Kar87, VM89]. Minis [Bak84]. Miscellaneous [Dal89b]. Missing [Ano84g, Don82, Fre80, JW82, LR83, Mur88, SZM87, Mur86]. Missing-Cell [Mur88]. Misspecifications [MT82]. MIX [Mac86]. mixed [HT83]. Mixture [SPC87, Gle89, PK82]. Mixtures [Mac86]. MLE [Lev85]. MLIB [Con87]. MLIN [Gro82]. Mo [SYC88]. MODBIN [DM87]. Model [Bel87, BCV88, Che88, GH84, Gro82, JM85, Lev87c, Lin86a, MLB80, Mul82, Pau81, Pau85, Pen88, Sch84a, SZM87, SM84, ST83, Bol86, CJT87, Her80, HT83, Kot86, ME84, SSM80, SSH81]. Model-Based [Bel87]. Modeling [Gre85, Lam83, PBM84, DBS^{+}85]. Modeling-From [Lam83]. Models [BF83, DM87, DBM86, Dri86, FM89, Gre83a, HO83, Lie80, Lit82, Neu85, Pei87, PC83, Rei80, Sta86, SD86, Sea84, SM83, SK85, Str80, War85, Cox84, GP89]. Modern [Low82, Var86]. Modest [Sti82]. Modified [Har80]. MODMULT [DM87]. Module [MRP^{+}89, Ste89a, Ste89b]. Modules [Fol82]. Moment [Arb89, KE83, PC85, CDFP81]. moment-generating [CDFP81]. Moments [BM86, Dic81, Fu84, Lin81, NA85, RJ82, CDFP81]. Monetary [Bar84]. Monitor [LNC^{+}81]. Monitoring [BMP^{+}85, Mil87, Pri87, Ros87, SC86, VR83]. Monotone [Rav86]. Monotonicity [FM84, Shi88a]. Monte [Gle86a]. MORPOL [Lin85a]. Mortality [MP80, MEP^{+}86, MRP^{+}89, Whi85]. Mosaic [HK84]. most [SS84]. Motion [Chu88, Neu85]. Moving [Tor86, AT86]. MR [Kar86, Kar87, SYC88, Zel85]. MS [LAH89, EL88, Gol89, Hin88, Hud87, LAH89, PH86]. MS-DOS [EL88, Hin88, Hud87]. MS/PC [Gol89]. MS/PC-DOS [Gol89]. MSUSTAT [Lun80, Lun84]. MSUSTAT-An [Lun80]. MTS [Rei87]. MUDIFT [DGS^{+}89]. Multibased [Cle84a]. Multidimensional [Dal87a, Lin85a, YTL80]. multifactor [GS86]. Multinomial [BB84, Fle82, HC86, Tho87, Ang89, Ang90, Cox84]. Multiparameter [Nov84]. Multiple [Ano84g, Don82, HT83, Job82, NZ88, Pav88, Sto81]. Multiple-Regression [Job82]. multiprocess [Bol86]. Multivariate [Ano85g, BW84, Bur82, DGS^{+}89, Daw89, FR86, FTL84, Gol82, Gro82, Hin80, IS85, Job82, Mer88, MM84, RJ82, Ray85, Roh87, RR85, Sch84b, TW86, Wil80, Rav85, RC88, TC88, WC87]. Must [Ehr83b, Ehr84, Shv84]. MVSPEC [MM84]. My [Bar83b]. MYSTAT [Dal87b, Dal88a, Wil87]. MYSTAT/SYSTAT [Dal87b].

N [Ano87f, SW86]. N-Dimensional [Ano87f, SW86]. Name [Smi84]. National [Coh82, Daw89, MGS82]. Nations [Rob83b]. Natural [SB85]. NCSD [Han83]. NCSD-Noncentral [Han83]. NCSS [Hin88]. NCSS-Graphics [Hin88]. nearness [Ped85]. Necessarist [Kas82, Zel82b, Zel82a]. necessary [MA84]. Need [Tuk80]. Negative [PC85, CDFP81]. Nested [GG80a]. Net [Bol85]. Neter [Dri86]. Networks [BS85]. Neyman [LR82]. NFD [Han85]. Nile [Ano89h, JN89]. Nimis [Ber87b]. No [Lou81, Ang90, CH86, GY90, Kar86, Kar87, SYC88, Spi87, Zel85]. Noise [FC87]. nonapplicability [PG82]. Noncentral [Coh88, Han83, Han85, Hjo88, Pos89, RD88]. Nonestimable [BL87]. NONLIN84 [Sta86]. Nonlinear [LV89, Sta86, Woo80]. NONLINWOOD [Woo80]. NONLINWOOD-Linear [Woo80]. NONMEM [BS80]. Nonnegative [Ame85, Kap87, MJ84, Mul88]. Nonnormality [Ham84]. Nonoptimally [BM88c]. NONPAR [HPM89, Zar80]. Nonparametric [Ami87, CI81a, CI81b, Die89, Fli81, HPM89, Kap87, Lev80, Noe81, REZ88, Shu82, Zar80]. Nonparametrics [Noe80, Noe84]. Nonprobabilistic [RBL86]. Nonuniform [ES86]. Nonunique [DJD85, Lev85]. Nonwhite [CS86]. Normal [GSW82, Gue81, Hef88, Ima82, Job82, LG85, Low83, Mag82, MT82, Nor89, Ray85, SS89, Sha85, Sti82, Whi86, Ham84]. Normality [Ano87g, Bro86, DW86, HW86, LV89, AO89]. Normals [Riv81]. North [Boa85, God85, Hah85c, Joi85a, Joi85b, Wac85]. Note [All87, Ang89, Ang90, Ano84f, Ano84j, Ano89e, BM86, BG87, BM88b, BN88, Bur82, Che82, Fre83b, FP89, Gen86, Hau83, Haw80, HW86, KHK89, KE83, KB84b, Lev85, MN88, MD83, SCR88, Tor86, WS88, AT86, Aso80, Buo87, CW88, Kah87, Ped85, dCS86]. Notes [Gle86b]. Nothing [Bro80b]. Novick [Gei84b]. NTSYS [Roh87]. NTSYS-pc [Roh87]. Nuisance [Alb89]. Null [MD83]. Number [DR83, DK88, Hod84a, Hin82]. Numbering [Hod87].

Numbers [RR83, Rea85, SSH81, Tom84]. Numeracy [Ehr81]. Numerical [KD87, Lin85a, Roh87]. Nursing [Bro80a].

Objective [Lin80, Mag82]. Observation [Whi86, Wil81]. Observations [Coh88, Sch89a, SZM87, Wat82]. Observing [Lou81]. Obtaining [Edg84]. Occupational [BH83, CA85, MP80, MEP^{+}86, ONC85]. Occurrence [PKG86]. OCMAP [MP80, MEP^{+}86, MRP^{+}89]. OCMAP/PC [MEP^{+}86]. OCTA [Dal87a]. ODDJOB [Dal89b]. Odds [Sav82]. Off [BS89]. Off-Campus [BS89]. Offensive [BF83]. Office [WBH^{+}88]. Old [Pen88]. Olds [Pos81]. Olympics [FC81]. One [Ano88g, DJ84b, Kap87, PG83, Sto81, Gue87, SD87, Sch85b]. one-parameter [Gue87]. One-Sample [PG83]. One-Sided [Kap87]. One-Way [Ano88g, Sto81, SD87, Sch85b]. Only [Smi84]. Ons [Dal88a]. Opening [Ben88]. Operators [Lin81, RJ82]. Opportunity [BS85]. Optical [CM83]. Optimal [Mer89, Pen81, WC87, Zel88b, Ber88, Hil88, Jay88, Kul88, Zel88a]. optimality [Shu82]. Optimization [Dem86, Pau81]. Optimum [SPC87]. Optional [Gam81]. Options [YTL80]. Order [BM86, FP81, Fie84, GY89, GY90, ME84]. order-statistics [GY89, GY90]. Ordered [BH86, Man83]. Orderings [Ano87e, Kru87]. Ordinal [Hod84b]. Ordinary [Har86c, Kem89b, PS89b, Sea89b, BvE88, MA84, PS89a]. ordinary-least-squares [BvE88]. Organization [Coh82, ED84]. Organizations [GF80]. Oriented [Boo84, MP80, MEP^{+}86, MD81, MM84]. Origin [Cas83, CS86]. ORSIM2 [HPB82]. Orthant [KHK89]. Orthogonal [HPB82, Lin85a, Pin88, RNT84, Sch85b]. Orthogonalization [SBA85]. Orthogonalization-Triangularization [SBA85]. Oscar [PS89a]. OSHA [Tho82]. OSIRIS [Van80]. Other [Kem89a, Kon87, LK86]. Otto [Smi81]. Ought [Mul80]. Outliers [Shi88b]. Output [Sea81, SHF83, Sea88a]. Overall [PN87]. Overlapping [Tam84]. Overview [Coo83, HHF82, Ros81b]. Ozone [Atk82, Sto82, Wat82]. Ozone-Observations [Wat82].

p [Lac84, Edg84, Fir82, MPH88, Wai80]. P-DIST [Edg84, MPH88]. P-STAT [Fir82, Wai80]. PACK [Chu88, Rei80]. Package [BH81, Chu88, Edw84, Ell88, Geb81, Geb89, GH88, GNS82, Ham85, HPM89, Hen83, Hin82, Hin88, LMM84, Len87, Lun80, Lun84, MPH88, Nor87, Orr80, Per84, Rah81, Roh84, Sau84b, SJ86, Sha86a, Sti80, Str80, YTL80, Zar80, Zel89]. Packages [BR84, Mul80, Nef83, Sea89a]. Paired [BB89, Ros89]. Pairs [KMM87]. Pairwise [Sto81]. Panel [Ano85h, MFG^{+}84, Nea82]. Papers [Ehr82, FGHK83, Ham81, Hof84]. Paradox [Gas89, Hil84, Kai89b, Kai89a, Pai85, Sch89b, Wag82]. Parallel [Sea88b]. Parameter [Ano82e, Ano85e, Ano89h, JN89, Lou81, Spi85, SK85, AS84, Bly80, Dah81, GD89, Gue87, Kah87]. Parameters [Alb89, FS89, Gre83a, LR83, LK86, Pau85, SZM87, SM83]. Parametric [CI81a, CI81b, Fli81, Noe81, BL87]. Parker [MMM82]. Part [Bas81, BS89, Hjo88, CH86, Kar86, Kar87]. Part-Time [BS89]. Partial [DK87, Flu89, Kor84, Saw84, MC87]. partitions [Mee86]. Partnership [BS85]. PASS [Ena88]. Path [AC82]. Pathological [KB84a]. Patriotism [FC81]. Pattern [Joh81, KG80a]. Patterns [Hod87]. pc [Roh87, Ano86h, Con87, Dal86a, Dal88b, Dal89a, Ell88, Gre85, MEP^{+}86, Rom87, Sau84a, Smi85, Zel89]. PC-DOS [Gol89]. PC-SIZE [Dal86a]. PCA [Ram86]. PCBRAP [Zel89]. PCNONLIN [Sta86]. PCPROBIT [Wal87]. PCT [Dal89c]. Peakedness [Hor83]. Pearson [Dav81, Gil81]. PEC [Pec80]. People [Hoo80]. Percentile [PN87]. Percentiles [Dal89c, Dic81, Lin86a]. Perception [Dun88b, FC81]. perfect [DK87]. Perform [Gre86a]. Performance [BF83, Mar81a, Mos80]. Personal [BW84, Col81, Gal84, Wil87]. Personality [Tro86]. Perspective [Did85, Ell83, KM88]. Physics [Lam83]. Pitfalls [Ano87i, Spi86, Spi87]. Pitman [Ped85]. PL [GD81]. PL/I [GD81]. Place [Smi84]. Places [BDMW87]. Planning [BFM^{+}88, BPT82, BH81, CBI80, Hei86, Pit80]. PLANUNG [Hei86]. Platform [Mos80]. Player [Cov89]. Plot [Ela84, Gil84, Hoa80, Hun88, MBM84a, MBM84b, Ros81a]. Plots [BG87, KL83, LG85, Mag82, Ros89, Sea88b, Wil81, MC87, Sti86]. Plus [LMM84]. Pocket [Nor89]. Pocket-Calculator [Nor89]. Point [Ano86f, Mas89, VR83, CH85, CH86]. Poisson [GY90, Aic85, Fro81, FM89, GY89, NW86, She84]. Poissonness [Hoa80]. Policy [Bar84, Bon88, Dun88a, Leh88, Rai82, Rey88, SB85, VPO^{+}85]. Polish [DHPR88]. Pollutants [Lod85]. Pollution [Dow85, VR83, WL85]. POLYGM [SZM87]. Polynomial [Pei87, SZM87]. Polynomials [Col83, Lin85a, Nag81]. PONC [Mas89]. Pool [Ano88h, Bal87]. Pooled [Die83a]. Popular [KH89]. Population [Bel87, Kap87, MCJ87, Pit80, SSM80, Sto88, VY85]. Populations [Ano87p, DCB88, Poo80, Buo87, PG82]. Portable [KF89]. possible [Mee86]. Posttest [Ano89g, BK80, Sta88]. Power [Ban88, Ena88, Gol89, HH87, Len87, PT80, PS81, Shi88a, Wai83]. Power-Law [HH87]. Powerful [KF89, SS84]. PowerPack [Len87]. Practical [KM85]. Practice [Ad 83, Bar83a, Boa83, Bro83b, Ell83, Fre83a, Geh83, Gre83b, Hun81, IHD^{+}83, Kem83, Kis83, Mar83d, Mos83, Rem83, Ric83, Rob83a, Sol83]. Practiced [Ehr83b, Ehr84, Sal85a, Shv84]. Preach [Ehr83b, Ehr84, Shv84]. Prediction [Kon87, LK86, Neu85, Pik85, SM84, Whi86]. Predictions [Car86]. Predictive [AJ88, Ano86f, CH85, CH86]. Preface [Can85, Lan85, Pat85, PHA83, Rus85a, Rus85b, Str82a]. Preference [FC81]. Pregnancy [CG82]. Preliminary [Gol82]. Preparation [Del86]. Preparing [BHH^{+}80, EWW^{+}82, Box80a, Bra82, Min80, Mos82, Par80, Sne82, Wat80, Zah82]. Presence [CA87, MCJ87]. Presentation [Cle84a, FHI^{+}88]. Presenting [FGHK83]. Preserving [FM84]. President [BCF^{+}81, Bon81, Dun81, Mar81b, Mos81, Tae81]. Pretest [Ano89g, BK80, Sta88]. Pretest-Posttest [Ano89g, BK80, Sta88].

Prevalence [Ami87]. PREVIEW [Lev87c]. Principal [Mul82, Stu82]. Principle [Coh86, Ber87c, Hil87]. Printer [TT80]. Prior [Ano85f, Ber84, Gei84a, Zel84, Gei84b, Zel85]. Privacy [Ano80e]. Pro [Sau84b]. Probabilistic [Gol86, Lam83]. Probabilities [Chu88, Edg84, HA86, MPH88]. Probability [BB85, Bry84, Hof84, Hud87, Jac84, KHK89, Lin81, LG85, Loy83, LW82, Mag82, MSGG87, PKG86, EG87, FF80, Boa82a, Pos82, Wat83]. Probit [IMS87, Wal87, Ste88a]. Problem [Ehr81, GW80, KM87, Pau81, Sch88b, Thi87, Wei82a, Wei83, Ano89h, JN89]. Problems [Boa82a, DBS^{+}85, HG83, Len87, McC80, Pit80, RR83, Rea85, Rue85, EW85]. Procedure [BB84, Hod84a, KGHS84, Lev80]. Procedures [AJ88, Ano84g, DCB88, Don82, Gam81, GW81, MRP^{+}89, Nov84, Pav88, Shu88]. Process [Mas89, GY89, GY90]. Processes [Chu88, MP88]. Processing [Ber88, Hil88, Hod87, Jay88, Kul88, Zel88a, Zel88b]. Processor [Int85]. PRODOS [IMS87]. Product [Mar84]. Production [ZR82]. Productivity [VD84]. Products [Rog84]. Profession [RM88]. Profession-Starting [RM88]. Professional [Bar83a, Boa83, Bro83b, Fre83a, Geh83, Gre83b, IHD^{+}83, Kem83, Kis83, Mar83d, Mos83, Rem83, Ric83, Rob83a, Sau84b, Sol83]. Prognostic [HG83]. Program [AC82, Ano86h, Ant80, BW84, Bro80a, Cam88, Che88, Cle81, Dal86a, Dal87a, Dal88b, Dal89a, Dal89d, Dem86, DK82, Edg84, ES86, Ena88, GG80a, Gil86, Gre86a, Gre83a, Gre85, GD81, Gro82, Gus86, Hal88, Har86a, HL87, Has81, HN86, Hin80, HPB82, Hun89, IS85, Lam80, Lie80, Mac86, MP80, MEP^{+}86, MD81, Mer88, Mer89, MM84, NZ88, Pec80, PBM84, PC83, Rah81, Rei80, RM81, Rom87, SD86, Smi85, TT80, Wal87, WM87, Wil80, dJ84a]. Programmable [Boa82a]. Programming [LM86, Rag86]. Programs [Ami87, Arm81, BS89, BF81, Cor87, Dal86b, DM87, FS89, Fra81b, Ham81, Hei81b, Hud87, Jer88, KF89, LK86, MD85, Nag81, NN82b, NN82a, NW81, Pay81, PH86, Roh84, Roh87, RM80, Ste85, SD89, Tur87b, VW82, Woo80, Hei81a]. Project [BCF^{+}81, Bon81, Dun81, Lev87c, Mar81b, Mos81, Shu83, Tae81]. Projection [Tuk85b]. Prologue [Wal83]. Promoting [BZ89]. Promotion [Lev87c]. Proof [JD89, Rag86, RD88, Sti84, Far89, GY89, GY90, PG82, Tom84]. Proofs [Ans85, FF80]. Propensity [RR85]. Properly [Hoo80]. properties [Ber81, Len83, SP85a]. Property [BE84, DS89, JD89, MN87, ME84]. Proportional [Gus86, Lin86a, MRP^{+}89]. Proportions [Fle82, HL87, Lit89, Tho87, Ang89, Ang90, Buo87, SR83]. Proposal [Sti82]. Psychometric [Ste87]. Public [Col81, RM88, WBH^{+}88]. Publications [Cle84b]. Purpose [Str80]. Purposes [Pit80]. Pursuit [Tuk85b]. Pyramid [Wai83].

QNumQ [Hod87]. Qualitative [Gre83a, IMS87]. Quality [Ano88h, Bal87, DP85, KM87, Mar84, Sal85b, Tan82a, VD84, VR83, Gup82, Hig82, Joh82a, Kna82, Leo82, Owe82, Swi82, Tan82b]. Quality/Value [KM87]. Quantitative [CA85, Juk82, PS83, Swa82, Moo81]. Quartiles [FP87]. Questionnaire [Fox83]. Questionnaires [Hod87]. Questions [Hod87, VR83]. Quick [Ans85, Ena88, HH87, Ang89, Ang90].

R [Gei84b, PS89a]. R. [Box80b]. Race [CS86, McC80]. Radiation [Bro83a, GL83, Ham83a, Ham83b, Joh83, Naj83]. Rain [FBC^{+}83]. Ramsdell [Rya82]. Random [BH86, BB84, CS83, DR83, DK88, ES86, Flu86, LW82, Mur88, Roc88, BG89, GP89, Ham84, PK82]. Random-Number [DK88]. Randomization [NZ88]. Randomized [Car81a, Car81b, Mur88, SD86, Mur86]. Range [Koc85]. Rangefinder [BG87]. Ranges [Kor84]. Rank [CI81a, FS87, KMM87, PN87, CI81b, Fli81, Noe81]. Ranking [Edw84]. rankings [DK87]. Ranks [PN87]. RANKSEL [Edw84]. Raster [Bea83]. Rated [BDMW87]. Rates [CS86, Pav88]. Rating [Ant80, BB85]. Ratings [AG88, HK84]. Ratio [ABS81, Che82, Dav80]. Ratios [Sav82, Kot86]. RATS [DL82]. RATS-Regression [DL82]. Real [Hun81, Wag82]. Really [Sch89a, SS84]. Reasoning [CBI80]. Recipe [GK84, SPC87]. Reckoning [Boe81]. Recollections [Col81, Noe84]. Recommendations [Ano80e, Bor82a, Bor82b, Dat82, Wal82, CGL83]. Records [Daw89]. Recounting [Har88]. Rectangle [Dun87, Dun88b]. Recurrence [Coh88]. Recursive [GH84, PKG86, CJT87]. Reducing [Mil84]. Reduction [Gus86, HE84]. Redundancy [Flu89]. Redundant [Ham87b]. Reexamined [Moo86]. Refereeing [Gle86b]. reference [Edw86]. Refinement [Ano85i, BL85]. Regarding [HA84]. Region [CCW^{+}86]. Regions [Car86, RC88]. Registered [KGHS84, KC84]. REGM [Wil80]. Regress [Sch89a]. Regression [Aic85, All87, Ami87, Ano84g, Ano86h, Ano87i, Ano89e, Ans85, BAP89, BN88, BCV88, Cas83, CA87, Cle81, Dal88b, Dal89a, DP87, Die89, DL82, Don82, DH87, Ehr83a, FC87, FR85, Flu89, Fre83b, FP89, Fro81, FM89, GH84, GNS82, Haw80, HO83, HO85, Job82, Lie80, McC80, Nel85, Pei87, Pin88, PC83, RT81, SF80, SK85, SCR88, SA87, Swi81, Thi87, Thu89, Tur87b, VW81, Wat81, Whe85, Whi89, WS88, Wil81, Zel89, CW88, Man82, Spi86, Spi87]. regressions [Duf86, TC88]. Regulatory [Atk82, Hut82]. Rejection [MD83]. Rejoinder [Bor82b, CI81b, Ham83b, Yel83a]. Related [Lev87c]. Relating [Die83b]. Relationship [BH83, KM87, Wil82]. Relationships [Lee86]. Relative [Ano84g, Ano87e, Don82, GH88, Kru87, KM89b, TVV85]. Release [NN86]. Relevance [KM85]. Reliability [BS85, Nel81]. Religion [Sal85a]. Remarks [BS87, Nea82]. Renewable [BMP^{+}85, DBS^{+}85]. Renewal [MP88]. Rensis [Kis82]. Reorganization [BCF^{+}81, Bon81, Dun81, Mar81b, Mos81, Tae81]. Repeated [BR84, Ela84, Gam81, Gil84, LS89, MBM84a, MBM84b]. replacement [Aso80]. Replicated [CA87]. Replicates [JSL89]. Reply [Ad 83, Bel84b, Bis84, Bon83b, Bra83b, Efr86a, Ehr84, Gei84a, Hah85b, Hei81b, Joi85b, Kai89a, MBM84b, NN82a, PHA84, PS89b, Tan82b, Zel82b, Zel88a, Efr86b, Gei84b, PS89a, Zel88b]. Report [Ad 83, Ano85h, Bar83a, BHH^{+}80, Boa83, BCF^{+}81, Bon81, Bor82a, Bor82b, Box80a, Bra82, Bro83b, Dat82, Dun81, EWW^{+}82, Fre83a, Geh83, Gre83b, Hog85, IHD^{+}83, Kem83, Kis83, Mar81b, Mar83d, MFG^{+}84, Min80, Mos81, Mos82, Mos83, Par80, Rem83, Ric83, Rob83a, Ros81b, Sne82, Sol83, Tae81, Wal82, Wat80, Zah82]. Reporting [HA84]. Reports [Ehr82]. Representation [Gas89, Kai89b, Kai89a, Pai85, Sch89b]. representing [Hid86]. Reproductive [NP82]. Requirements [Lod85]. Research [BK80, Bro80a, HO83, SB85]. Residual [Sea88b, MC87]. Residuals [DJ84b, GH84, Lie80]. Resolving [Ste82]. Resources [BMP^{+}85, DH85, Shu83, DBS^{+}85]. Response [Ami87, Ano84f, CCW^{+}86, Hau83, IMS87, KC84, Mar83c, NZ88, SD86, MS87]. Responses [CA87]. Restriction [MLB80]. Restrictions [Sea84]. Results [HA89, McC80, Thu89, CJT87]. Retirements [DS89]. Retrievable [GK84]. Return [FP89]. Reverse [Fie84]. Review [Ano80f, BM88a, Bar86, Bel80, Ber89, Bes87, BF81, Boa82a, Boe81, Bur87, BG80, CJ81, CT89, Chr81, Con89, Coo87, Dav85b, Dav89, Dix88, Ell87, Fol82, GW80, Gil80, Gol88, Hal89, HB86, Hil85, Hof84, KM89a, KP87, KG80b, KSB^{+}81, Koe81, Koe87, Koe86, Lac85, LeP87, Lev87b, Lig85, Lig88, Lin85b, Lin87, Lin88, LS89, Man88, Mar85, Mar83a, McB80, Mil89, Moo81, Mor87a, Mor87b, Nas86a, Nas86b, Nel81, O’B88, Par89, Pla86, Pos80, Pos81, Pos82, Pos83, PF89, Reg86, Sch85a, Shr87, Shu83, Shu86, Ste88b, Ste86, Sto82, Tan89, Tho82, Tur87a, Vin89, Wat83, Wei88, WDS84, WM85, Woo87]. Reviews [Ber87b]. Revisited [Nov84]. Reweighted [Ano89e, SCR88, Wil81]. RFACTOR [Hin80]. Ridge [Ano86e, DH87, Hoe85, Swi81]. Ridge-Regression [DH87]. Right [Man81]. Risk [ABS81, CA85, Fer88, FSL82, Joh82b, Mar82, Par82, PS83, Rai82, Str82b, Swa82, Wei82a, Nai87]. Robert [PS82]. robin [DK87]. Robust [Ano89e, Cle81, Gam81, SCR88, SI82]. Robustness [Che82, CW88]. Role [BPT82, Boa85, CSM86, God85, Hah85c, Joi85a, Joi85b, Neu85, Noe80, Var86, Wac85, Ber87c, JH83]. Room [Mil87, Pri87, Ros87]. Roots [DBM86]. Rotation [HPB82]. Rough [Cha84]. round [DK87]. round-robin [DK87]. RRCOV [SD86]. Rubin [Hin80]. Rule [Rob83b, Tar81]. Rules [Cha84, Poo80, RR84, SS89]. ruses [VY85].

S [Dav81, Zel85, KLV89]. Salvo [Tuk86]. same [FK86]. Sample [Ano87p, Ban88, Bol85, Dal86a, Ena88, Fle82, Gol89, Gue81, HA86, KH89, Len87, MJ84, PG83, Pen81, RT81, Roy86, Thi87, Tho87, WT80, Ber81, CGL83, ME84, Shu82]. Sample-Size [Dal86a, Ena88, Len87, Thi87]. Samples [Che82, Fox89, IS85, Tam84]. Sampling [AJ88, Bel87, Bol85, FM84, Gle86a, RR85, SC81, Aso80, SR83, WC87]. Samuelson [Far89, WS80]. SAS [KGHS84, LM86, NZ88, Sal88, SA87].

SAS/IML [NZ88, Sal88]. SAS(R) [Ami87]. Savings [Lev87c]. SCA [Ass86]. Scale [Cle84a, Ros81a]. Scales [Ant80]. Scaling [YTL80]. Scatterplots [Cle81]. School [Bor87, Hac87, Hof84, Pos82]. Schools [DSAGR87, RM88]. Science [Hac87, Hun83a, Lam83, Mar83b, Rai82, RM88, Mar83c, Yel83a]. Sciences [AC82, Lun88, Moo81]. Scientific [Cle84b, Die83b, KM89b, WBH^{+}88]. Scientists [BS85]. Scope [Mos88]. Score [Ano85i, BL85, RR85]. Scores [Man81, Shi88b]. Scoring [Wai83]. Screening [Fre83b, FP89, GW81]. SCSS [NW80]. Sealy [Bol84]. Search [SPC87]. Searching [Wai83]. Searle [PS89a]. Seasonal [Che88, Pie80]. Secant [MN88]. Second [Sea88a, SZM87]. Second-Degree [SZM87]. Section [BHH^{+}80, Box80a, Bra82, EWW^{+}82, Gen86, Min80, Mos82, Par80, Sne82, Wat80, Zah82]. Sectional [Die83a]. Seemingly [BN88, SK85]. Segmented [Sha86b]. SELECT [HG83]. Selected [FS89, Moo81]. Selecting [FC87, Gus86]. Selection [Edw84, Fox89, HG83, HH87, Pei87, VY85]. SELECTOR [Gus86]. Self [Boe82, Hid86]. self-representing [Hid86]. Self-Supporting [Boe82]. Semi [MP88]. Semi-Markov [MP88]. Semilogarithmic [Sch86]. Sensitivity [RR84]. Separating [Ano85i, BL85, Mar83b, Mar83c, Yel83a]. Separation [Rai82]. serial [Sti86]. Series [Bak84, DBM86, Die83a, DL82, Gam81, Hen83, Int85, Lev87a, Rei84, Rei87]. Service [BS89, BK86]. Set [GW80, Har86a, HL87, Pos83, RBL86]. Set-Theoretic [RBL86]. Sets [SF80, Nai87]. Several [IMS87, YTL80]. Sex [McC80]. Shading [Dun88b]. Shaffer [Lev80]. Shall [Yel83b]. Shaw [TG89]. Shayle [PS89a]. SHAZAM [WM87]. Short [Mar83a, Wat83, Ped85]. Should [Ame85, Ehr84, Gup82, Hig82, Joh82a, Kna82, Leo82, Owe82, Sei89, Shv84, Swi82, Tan82a, Tan82b, TVV85]. SIBYL [Mak81]. SIBYL/TFS [Mak81]. Sided [Kap87]. Sidelight [DH87]. Sidney [DS81]. Signal [Kra88, Ste89b]. Signal-Detection [Kra88]. Significance [Roy86]. SIGSTAT [Gal84]. SIMBOXJ [Rei87]. SIMCA [Gre86a]. Simon [Ham87a]. Simple [Gre86a, JD89, LSWB89, RT81, RR83, Rea85, SF80, Thi87, Thu89, Wil82, Aso80, FF80, GP89, Nai87, PG82]. Simpler [Hef88, Sha85]. SIMPLEX [Dem86]. SIMPLEX-V [Dem86]. Simplification [Fle82]. Simplified [SM83]. Simpson [Pai85, Wag82]. Simulating [ES86]. Simulation [FC87, Gle86a, HA84]. Simulations [Gle88, IS85]. SIMULATOR [Rei84]. Simultaneous [BB89, Car86, MRP^{+}89, Sch81a, Sto81, Ang89, Ang90, Wom85]. Sine [Bur86]. Single [BR84, Rei80, Wil81]. Single-Equation [Rei80]. Singpurwalla [HA89]. Singular [Nel85, EW85, Man82]. Singular-Value [Nel85, EW85]. Size [AG88, Dal86a, Ena88, Fox89, Gol89, Gue81, KH89, Len87, MCJ87, Roy86, Thi87, Tho87, Dal86a]. Size-Biased [Fox89]. Sizes [Ban88, Pen81, WT80]. Skip [Hod87]. Small [Che82, Hod84a, Ber81]. small-sample [Ber81]. Smart [Rei84, Rei87]. SmartForecasts [Wil86]. Smirnov [PG83]. Smoothing [Cle81, Sha86a, MS87]. Social [AC82, Moo81]. Software [Ano80f, BS85, Ber87a, Ber87b, Boo84, DP85, Dus86, Edm85, Fer88, Fox83, Gen85, Hah85a, Hah85b, Har86c, Hod87, Mas89, Rei84, Rei87, Sta86, Sha84, Tro86, Tuk85c, Vel85, Whe85, Tuk85a]. Soiling [Rue85]. Solving [GW80, EW85]. Some [Ans85, Ber81, BE84, BS87, Coh88, Col81, DCB88, FTL84, FHI89, Gle86b, Jer88, Joi81, Kor84, Len83, Lit88, Pit80, SSH81, Sea89a, BL87, CJT87, VY85]. Sometimes [Ham87b, Sch84c]. SORITEC [Bak84]. Sources [VR83]. SP [Nor87]. Space [KS88, EG87]. Spanish [CS86]. Spatial [Mas89]. SPCTL [Lin86a]. Speaking [Bro80b]. Specification [Arm81, Fra81b, Hei81b, NW81, Pay81, Hei81a]. Spectral [MM84]. SPIDA [Geb89]. Split [Ela84, Gil84, MBM84a, MBM84b]. Spread [Hef88]. SPSS [Sau84a, Sau84b]. SPSS/PC [Sau84a]. SPSS/Pro [Sau84b]. SQC [Var86]. Square [Coh88, Hjo88, Mur86]. Squared [Gil81, Han83, HW86, PT80, Pos89, WB81]. Squares [Ano89e, BM88c, Ehr83a, EL88, Har86c, HO85, Hun89, Kem89b, PS89b, Sea89b, SCR88, WS88, Wil81, Woo80, BvE88, MA84, PS89a, SSM80, SP85a, Wom85]. SSF [Sch84b]. SSI [Sch84b]. SSI/SSF [Sch84b]. Stage [BB84, CBI80, WC87]. Stagewise [All87]. Standard [Ano87i, Fle82, FR86, Nor89, Sha85, Sti82, Spi86, Spi87, Tho82]. Standardization [Lit82]. Starting [RM88]. STAT [Fir82, Gus86, Per84, PH86, Wai80, Sug86]. STAT80 [Ham85]. STATCALC [LMM84]. State [Bon88, Dun88a, Leh88, PS83, Rey88]. State-of-the-Art [PS83]. Statements [Ano85e, AS84]. STATGRAPHICS [KJ89, Pol82]. Statist [Ang90, CH86, GY90, Kar86, Kar87, SYC88, Spi87, Zel85]. Statistic [Ano87g, Cox82, DW86, Gil81, Haw89, Mee87, PG83, Som86, Har80]. Statistical [Ad 83, AG88, Ano80f, Ano89f, Ass86, BFM^{+}88, Bar83a, BR84, Ber87a, Ber87b, BF81, BHH^{+}80, Boa82b, Boa83, Boe82, BCF^{+}81, Bon81, Bon83a, Bon83b, Boo84, Box80a, Bra82, Bro83b, Cha80, CBI80, CM83, Con87, Cor87, Dal86b, Dal88c, Dal89b, DeM83, Die83a, Die83b, Dun81, Dun87, Dun88b, Dus89, EWW^{+}82, Ell83, Ell88, ED84, Fie83, Fre83a, FGHK83, Geb81, Geb89, Geh83, Gen85, GF80, Gre83b, Gue87, Gup82, Hah85a, Hah85b, Ham85, HN86, Hig82, Hin82, Hin88, Hog85, Hu88, IHD^{+}83, Joh82a, Kem83, Kis83, Kna82, KGHS84, LMM84, Leh88, Leo82, Lit88, Lun80, Lun84, LH81, MD85, Mar81a, Mar81b, Mar83d, Mas89, MBM^{+}85, Min80, Mos81, Mos82, Mos83, Mos88, Mul80, Nef83, Neu85, NP82, Orr80, Owe82, Par80, Pol82, Pos81, Rem83, Ric83, Rob83a, Roh84]. Statistical [Ros81b, RBL86, Rue85, Sta86, Sug86, Sal85b, Sch84b, SBA85, Sea89a, Sla83, Sne82, Sol83, SP84, SP85b, SC86, Sto80, Str80, SD89, Swi82, Tae81, Tan82a, Tan82b, TW86, Tia83, Tro86, Tru81, Tuk85c, Van80, VR83, Vel85, Wai80, Wal83, Wat80, Zah82, Hel88, Tuk85a, Bon88, Dun88a, Rey88]. Statistician [BPT82]. Statisticians [BHH^{+}80, Boa85, Box80a, Bra82, EWW^{+}82, God85, Hah85c, Joi85a, Joi85b, Min80, Mos82, Par80, Sne82, Sne84, Wac85, Wat80, Zah82]. Statistics [Ano87f, BM86, BMM83, Bas81, BS89, BZ89, Bra83a, Bra83b, Bro80a, BK86, Bro83a, BE84, Bry84, Bur86, Car83, CSM86, CA87, Cho86, CI81a, CI81b, Die89, Dus86, EJKS87, FP81, Fie84, Fli81, Flu89, Gal84, GL83, Hac87, Hah89, Ham83a, Ham83b, HPM89, Haw89, Hil83, Hof84, Hoo80, Hud87, Hun81, Hun83a, Hun83b, Ive85, Joh83, Kem80, KB84a, KS88, Leo83, Lev85, Low82, Loy83, MD85, Mar83a, McD83, Mil87, MSGG87, MO84, MR89, Mos88, Naj83, Noe80, Noe81, Ott83, Pri87, RR83, Ros87, RM88, Sal85a, Sal85b, SW86, SJ86, Shu83, Ste85, Sto88, VD84, WMB82, WDS84, WB81, Zar80, FK86, FF80, GY89, GY90, KD87, Len83, Sch81b, MGS82, GW80, Pos83]. Statistics-Aided [Hah89]. STATISTIX [HN86, Ber87b]. STATOOLS [Dal86b]. STATPAC [Str80]. Status [Sto81]. Stein [KM88, Whi89]. Step [Rog84]. Stepwise [Tur87b]. Stevens [Bol86]. Stirling [Ano89f, Hu88]. Stochastic [Fer88, RM81]. Stochasticity [MCJ87]. Stopping [RR84]. STPLAN [BH81]. Straight [MSTY81]. Straightening [Sha86b]. Strategic [BFM^{+}88]. Stratospheric [Atk82, Sto82, Wat82]. stratum [Hid86]. strong [Tom84]. Stronger [Cov89]. Structure [Sch88a, SZM87, TH81, VPO^{+}85]. Structures [Arb89]. Student [Ber81]. Students [Bro80a, Pos82]. Studies [CA85, FS87, FM89, GD81, HA84, KGHS84, NP82, War85, MS87]. Study [Bas81, BH83, BH81, FC87, SC86, SC86]. Stupidities [Kem89a]. subclass [SSH81]. Subject [Man81]. Subnational [Ban88]. Subroutine [Chu88, GG80b]. Subset [FC87, Pin88]. Successes [Lou81]. Sufficiency [Mee86]. Sufficient [Lev85, Mee87, MA84]. Suitable [FHI^{+}88]. Sukhatme [Ful80]. Sum [CS83, Riv81]. summary [KD87]. Summated [Ant80]. Sums [Flu86, LW82, SP85a]. Sunset [Tuk86]. Supplementary [Ste89a, Ste89b]. Supplements [Ste88a]. Support [MO84, SJ86, GD89]. Supporting [Boe82]. Supreme [Tho82]. SUREX [Swa84]. SUREXPL [Lin80]. Surface [CCW^{+}86, HO85, KC84, Lin80]. Surprising [Sch89a, NW86, VY85]. Survey [Coh82, CBJ86, Die83a, Fox83, HA84, Lit88, Nor87, Sha84, Hid86, Pie80]. Surveys [BMP^{+}85]. Survival [GH88, Gre86b, JM85, Lin86a, PC83, Ste89a, Swa84]. SURVREG [PC83]. SURVTAB [Fox83]. Symbolization [TH81]. Symmetric [BE84, CS83, HPB82]. Synthesis [HO83]. SYSTAT [Dal87b, Ste88a, Ste89a, Ste89b, Wil87]. System [Ass86, BS80, BCF^{+}81, Bon81, DKY80, Dun81, KG80a, Lev87a, Mak81, Mar81b, Mos81, NW80, Pol82, Rel85, Ros81b, Sau84a, Tae81, Tro86, Wai80, Web87, BCF^{+}81]. Systematic [Bas81]. Systems [Gen85, Hah85a, Hah85b, Lam80, Tuk85c, Vel85, Tuk85a].

T [FR85, Gue81, NP82, Par82]. T/Dioxin [Nea82]. Table [Cox82, Pai85, RM81, Sto80, Gol86]. Tables [Dal87a, Dav86, Del86, Gil81, Hod84a, Hod84b, Sha82, Sto80]. Tabulation [Fox83, Wai80]. Tail [Man83]. Tally [Hsi81]. Task [Hog85, Wat81]. Task-Analysis [Wat81]. Taxonomy [Roh87]. TCDD [NP82]. TDDA [Dus86]. Teacher [AG88]. Teaching [AG88, Ano87f, BAP89, Bor87, Bro80a, Die89, Hof84, Kem80, Lit82, MD85, MSGG87, Pos81, SW86, Tor86, AT86]. Technical [Ano85h, Ehr82, Mar84, MFG^{+}84]. technique [SI82].

Techniques [Dal89b, GSW82, Rah81, SPC87]. Technology [Mar83b, Mar83c, Yel83a]. Television [HK84]. Tell [Bro80b]. Template [Dal88a]. Term [Haw80]. Test [Ano87g, BB89, Bur82, Che82, Cox82, DW86, FS87, Har86a, HL87, Ima82, Job82, Kin89, Man83, Man87, PS81, Rob83b, Ste87, Wai83, WB81, AO89, Dav80, Har80]. TESTAT [Ste87]. Testing [DCB88, DR83, DK88, Hod84b, Jer88, Job82, JSL89, Lit89, Loy83, Mag82, Swa84, Tro86, MS87, Nai87]. TESTRAND [DR83, DK88]. Tests [Ami87, BR84, Bor82a, Bor82b, Dat82, Dav86, DBM86, FR85, Gue81, Han89, KMM87, Lac84, PT80, Roy86, Sch81a, SM83, Wal82, GS86, Hel88, SS84]. Texas [Hac87]. Textbooks [BA87]. Textile [BS83]. TFS [Mak81]. Theil [Did85]. Their [Dri80, FR85, Rai82]. Theme [Bry84]. Theodore [FK87]. Theorem [AL86, Ber88, DG86, Hil88, Jay88, KK84, Kul88, NA85, Pau81, RD88, Cox84, PG82, Tar81, Yan88, Zel88b, Zel88a]. Theorems [Ans85]. Theoretic [RBL86, TC88]. Theory [Gue81, MP88, Tru81]. There [Hin80]. Thirteen [RN88]. Three [BR84, DHPR88, FF80, NP82, Pai85, SC81]. Three-Dimensional [DHPR88]. Three-Generation [NP82]. Three-Way [Pai85]. Thresholds [Rue85]. Thumb [SS89]. Time [Bak84, BS89, DBM86, Die83a, DL82, Gre86b, Hen83, Hun81, Int85, Lev87a, LK86, Rei84, Rei87, SB85]. Time-Varying [LK86]. TM [Pol82, Smi84]. TOBIT [Ste88a]. Today [Bon83a, Bon83b, DeM83, Fie83, Sla83, Wal83]. tolerance [RC88]. Tool [HE84, KF89, Lit82, MD85, Whe85, EW85]. Toolbox [Sug86]. Total [Mak81]. Toxicity [NP82]. TPATH [AC82]. Track [Daw89, Dal89d]. Tracking [Dal89d]. Trademark [KC84, KGHS84, Bak84]. Training [Bas81, BHH^{+}80, Box80a, Bra82, EWW^{+}82, Min80, Mos82, Par80, Sne82, Wat80, Zah82]. Transformation [Boa85, God85, Hah85c, HW86, HH87, Joi85a, Joi85b, Mil84, Wac85]. Transformations [CI81a, CI81b, Fli81, Lie80, Noe81, Sha86b, RC88]. Transforming [LV89]. Transparency [Pos83]. Trap [DS89]. Treatment [Har86c, SM83]. Trees [Fai86]. Trends [BMP^{+}85, Cha80, Hod84b]. Trial [Sei89]. Trials [Ano85f, Ber84, Gei84a, Zel84, Ber87c, Gei84b, NW86, Zel85]. Triangularization [SBA85]. Tropical [Neu85, Pik85]. truncated [Mul88]. Trusted [BA87]. Truth [Bro80b]. TSA [Hen83]. TSP [Hal88]. Tumor [Ami87]. Tumor-Prevalence [Ami87]. TURBO RAND [Gle86a]. Tutorial [AC82]. Two [Ami87, Ano87i, BB84, DCB88, DHPR88, Dun89, FR85, Gil81, HA86, Lit89, LS89, Poo80, Ros81a, SS89, Spi86, Spi87, WF80, BvE88, FK86, Gue87, WC87]. Two- [DHPR88]. Two-Group [FR85]. Two-Scale [Ros81a]. Two-Stage [BB84, WC87]. Two-Variable [Dun89, WF80]. Two-Way [Gil81]. Type [TT80, Pav88].

U [Nor87]. U-SP [Nor87]. U.S.S.R. [KS88]. ubiquitous [JH83]. UMAP [Fol82]. Umbrella [KM87]. Unbalanced [GG80a, Lit82, SM83, Her86]. Unbiased [Ano86g, Ano87h, BM88b, HTM89, Kem89b, MN87, PS89b, Sea89b, BvE88, GD89, Kar85, Kar86, Kar87, Kot86, MA84, PS89a, SS84, VM89]. Uncertainty [AL86, Coh86, Sto88]. Unclassed [GH82]. Unconditional [Shu88, Thu89]. Uncorrelated [RNT84]. Undercount [Ano85h, MFG^{+}84]. Understanding [Flu89, Man87, MS83, Mul82, WF80]. Undue [Bis84, Bro84, PHA84]. unequal [SSH81]. unequal-subclass-numbers [SSH81]. Ungrouped [FP87]. Uniform [LW82, EG87, GM86]. Uniformly [GD89, SS84]. Union [Jac84]. Unique [MN87]. Unisex [Coh86]. Unit [DBM86]. Units [CSM86, Aso80, Hid86]. Univariate [BW84, Edg84, FR86, Job82, Lee86, Mer89, Whi86, TC88]. Universities [GF80]. University [Boe82, Bor87, Hac87, Sne84]. UNIX [Per84, PH86]. Unrelated [BN88, SK85]. Unsurprising [Sch89a]. untitled [Bar86, Bel80, Ber89, Bes87, Bur87, BG80, CJ81, CT89, Chr81, Con89, Coo87, Dav85b, Dav89, Dix88, Ell87, Gil80, Gol88, Hal89, HB86, Hil85, KM89a, KP87, KG80b, KSB^{+}81, Koe87, Koe86, Lac85, LeP87, Lev87b, Lig85, Lig88, Lin85b, Lin87, Lin88, Man88, Mar85, McB80, Mil89, Mor87a, Mor87b, Nas86a, Nas86b, O’B88, Par89, Pla86, Pos80, PF89, Reg86, Sch85a, Shr87, Shu86, Ste88b, Ste86, Tan89, Tur87a, Vin89, Wei88, WM85, Woo87]. Update [LS89]. Urn [Pau81]. Use [Alb89, Ano85g, Bol85, Dav80, DP87, GH84, Hoo80, Ima82, LAH89, LG85, Man82, MD83, MGS85, Pen81, Pik85, Tia83, WS88, Her80, Rav85, RC88]. Used [Ano84g, Don82, Gon80]. Useful [BE84, KK84]. User [MP80, MEP^{+}86, MD81, MM84, Nor87, Wal87]. User-Friendly [Nor87, Wal87]. User-Oriented [MP80, MEP^{+}86, MD81, MM84]. Uses [Dri80]. Using [Ano86f, Ano87f, Ano87i, CA85, DS89, GSW82, Gus86, Har86c, Haw89, Job82, Lin81, Mag82, Mee87, Nor84, RJ82, RR85, SW86, Shi88a, Smi84, SA87, Whi89, CH85, CH86, KD87, SP85a, SR83, Spi86, Spi87]. usual [Bly80]. Utility [Dal86b, KF89]. Utilized [BZ89].

V [Ful80, Dem86]. V1.2 [Dus86]. validation [EG83]. Validity [Hil87]. Value [KM87, Kin89, Mar83b, Mar83c, Nel85, Yel83a, EW85, Man82, MC87]. Valued [BZ89]. Values [Ano84g, Don82, Kor84, Lac84, Har80]. Variable [Ano86h, DM87, Dun87, Dun89, Gre83a, Pei87, Sei89, WF80]. Variable-Width [Dun87]. Variables [CS83, FC87, Flu86, Flu89, HG83, Ham87b, HW86, Joi81, LK86, LW82, Roc88, RNT84, Sav82, Whi89, BG89, Ham84, PK82]. Variance [Ano86g, Ano87h, Ano88e, BW84, Cam88, Car81a, Car81b, CBJ86, CC89, GG80a, LSWB89, Sea81, Ste82, TH81, CGL83, GD89, Kar85, Kar86, Kar87, Mul88, SYM84, SYC88, Sch85b, SSH81, Shu82, VM89]. Variances [BB89, Ros81a, NW86]. Variates [BB84, ES86, HPB82, Wil82]. Variations [Bry84]. Various [Bol85]. Varying [LK86]. VAX [DK88]. VAX-11 [DK88]. Version [Ber87b, Ell89, Fra81a, Geb89, Int85, KLV89, KJ89, KF89, WM87, Wil87]. Versus [PN87, Kem80, Nai87]. Via [Ano89e, SCR88, Ans85, Gol89, Pau85, SPC87]. VIC [Swa84]. VIC-20 [Swa84]. Video [GW80, KC84, Koe81]. Videotape [Mar83a, Nel81, Wat83, WDS84]. View [KS88]. Violated [Gam81]. Violations [Ano88g, SD87]. Visual [BAP89, SP84, SP85b, Sti86]. Volume [Ano80l, Ano81i, Ano82k, Ano84o, Ano86n, Ano87o, Ano88n, Ano89n]. Volumes [Moo81].

Wald [Man87]. Walks [BH86]. Wallbanger [SPC87]. Washington [Bro80b]. Wasserman [Dri86]. Way [Ano88g, Cam88, Cox82, Gil81, Pai85, Sto81, SD87, Sch85b]. Ways [RN88]. Weighted [BM88c, Cle81, WS88, SP86]. Whatever [Sch86]. while [FM84]. White [CS86]. Who [Yel83b]. Whole [Bro80b]. Whose [Hun81]. Width [Dun87]. WIENER [Chu88]. Wilcoxon [Man81]. William [Bol84, DM81, Rya82, TG89]. Winter [FC81]. without [Haw80]. Witness [Bar83b]. Wolfowitz [Wei82b]. Words [Sea89a]. Workers [BS83]. Worksite [Lev87c]. Workspace [Sav82]. World [Hun81, Lit88]. Worse [Kem89a]. Writing [Ehr82]. WUBA [Dal87b]. WUBA-WUBA [Dal87b].

X [Ano87p]. X-BAR [Ano87p]. XTREE [TW86].

Year [DH85, Pos81]. Years [Ano86e, Hoe85, MGS82, Her86, Noe84]. yield [DBS^{+}85]. YREG [Yak88].

Zellner [Gei84b]. Zero [Bro86, CC89, Job82]. ZIP [Smi84]. Zonal [Nag81].

References

Aitchison:1981:BRR

[ABS81] John Aitchison and John Bacon-Shone. Bayesian risk ratio analysis. The American Statistician, 35(4):254–257, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

           URL http://www.jstor.org/stable/2683302.

Anderton:1982:TTP

[AC82] Douglas L. Anderton and Joseph Conaty. TPATH — a tutorial path analysis program for the social sciences. The American Statistician, 36(1):61, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

           org/stable/2684099.

AHCPE:1983:EGS

[Ad 83] The Ad Hoc Committee on Professional Ethics. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Reply. The American Statistician, 37(1):19–20, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685835.

Ashikaga:1981:LE

[ADS81] Taka Ashikaga, W. Edwards Deming, and Alan Stuart. Letters to the Editor. The American Statistician, 35(2):115, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683162.

Atkins:1984:BBF

[AG84] John Atkins and Erdogan Gunel. BAYCAT: Bayes factors for categorical data analysis. The American Statistician, 38(2):156, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683259.

Anderson:1988:SEC

[AG88] Larry R. Anderson and David R. Guichard. Statistical effects of class size and teaching ability in determining teacher ratings. The American Statistician, 42(1):55–59, February 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685262.

Aickin:1985:PR

[Aic85] Mikel Aickin. Poisson regression. The American Statistician, 39(2): 145, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682825.

Agrawal:1988:PED

[AJ88] M. C. Agrawal and Nirmal Jain. Predictive estimation in double sampling procedures. The American Statistician, 42(3):184–186, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684997.

Alaouze:1986:GGT

[AL86] Chris M. Alaouze and Peter J. Lloyd. A generalization of Gurland’s theorem, with applications to economic behavior under uncertainty. The American Statistician, 40(1):70–71, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683136.

Albert:1989:NPU

[Alb89] James H. Albert. Nuisance parameters and the use of exploratory graphical methods in a Bayesian analysis. The American Statistician, 43(4):191–196, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685360.

Alley:1987:NSR

[All87] William M. Alley. A note on stagewise regression. The American Statistician, 41(2):132–134, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684225.

Amemiya:1985:WSD

[Ame85] Yasuo Amemiya. What should be done when an estimated between-group covariance matrix is not nonnegative definite? The American Statistician, 39(2):112–117, May 1985. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682808.

Amini:1987:TMS

[Ami87] Saeid B. Amini. Two mainframe SAS(R) programs for nonparametric tests for homogeneity of dose-response growth curves and regression analysis of tumor-prevalence data. The American Statistician, 41(3):237, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685116.

Angers:1989:NQS

[Ang89] Claude Angers. Note on quick simultaneous confidence intervals for multinomial proportions. The American Statistician, 43(2):91, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684505. See

correction [Ang90].

Angers:1990:CNQ

[Ang90] Claude Angers. Correction: “Note on quick simultaneous confidence intervals for multinomial proportions” [Amer. Statist. 43 (1989), no. 2, 91]. The American Statistician, 44(1):65, ???? 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Ang89].

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. The American Statistician, 34(1):??, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683008.

Anonymous:1980:BMb

[Ano80b] Anonymous. Back matter. The American Statistician, 34(2):??, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684127.

Anonymous:1980:BMc

[Ano80c] Anonymous. Back matter. The American Statistician, 34(3):??, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683893.

Anonymous:1980:BMd

[Ano80d] Anonymous. Back matter. The American Statistician, 34(4):??, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684080.

Anonymous:1980:BDF

[Ano80e] Anonymous. Business directories: Findings and recommendations of the ASA Committee on Privacy and Confidentiality. The American Statistician, 34(1):8–10, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682987.

Anonymous:1980:CRS

[Ano80f] Anonymous. A comparative review of statistical software. The American Statistician, 34(4):248, November 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684077.

Anonymous:1980:EC

[Ano80g] Anonymous. Editorial collaborators. The American Statistician, 34(4):256–257, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684079.

Anonymous:1980:FMa

[Ano80h] Anonymous. Front matter. The American Statistician, 34(1):??, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682985.

Anonymous:1980:FMb

[Ano80i] Anonymous. Front matter. The American Statistician, 34(2):??, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684105.

Anonymous:1980:FMc

[Ano80j] Anonymous. Front matter. The American Statistician, 34(3):??, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683867.

Anonymous:1980:FMd

[Ano80k] Anonymous. Front matter. The American Statistician, 34(4):??, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684059.

Anonymous:1980:VI

[Ano80l] Anonymous. Volume information. The American Statistician, 34(4):255–256, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684058.

Anonymous:1981:BMa

[Ano81a] Anonymous. Back matter. The American Statistician, 35(1):59–60, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683595.

Anonymous:1981:BMb

[Ano81b] Anonymous. Back matter. The American Statistician, 35(2):116, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683163.

Anonymous:1981:BMc

[Ano81c] Anonymous. Back matter. The American Statistician, 35(3):??, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684002.

Anonymous:1981:BMd

[Ano81d] Anonymous. Back matter. The American Statistician, 35(4):??, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683311.

Anonymous:1981:FMa

[Ano81e] Anonymous. Front matter. The American Statistician, 35(1):??, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683574.

Anonymous:1981:FMb

[Ano81f] Anonymous. Front matter. The American Statistician, 35(2):??, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683141.

Anonymous:1981:FMc

[Ano81g] Anonymous. Front matter. The American Statistician, 35(3):??, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683972.

Anonymous:1981:FMd

[Ano81h] Anonymous. Front matter. The American Statistician, 35(4):267, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683284.

Anonymous:1981:VI

[Ano81i] Anonymous. Volume information. The American Statistician, 35(4):271–274, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683283.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. The American Statistician, 36(1):??, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684104.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. The American Statistician, 36(2):143, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684037.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. The American Statistician, 36(3):??, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683852.

Anonymous:1982:BMd

[Ano82d] Anonymous. Back matter. The American Statistician, 36(4):401, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683103.

Anonymous:1982:CIM

[Ano82e] Anonymous. Corrigendum: An improved method of estimating in integer-parameter by maximum likelihood. The American Statistician, 36(1):65, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684103.

Anonymous:1982:EC

[Ano82f] Anonymous. Editorial collaborators. The American Statistician, 36(4):399–400, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683102.

Anonymous:1982:FMa

[Ano82g] Anonymous. Front matter. The American Statistician, 36(1):??, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684081.

Anonymous:1982:FMb

[Ano82h] Anonymous. Front matter. The American Statistician, 36(2):121, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684003.

Anonymous:1982:FMc

[Ano82i] Anonymous. Front matter. The American Statistician, 36(3):??, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683830.

Anonymous:1982:FMd

[Ano82j] Anonymous. Front matter. The American Statistician, 36(4):??, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683077.

Anonymous:1982:VI

[Ano82k] Anonymous. Volume information. The American Statistician, 36(4):397–399, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683076.

Anonymous:1983:BMa

[Ano83a] Anonymous. Back matter. The American Statistician, 37(1):??, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685854.

Anonymous:1983:BMb

[Ano83b] Anonymous. Back matter. The American Statistician, 37(2):??, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685892.

Anonymous:1983:BMc

[Ano83c] Anonymous. Back matter. The American Statistician, 37(3):??, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683389.

Anonymous:1983:BMd

[Ano83d] Anonymous. Back matter. The American Statistician, 37(4):??, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683516.

Anonymous:1983:FMa

[Ano83e] Anonymous. Front matter. The American Statistician, 37(1):??, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685818.

Anonymous:1983:FMb

[Ano83f] Anonymous. Front matter. The American Statistician, 37(2):??, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685867.

Anonymous:1983:FMc

[Ano83g] Anonymous. Front matter. The American Statistician, 37(3):??, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683364.

Anonymous:1983:FMd

[Ano83h] Anonymous. Front matter. The American Statistician, 37(4):??, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683492.

Anonymous:1984:BMa

[Ano84a] Anonymous. Back matter. The American Statistician, 38(1):??, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683573.

Anonymous:1984:BMb

[Ano84b] Anonymous. Back matter. The American Statistician, 38(2):??, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683266.

Anonymous:1984:BMc

[Ano84c] Anonymous. Back matter. The American Statistician, 38(3):??, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683671.

Anonymous:1984:BMd

[Ano84d] Anonymous. Back matter. The American Statistician, 38(4):??, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683431.

Anonymous:1984:CLE

[Ano84e] Anonymous. Corridenda: Letters to the Editor. The American Statistician, 38(1):72, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683572.

Anonymous:1984:CNC

[Ano84f] Anonymous. Corrigenda: a note on confidence bands for the logistic response curve. The American Statistician, 38(1):72, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683570.

Anonymous:1984:CRE

[Ano84g] Anonymous. Corrigenda: The relative effectiveness of procedures commonly used in multiple regression for dealing with missing values. The American Statistician, 38(1):72, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683571.

Anonymous:1984:CMM

[Ano84h] Anonymous. Corrigendum: In Memoriam: Maurice George Kendall, 1907–1983. The American Statistician, 38(3):234, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683670.

Anonymous:1984:EC

[Ano84i] Anonymous. Editorial collaborators. The American Statistician, 38(4):333–334, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683430.

Anonymous:1984:EN

[Ano84j] Anonymous. Editor’s note. The American Statistician, 38(3):229, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683668.

Anonymous:1984:FMa

[Ano84k] Anonymous. Front matter. The American Statistician, 38(1):??, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683548.

Anonymous:1984:FMb

[Ano84l] Anonymous. Front matter. The American Statistician, 38(2):??, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683235.

Anonymous:1984:FMc

[Ano84m] Anonymous. Front matter. The American Statistician, 38(3):??, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683645.

Anonymous:1984:FMd

[Ano84n] Anonymous. Front matter. The American Statistician, 38(4):??, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683391.

Anonymous:1984:VI

[Ano84o] Anonymous. Volume information. The American Statistician, 38(4):335–337, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683390.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. The American Statistician, 39(1):??, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683921.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. The American Statistician, 39(2):??, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682834.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. The American Statistician, 39(3):??, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683952.

Anonymous:1985:BMd

[Ano85d] Anonymous. Back matter. The American Statistician, 39(4):??, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683029.

Anonymous:1985:CIC

[Ano85e] Anonymous. Correction: Improved confidence statements for the binomial parameter. The American Statistician, 39(1):81, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683920.

Anonymous:1985:CPD

[Ano85f] Anonymous. Corrections: On prior distributions for binary trials. The American Statistician, 39(2):158, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682832.

Anonymous:1985:CUI

[Ano85g] Anonymous. Corrections: On the use of the inverse of the correlation matrix in multivariate data analysis. The American Statistician, 39(3):241, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683951.

Anonymous:1985:CRA

[Ano85h] Anonymous. Corrections: Report of the ASA technical panel on the census undercount. The American Statistician, 39(3):241, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683950.

Anonymous:1985:CSB

[Ano85i] Anonymous. Corrections: Separating the Brier score into calibration and refinement components: a graphical analysis. The American Statistician, 39(2):158, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682833.

Anonymous:1985:FMa

[Ano85j] Anonymous. Front matter. The American Statistician, 39(1):??, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683894.

Anonymous:1985:FMb

[Ano85k] Anonymous. Front matter. The American Statistician, 39(2):??, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682800.

Anonymous:1985:FMc

[Ano85l] Anonymous. Front matter. The American Statistician, 39(3):??, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683922.

Anonymous:1985:FMd

[Ano85m] Anonymous. Front matter. The American Statistician, 39(4):??, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683009.

Anonymous:1986:BMa

[Ano86a] Anonymous. Back matter. The American Statistician, 40(1):80–87, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683140.

Anonymous:1986:BMb

[Ano86b] Anonymous. Back matter. The American Statistician, 40(2):??, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684888.

Anonymous:1986:BMc

[Ano86c] Anonymous. Back matter. The American Statistician, 40(3):??, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684565.

Anonymous:1986:BMd

[Ano86d] Anonymous. Back matter. The American Statistician, 40(4):??, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684623.

Anonymous:1986:CRA

[Ano86e] Anonymous. Correction: Ridge analysis 25 years later. The American Statistician, 40(1):79, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683139.

Anonymous:1986:CBP

[Ano86f] Anonymous. Corrections: Bayesian point estimation using the predictive distribution. The American Statistician, 40(3):258, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684563.

Anonymous:1986:CMV

[Ano86g] Anonymous. Corrections: On minimum variance unbiased estimators. The American Statistician, 40(3):258, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684564.

Anonymous:1986:CPL

[Ano86h] Anonymous. CRAWTRAN-PC: Limited dependent variable regression program. The American Statistician, 40(1):56, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683129.

Anonymous:1986:EC

[Ano86i] Anonymous. Editorial collaborators. The American Statistician, 40(4):332–333, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684622.

Anonymous:1986:FMa

[Ano86j] Anonymous. Front matter. The American Statistician, 40(1):??, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683104.

Anonymous:1986:FMb

[Ano86k] Anonymous. Front matter. The American Statistician, 40(2):??, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684862.

Anonymous:1986:FMc

[Ano86l] Anonymous. Front matter. The American Statistician, 40(3):??, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684531.

Anonymous:1986:FMd

[Ano86m] Anonymous. Front matter. The American Statistician, 40(4):??, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684594.

Anonymous:1986:VI

[Ano86n] Anonymous. Volume information. The American Statistician, 40(4):334–336, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684593.

Anonymous:1987:BMa

[Ano87a] Anonymous. Back matter. The American Statistician, 41(1):??, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684340.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. The American Statistician, 41(2):??, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684238.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. The American Statistician, 41(3):??, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685126.

Anonymous:1987:BMd

[Ano87d] Anonymous. Back matter. The American Statistician, 41(4):??, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684772.

Anonymous:1987:CRI

[Ano87e] Anonymous. Correction: Relative importance by averaging over orderings. The American Statistician, 41(4):341, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684770.

Anonymous:1987:CMT

[Ano87f] Anonymous. Corrections: a method for teaching statistics using N-dimensional geometry. The American Statistician, 41(3):248, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685124.

Anonymous:1987:CAA

[Ano87g] Anonymous. Corrections: An analytic approximation to the distribution of Lilliefors’s test statistic for normality. The American Statistician, 41(3):248, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685125.

Anonymous:1987:CMV

[Ano87h] Anonymous. Corrections: On minimum variance unbiased estimators. The American Statistician, 41(1):92, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684338.

Anonymous:1987:CTP

[Ano87i] Anonymous. Corrections: Two pitfalls of using standard regression diagnostics when both X and Y have measurement error. The American Statistician, 41(1):92, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684339.

Anonymous:1987:EC

[Ano87j] Anonymous. Editorial collaborators. The American Statistician, 41(4):342–343, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684771.

Anonymous:1987:FMa

[Ano87k] Anonymous. Front matter. The American Statistician, 41(1):??, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684308.

Anonymous:1987:FMb

[Ano87l] Anonymous. Front matter. The American Statistician, 41(2):??, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684217.

Anonymous:1987:FMc

[Ano87m] Anonymous. Front matter. The American Statistician, 41(3):??, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685096.

Anonymous:1987:FMd

[Ano87n] Anonymous. Front matter. The American Statistician, 41(4):??, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684738.

Anonymous:1987:VI

[Ano87o] Anonymous. Volume information. The American Statistician, 41(4):344–346, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684737.

Anonymous:1987:XBD

[Ano87p] Anonymous. X-BAR DIST: Sample mean distributions from discrete populations. The American Statistician, 41(1):78, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684330.

Anonymous:1988:BMa

[Ano88a] Anonymous. Back matter. The American Statistician, 42(1):??, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685276.

Anonymous:1988:BMb

[Ano88b] Anonymous. Back matter. The American Statistician, 42(2):??, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684498.

Anonymous:1988:BMc

[Ano88c] Anonymous. Back matter. The American Statistician, 42(3):??, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685033.

Anonymous:1988:BMd

[Ano88d] Anonymous. Back matter. The American Statistician, 42(4):??, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685154.

Anonymous:1988:CCI

[Ano88e] Anonymous. Corrections: Chebyshev inequality with estimated mean and variance. The American Statistician, 42(2):166, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684497.

Anonymous:1988:CLE

[Ano88f] Anonymous. Corrections: Letter to the Editor. The American Statistician, 42(4):291, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685151.

Anonymous:1988:CEV

[Ano88g] Anonymous. Corrections: The effects of violations of independence assumptions in the one-way ANOVA. The American Statistician, 42(4):292–293, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685152.

Anonymous:1988:CPP

[Ano88h] Anonymous. Corrections: To pool or not to pool: The quality bank case. The American Statistician, 42(2):166, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684496.

Anonymous:1988:EC

[Ano88i] Anonymous. Editorial collaborators. The American Statistician, 42(4):294–296, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685153.

Anonymous:1988:FMa

[Ano88j] Anonymous. Front matter. The American Statistician, 42(1):??, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685251.

Anonymous:1988:FMb

[Ano88k] Anonymous. Front matter. The American Statistician, 42(2):??, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684478.

Anonymous:1988:FMc

[Ano88l] Anonymous. Front matter. The American Statistician, 42(3):??, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684992.

Anonymous:1988:FMd

[Ano88m] Anonymous. Front matter. The American Statistician, 42(4):??, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685128.

Anonymous:1988:VI

[Ano88n] Anonymous. Volume information. The American Statistician, 42(4):297–299, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685127.

Anonymous:1989:BMa

[Ano89a] Anonymous. Back matter. The American Statistician, 43(1):??, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685182.

Anonymous:1989:BMb

[Ano89b] Anonymous. Back matter. The American Statistician, 43(2):??, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684530.

Anonymous:1989:BMc

[Ano89c] Anonymous. Back matter. The American Statistician, 43(3):??, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685074.

Anonymous:1989:BMd

[Ano89d] Anonymous. Back matter. The American Statistician, 43(4):??, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685394.

Anonymous:1989:CNC

[Ano89e] Anonymous. Correction: “A Note on Computing Robust Regression Estimates Via Iteratively Reweighted Least Squares”. The American Statistician, 43(1):69, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685181. See [SCR88].

Anonymous:1989:CSM

[Ano89f] Anonymous. Correction: a statistical method of approach to Stirling’s formula. The American Statistician, 43(3):188, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685073.

Anonymous:1989:CCP

[Ano89g] Anonymous. Correction: Choosing a pretest-posttest analysis. The American Statistician, 43(2):132, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684529.

Anonymous:1989:CLE

[Ano89h] Anonymous. Correction: Linear estimators for the parameter in the Problem of the Nile. The American Statistician, 43(4):292, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685392.

See [JN89].

Anonymous:1989:EC

[Ano89i] Anonymous. Editorial collaborators. The American Statistician, 43(4):293–295, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685393.

Anonymous:1989:FMa

[Ano89j] Anonymous. Front matter. The American Statistician, 43(1):??, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685155.

Anonymous:1989:FMb

[Ano89k] Anonymous. Front matter. The American Statistician, 43(2):??, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684499.

Anonymous:1989:FMc

[Ano89l] Anonymous. Front matter. The American Statistician, 43(3):??, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685057.

Anonymous:1989:FMd

[Ano89m] Anonymous. Front matter. The American Statistician, 43(4):??, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685358.

Anonymous:1989:VI

[Ano89n] Anonymous. Volume information. The American Statistician, 43(4):296–298, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685357.

Ansley:1985:QPS

[Ans85] Craig F. Ansley. Quick proofs of some regression theorems via the QR algorithm. The American Statistician, 39(1):55–59, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683909.

Antonak:1980:FPA

[Ant80] Richard F. Antonak. A FORTRAN program to analyze summated rating scales. The American Statistician, 34(1):59, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683004.

Arizono:1989:TNB

[AO89] Ikuo Arizono and Hiroshi Ohta. A test for normality based on Kullback–Leibler information. The American Statistician, 43(1): 20–22, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685161.

Arbuckle:1989:AAM

[Arb89] James L. Arbuckle. AMOS: Analysis of moment structures. The American Statistician, 43(1):66, February 1989. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685178.

Armor:1981:SED

[Arm81] David J. Armor. [The specification of experimental designs to ANOVA programs]: Comment. The American Statistician, 35(2): 104, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683150.

Angus:1984:ICS

[AS84] John E. Angus and Ray E. Schafer. Improved confidence statements for the binomial parameter. The American Statistician, 38 (3):189–191, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683650.

Abernathy:1986:MBG

[AS86] James R. Abernathy and Pranab K. Sen. In memoriam: Bernard George Greenberg, 1919–1985. The American Statistician, 40 (3):183–184, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684532.

Asok:1980:NCB

[Aso80] C. Asok. A note on the comparison between simple mean and mean based on distinct units in sampling with replacement. The American Statistician, 34(3):158, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683874.

Andrews:1980:GDM

[ASS80] Horace P. Andrews, Ronald D. Snee, and Margaret H. Sarner. Graphical display of means. The American Statistician, 34(4): 195–199, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684060.

Associates:1986:SSS

[Ass86] Scientific Computing Associates. The SCA statistical system. The American Statistician, 40(1):58–59, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683133.

Aguirre-Torres:1986:NTI

[AT86] Victor Aguirre-Torres. A note on teaching infinite moving averages. The American Statistician, 40(1):40–41, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Atkinson:1982:CSO

[Atk82] Edward Stevens Atkinson, Jr. Chlorofluorocarbons and stratospheric ozone: Regulatory background. The American Statistician, 36(3):301–302, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683848.

Binkley:1987:FAE

[BA87] James K. Binkley and Philip C. Abbott. The fixed X assumption in econometrics: Can the textbooks be trusted? The American Statistician, 41(3):206–214, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685106.

Bailar:1984:CCB

[Bai84] Barbara A. Bailar. Comment from the Census Bureau. The American Statistician, 38(4):257–260, November 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683399.

Baker:1984:SET

[Bak84] W. E. Baker. SORITEC[trademark]: Econometric and time series analysis for micros, minis, and mainframes. The American Statistician, 38(4):316, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683417.

Baltagi:1987:PPQ

[Bal87] Badi H. Baltagi. To pool or not to pool: The quality bank case. The American Statistician, 41(2):150–152, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684234.

Bankier:1988:PAD

[Ban88] Michael D. Bankier. Power allocations: Determining sample sizes for subnational areas. The American Statistician, 42(3):174–177, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684995.

Bajgier:1989:VFT

[BAP89] Steve M. Bajgier, Maryanne Atkinson, and Victor R. Prybutok. Visual fits in the teaching of regression concepts. The American Statistician, 43(4):229–234, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685367.

Barabba:1983:EGS

[Bar83a] Vincent P. Barabba. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):18–19, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685834.

Barton:1983:MEE

[Bar83b] Anne Barton. My experience as an expert witness. The American Statistician, 37(4):374–376, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683499.

Barnett:1984:RMP

[Bar84] William A. Barnett. Recent monetary policy and the divisia monetary aggregates. The American Statistician, 38(3):165–172, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683646.

Bartels:1985:IE

[Bar85] Robert Bartels. Identification in econometrics. The American Statistician, 39(2):102–104, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682804.

Barcikowski:1986:BRB

[Bar86] Robert S. Barcikowski. Book review: [untitled]. The American Statistician, 40(3):225–228, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684543.

Baskerville:1981:SSC

[Bas81] J. C. Baskerville. A systematic study of the consulting literature as an integral part of applied training in statistics. The American Statistician, 35(3):121–123, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683974.

Brown:1984:ETS

[BB84] Morton B. Brown and Judith Bromberg. An efficient two-stage procedure for generating random variates from the multinomial distribution. The American Statistician, 38(3):216–219, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683660.

Berry:1985:PFG

[BB85] Donald A. Berry and Timothy D. Berry. The probability of a field goal: Rating kickers. The American Statistician, 39(2):152–155, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682830.

Bradley:1989:CPD

[BB89] Edwin L. Bradley and Larry G. Blackwood. Comparing paired data: a simultaneous test for means and variances. The American Statistician, 43(4):234–235, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685368.

Blough:1989:LE

[BBC^{+}89] David K. Blough, Daniel G. Brooks, Steven Carroll, William A. Verdini, John T. Kent, S. R. Searle, Ravindra Khattree, Andrew R. Solow, Aaron A. Aardvark, I. J. Good, Gordon H. Fick, Hans Riedwyl, George D. Swanson, and Ralph B. D’Agostino. Letters to the Editor. The American Statistician, 43(3):186–188, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685072.

Bonnen:1981:IFSa

[BCF^{+}81] James T. Bonnen, Theodore G. Clemence, Ivan P. Fellegi, Thomas B. Jabine, Ronald E. Kutscher, Larry K. Roberson, and Charles A. Waite. Improving the federal statistical system: Report of the President’s Reorganization Project for the Federal Statistical System. The American Statistician, 35(4):184–196, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683286.

Brooks:1988:CDR

[BCV88] Daniel G. Brooks, Steven S. Carroll, and William A. Verdini. Characterizing the domain of a regression model. The American Statistician, 42(3):187–190, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684998.

Becker:1987:ADP

[BDMW87] Richard A. Becker, Lorraine Denby, Robert McGill, and Allan R. Wilks. Analysis of data from the places rated almanac. The American Statistician, 41(3):169–186, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685098.

Bowerman:1988:LE

[BDR^{+}88] Bruce L. Bowerman, John S. Dowd, Harry V. Roberts, Chester L. Olson, and A. S. Hedayat. Letters to the Editor. The American Statistician, 42(3):239–240, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685032.

Brown:1984:UPS

[BE84] Timothy C. Brown and Geoffrey K. Eagleson. A useful property of some symmetric statistics. The American Statistician, 38(1):63–65, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683564.

Beatty:1983:RGC

[Bea83] John C. Beatty. Raster graphics and color. The American Statistician, 37(1):60–75, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685849.

Bechhofer:1982:MJC

[Bec82] Robert Bechhofer. In memoriam: Jack Carl Kiefer, 1924–1981. The American Statistician, 36(4):356–357, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683084.

Bell:1980:BRB

[Bel80] C. B. Bell. Book review: [untitled]. The American Statistician, 34(4):236, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684069.

Belsley:1984:DCDa

[Bel84a] David A. Belsley. Demeaning conditioning diagnostics through centering. The American Statistician, 38(2):73–77, May 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683236.

Belsley:1984:DCDb

[Bel84b] David A. Belsley. [Demeaning conditioning diagnostics through centering]: Reply. The American Statistician, 38(2):90–93, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683241.

Bellhouse:1987:MBE

[Bel87] D. R. Bellhouse. Model-based estimation in finite population sampling. The American Statistician, 41(4):260–262, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684742.

Benjamini:1988:OBB

[Ben88] Yoav Benjamini. Opening the box of a boxplot. The American Statistician, 42(4):257–262, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685133.

Berengut:1981:SSS

[Ber81] David Berengut. Some small-sample properties of the Student t confidence intervals. The American Statistician, 35(3):144–147, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683981.

Bernardo:1984:PDB

[Ber84] Jose M. Bernardo. [On prior distributions for binary trials]: Comment. The American Statistician, 38(4):247–248, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683394.

Berk:1987:EMS

[Ber87a] Kenneth N. Berk. Effective microcomputer statistical software. The American Statistician, 41(3):222–228, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685112.

Berk:1987:SCS

[Ber87b] Kenneth N. Berk. Statistical computing software reviews: STATIS-TIX, Version 1.1, by Dennis Heisey and Gerard Nimis. The American Statistician, 41(3):229–231, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685113.

Berry:1987:IAC

[Ber87c] Donald A. Berry. Interim analysis in clinical trials. the role of the likelihood principle. The American Statistician, 41(2):117–122, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684222.

Bernardo:1988:OIP

[Ber88] Jose M. Bernardo. [Optimal information processing and Bayes’s theorem]: Comment. The American Statistician, 42(4):282, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685146.

Berk:1989:BRB

[Ber89] Richard A. Berk. Book review: [untitled]. The American Statistician, 43(1):55–56, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685174.

Best:1987:BRB

[Bes87] Al M. Best. Book review: [untitled]. The American Statistician, 41(4):321–324, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684755.

Beus:1981:RSF

[BF81] Gary B. Beus and J. Earl Faulkner. Review of statistical filmstrip programs. The American Statistician, 35(3):158–159, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683988.

Bennett:1983:EML

[BF83] Jay M. Bennett and John A. Flueck. An evaluation of major league baseball offensive performance models. The American Statistician, 37(1):76–82, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685850.

Bailar:1988:SPA

[BFM^{+}88] Barbara A. Bailar, Stephen E. Fienberg, Donald W. Marquardt, John Neter, and Janet L. Norwood. Strategic planning for the American Statistical Association, 1984–1987. The American Statistician, 42(1):1–9, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685252.

Butler:1980:BRB

[BG80] Michael D. Butler and Donald Guthrie. Book review: [untitled]. The American Statistician, 34(2):112–113, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684118.

Becketti:1987:RBP

[BG87] Sean Becketti and William Gould. Rangefinder box plots: a note. The American Statistician, 41(2):149, May 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684233.

Benichou:1989:DMI

[BG89] Jacques Benichou and Mitchell H. Gail. A delta method for implicitly defined random variables. The American Statistician, 43 (1):41–44, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685169.

Bancroft:1981:LE

[BGF^{+}81] T. A. Bancroft, Jean D. Gibbons, Rudolf J. Freund, Charles S. Davis, H. Leon Harter, A. P. Grieve, Joel R. Levin, William D. Schafer, Michael Kutner, Donna Brogan, and Colin R. Blyth. Letters to the Editor. The American Statistician, 35(3):176–179, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684001.

Brown:1981:SIS

[BH81] Barry W. Brown and Jay Herson. STPLAN: An interactive study planning package. The American Statistician, 35(3):164, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683994.

Bishop:1983:SRB

[BH83] Lane Bishop and William J. Hill. A study of the relationship between blood lead levels and occupational air lead levels. The American Statistician, 37(4):471–475, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683514.

Blom:1986:RWO

[BH86] Gunnar Blom and Lars Holst. Random walks of ordered elements with applications. The American Statistician, 40(4):271– 274, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684598.

Boardman:1980:PSC

[BHH^{+}80] Thomas J. Boardman, Gerald J. Hahn, William J. Hill, Ronald R. Hocking, William G. Hunter, William H. Lawton, R. Lyman Ott, Ronald D. Snee, and William E. Strawderman. Preparing statisticians for careers in industry: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Industry. The American Statistician, 34(2):65–75, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684106.

Bisgyer:1984:HAU

[Bis84] Edgar M. Bisgyer. [How to avoid undue influence of federal agencies on the editorial content of ASA journals]: Reply. The American Statistician, 38(3):228, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683666.

Brogan:1980:CAP

[BK80] Donna R. Brogan and Michael H. Kutner. Comparative analyses of pretest-posttest research designs. The American Statistician, 34(4):229–232, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684066.

Brogan:1986:GSS

[BK86] Donna Brogan and Michael H. Kutner. Graduate statistics service courses. The American Statistician, 40(3):252–254, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684561.

Blattenberger:1985:SBS

[BL85] Gail Blattenberger and Frank Lad. Separating the Brier score into calibration and refinement components: a graphical exposition. The American Statistician, 39(1):26–32, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683902.

Bar-Lev:1987:NPF

[BL87] Shaul K. Bar-Lev. Nonestimable parametric functions for some discrete distributions. The American Statistician, 41(3):199–200, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685102.

Blyth:1980:EAE

[Bly80] Colin R. Blyth. Expected absolute error of the usual estimator of the binomial parameter. The American Statistician, 34(3):155–157, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683873.

Balakrishnan:1986:NMO

[BM86] N. Balakrishnan and Henrick J. Malik. A note on moments of order statistics. The American Statistician, 40(2):147–148, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684877.

Balanda:1988:KCR

[BM88a] Kevin P. Balanda and H. L. MacGillivray. Kurtosis: a critical review. The American Statistician, 42(2):111–119, May 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684482.

Bickel:1988:NUB

[BM88b] P. J. Bickel and C. L. Mallows. A note on unbiased Bayes estimates. The American Statistician, 42(2):132–134, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684486.

Bloch:1988:NWL

[BM88c] Daniel A. Bloch and Lincoln E. Moses. Nonoptimally weighted least squares. The American Statistician, 42(1):50–53, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685260.

Barbara:1983:FSC

[BMM83] Vincent P. Barbara, Richard O. Mason, and Ian I. Mitroff. Federal statistics in a complex environment: The case of the 1980 Census. The American Statistician, 37(3):203–212, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683371.

Barnard:1985:SMC

[BMP^{+}85] Joseph Barnard, Wayne Myers, John Pearce, Fred Ramsey, Michael Sissenwine, and Woollcott Smith. Surveys for monitoring changes and trends in renewable resources: Forests and marine fisheries. The American Statistician, 39(4):363–373, November 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683016.

Binkley:1988:NES

[BN88] James K. Binkley and Carl H. Nelson. A note on the efficiency of seemingly unrelated regression. The American Statistician, 42(2):137–139, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684488.

Boardman:1982:BRB

[Boa82a] Thomas J. Boardman. Book review: Probability Problems for Programmable Calculators. The American Statistician, 36(2):134, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684028.

Boardman:1982:FSC

[Boa82b] Thomas J. Boardman. The future of statistical computing on desktop computers. The American Statistician, 36(1):49–58, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684095.

Boardman:1983:EGS

[Boa83] Thomas J. Boardman. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):11–12, February 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685825.

Boardman:1985:KRS

[Boa85] Thomas J. Boardman. [The key role of statisticians in the transformation of North American industry]: Comment. The American Statistician, 39(3):232–233, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

www.jstor.org/stable/2683947. See [Joi85a, Wac85, Hah85c,

God85].

Boen:1981:FRB

[Boe81] James R. Boen. Film review: The Reckoning. The American Statistician, 35(1):51–52, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683589.

Boen:1982:SSU

[Boe82] James R. Boen. A self-supporting university statistical consulting center. The American Statistician, 36(4):321–325, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683078.

Boland:1984:BGW

[Bol84] Philip J. Boland. A biographical glimpse of William Sealy Gosset. The American Statistician, 38(3):179–183, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683648.

Bolland:1985:UEN

[Bol85] Thomas W. Bolland. The use of the expected-net-gain chart to illustrate various aspects of sampling and sample design. The American Statistician, 39(1):47–52, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683906.

Bolstad:1986:HSF

[Bol86] William M. Bolstad. Harrison–Stevens forecasting and the multiprocess dynamic linear model. The American Statistician, 40(2): 129–135, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684871.

Bonnen:1981:IFSb

[Bon81] James T. Bonnen. [Improving the federal statistical system: Report of the President’s reorganization project for the federal statistical system]: Comment. The American Statistician, 35(4): 205–209, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683290.

Bonnen:1983:FSCa

[Bon83a] James T. Bonnen. Federal statistical coordination today: a disaster or a disgrace? The American Statistician, 37(3):179–192, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683366.

Bonnen:1983:FSCb

[Bon83b] James T. Bonnen. [Federal statistical coordination today: a disaster or a disgrace?]: Reply. The American Statistician, 37(3):199– 202, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2683370.

Bonnen:1988:SPS

[Bon88] James Bonnen. [Statistical policy for state and local governments]: Comment. The American Statistician, 42(1):16–18, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685254.

Bookbinder:1984:IOS

[Boo84] Albert I. A. Bookbinder. Investment-oriented statistical software. The American Statistician, 38(3):224, August 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683662.

Boruch:1982:ETEa

[Bor82a] Robert F. Boruch. Experimental tests in education: Recommendations from the holtzman report. The American Statistician, 36(1):1–8, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684082.

Boruch:1982:ETEb

[Bor82b] Robert F. Boruch. [Experimental tests in education: Recommendations from the holtzman report]: Rejoinder. The American Statistician, 36(1):13–14, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684085.

Boring:1987:TBE

[Bor87] John R. Boring III. Teaching biostatistics at the emory university school of medicine. The American Statistician, 41(4):269, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684746.

Box:1980:PSC

[Box80a] George E. P. Box. [Preparing statisticians for careers in industry: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Industry]: Comment. The American Statistician, 34(2):76–77, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684107.

Box:1980:RFD

[Box80b] Joan Fisher Box. R. A. Fisher and the design of experiments, 1922–1926. The American Statistician, 34(1):1–7, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682986.

Box:1981:GFD

[Box81] Joan Fisher Box. Gosset, Fisher, and the t distribution. The American Statistician, 35(2):61–66, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683142.

Bishop:1982:ALS

[BPT82] Thomas Bishop, Bruce Petersen, and David Trayser. Another look at the Statistician’s role in experimental planning and design. The American Statistician, 36(4):387–389, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683098.

Barcikowski:1984:DSG

[BR84] Robert S. Barcikowski and Randall R. Robey. Decisions in single group repeated measures analysis: Statistical tests and three computer packages. The American Statistician, 38(2):148–150, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683254.

Brauer:1982:PSC

[Bra82] Donald G. Brauer. [Preparing statisticians for careers in the federal government: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Government]: Comment. The American Statistician, 36(2):89, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684008.

Bradley:1983:WICa

[Bra83a] Ralph A. Bradley. What industry can do for statistics. The American Statistician, 37(1):21–24, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685836.

Bradley:1983:WICb

[Bra83b] Ralph A. Bradley. [What industry can do for statistics]: Reply. The American Statistician, 37(1):34–35, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685843.

Brogan:1980:PTC

[Bro80a] Donna R. Brogan. A program of teaching and consultation in research methods and statistics for graduate students in nursing. The American Statistician, 34(1):26–33, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682992.

Bross:1980:WSW

[Bro80b] Irwin D. J. Bross. When speaking to washington, tell the truth, the whole truth, and nothing but the truth, and do so intelligibly. The American Statistician, 34(1):34–38, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682993.

Bross:1983:AIS

[Bro83a] Irwin D. J. Bross. [Alternative interpretations of statistics on health effects of low-level radiation]: Comment. The American Statistician, 37(4):451–452, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683508.

Bross:1983:EGS

[Bro83b] Irwin D. J. Bross. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):12–13, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685826.

Bross:1984:HAU

[Bro84] Irwin Bross. How to avoid undue influence of federal agencies on the editorial content of ASA journals. The American Statistician, 38(3):226–228, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683665.

Broffitt:1986:ZCI

[Bro86] James D. Broffitt. Zero correlation, independence, and normality. The American Statistician, 40(4):276–277, November 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684600.

Bryant:1984:GSP

[Bry84] Peter Bryant. Geometry, statistics, probability: Variations on a common theme. The American Statistician, 38(1):38–48, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683558.

Beal:1980:NS

[BS80] Stuart Beal and Lewis Sheiner. The NONMEM system. The American Statistician, 34(2):118–119, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684123.

Beck:1983:ECL

[BS83] Gerald J. Beck and E. Neil Schachter. The evidence for chronic lung disease in cotton textile workers. The American Statistician, 37(4):404–412, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683503.

Barlow:1985:ARC

[BS85] Richard E. Barlow and Nozer D. Singpurwalla. Assessing the reliability of computer software and computer networks: An opportunity for partnership with computer scientists. The American Statistician, 39(2):88–94, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682802.

Buhler:1987:SRE

[BS87] Wolfgang J. Bühler and Joachim Sehr. Some remarks on exponential families. The American Statistician, 41(4):279–280, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684748.

Belli:1989:GSS

[BS89] Gabriella M. Belli and William L. Seaver. Graduate statistics service courses in part-time off-campus programs. The American Statistician, 43(2):86–90, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684504.

Bailey:1986:LE

[BSB^{+}86] Kent R. Bailey, David S. Salsburg, J. Calvin Berry, S. D. Peddada, D. V. Lindley, and George Casella. Letters to the Editor. The American Statistician, 40(3):255–258, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684562.

Bhrolchain:1986:MHE

[BT86] M. Ni Bhrolchain and I. Timaeus. The machine handling of event histories. The American Statistician, 40(1):56, February 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683128.

Buonaccorsi:1987:NCI

[Buo87] John P. Buonaccorsi. A note on confidence intervals for proportions in finite populations. The American Statistician, 41(3):215–218, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685108.

Burdick:1982:NMG

[Bur82] Richard K. Burdick. A note on the multivariate general linear test. The American Statistician, 36(2):131–132, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684026.

Burrows:1986:ESS

[Bur86] Peter M. Burrows. Extreme statistics from the sine distribution. The American Statistician, 40(3):216–217, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684539.

Burn:1987:BRB

[Bur87] David A. Burn. Book review: [untitled]. The American Statistician, 41(4):324–327, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684756.

Baksalary:1988:CTC

[BvE88] Jerzy K. Baksalary and Adrian C. van Eijnsbergen. A comparison of two criteria for ordinary-least-squares estimators to be best linear unbiased estimators. The American Statistician, 42(3):205–208, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685006.

Brecht:1984:GGP

[BW84] Mary-Lynn Brecht and J. Arthur Woodward. GANOVA: a general personal computer program for univariate and multivariate analysis of variance. The American Statistician, 38(1):69, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683567.

Boroto:1989:EIC

[BZ89] Daniel R. Boroto and Douglas A. Zahn. Editor’s invited column: Promoting statistics: On becoming valued and utilized. The American Statistician, 43(2):71–72, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684500.

Crump:1985:MQR

[CA85] Kenny S. Crump and Bruce C. Allen. Methods for quantitative risk assessment using occupational studies. The American Statistician, 39(4):442–450, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683028.

Chang:1987:GFS

[CA87] Potter C. Chang and A. A. Afifi. Goodness-of-fit statistics for general linear regression equations in the presence of replicated responses. The American Statistician, 41(3):195–199, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685101.

Campbell:1988:BWA

[Cam88] Robert A. Campbell. Balanced N-way analysis-of-variance program. The American Statistician, 42(3):226, August 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685013.

Cantor:1985:P

[Can85] Kenneth P. Cantor. Preface. The American Statistician, 39(4):431, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683025.

Carmer:1981:FEA

[Car81a] Samuel G. Carmer. Factorial experiments: Analysis of variance and randomized layouts. The American Statistician, 35(1):55, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683593.

Carmer:1981:LDA

[Car81b] Samuel G. Carmer. Lattice designs: Analysis of variance and randomized layouts. The American Statistician, 35(1):54–55, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683592.

Carroll:1983:WIC

[Car83] Mavis B. Carroll. [What industry can do for statistics]: Comment. The American Statistician, 37(1):31–32, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685841.

Carlstein:1986:SCR

[Car86] E. Carlstein. Simultaneous confidence regions for predictions. The American Statistician, 40(4):277–279, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684601.

Carter:1987:BIM

[Car87] Richard A. Carter. Biostatistics instruction at meharry medical college. The American Statistician, 41(4):268–269, November 1987.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684745.

Casella:1983:LRT

[Cas83] George Casella. Leverage and regression through the origin. The American Statistician, 37(2):147–152, May 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685876.

Casella:1985:IEB

[Cas85] George Casella. An introduction to empirical Bayes data analysis. The American Statistician, 39(2):83–87, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682801.

Chervany:1980:PSS

[CBI80] Norman L. Chervany, P. George Benson, and Raja K. Iyer. The planning stage in statistical reasoning. The American Statistician, 34(4):222–226, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684064.

Cohen:1986:EVE

[CBJ86] Steven B. Cohen, Vicki L. Burt, and Gretchen K. Jones. Efficiencies in variance estimation for complex survey data. The American Statistician, 40(2):157–164, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684880.

Collins:1989:ACD

[CC89] William H. Collins and Carol B. Collins. Analysis by computer of designed experiments with zero degrees of freedom for error variance. The American Statistician, 43(3):165–168, August 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685066.

Carter:1986:ACR

[CCW^{+}86] Walter H. Carter, Jr., Vernon M. Chinchilli, John D. Wilson, Eleanor D. Campbell, Fay K. Kessler, and Richard A. Carchman. An asymptotic confidence region for the ED_{100p} from the logistic response surface for a combination of agents. The American Statistician, 40(2):124–128, May 1986. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684870.

Cressie:1981:MGF

[CDFP81] Noel Cressie, Anne S. Davis, J. Leroy Folks, and George E. Policello II. The moment-generating function and negative integer moments. The American Statistician, 35(3):148–150, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683982.

Colton:1982:EEA

[CG82] Theodore Colton and E. Robert Greenberg. Epidemiologic evidence for adverse effects of DES exposure during pregnancy. The American Statistician, 36(3):268–272, August 1982. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683841.

Chan:1983:ACS

[CGL83] Tony F. Chan, Gene H. Golub, and Randall J. LeVeque. Algorithms for computing the sample variance: Analysis and recommendations. The American Statistician, 37(3):242–247, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683386.

Christensen:1985:BPE

[CH85] Ronald Christensen and Michael D. Huffman. Bayesian point estimation using the predictive distribution. The American Statistician, 39(4):319–321, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [CH86].

Christensen:1986:CBP

[CH86] Ronald Christensen and Michael D. Huffman. Correction to: “Bayesian point estimation using the predictive distribution” [amer. statist. 39 (1985), no. 4, part 1, 319–321]. The American Statistician, 40(3):258, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [CH85].

Chambers:1980:SCH

[Cha80] John M. Chambers. Statistical computing: History and trends. The American Statistician, 34(4):238–243, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684072.

Chao:1982:AAI

[Cha82] Anne Chao. Another approach to incomplete integrals. The American Statistician, 36(1):48, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684094.

Chatillon:1984:BRR

[Cha84] Guy Chatillon. The balloon rules for a rough estimate of the correlation coefficient. The American Statistician, 38(1):58–60, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683562.

Chernick:1982:NRD

[Che82] Michael R. Chernick. A note on the robustness of Dixon’s ratio test in small samples. The American Statistician, 36(2):140, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684033.

Chernoff:1986:WIE

[Che86] Herman Chernoff. [Why isn’t everyone a bayesian?]: Comment. The American Statistician, 40(1):5–6, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683106.

Chen:1988:FPF

[Che88] Shaw K. Chen. FDSM: a program for a fractional-differencing seasonal model. The American Statistician, 42(3):226, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685014.

Chow:1986:CS

[Cho86] Gregory C. Chow. Chinese statistics. The American Statistician, 40(3):191–196, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684534.

Cornelius:1989:LE

[CHP^{+}89] Wayne L. Cornelius, David C. Hoaglin, John W. Pratt, Samuel H. Cox, Robert M. Norton, Neil S. Fleming, Stephen M. Stigler, R. W. Farebrother, James L. Kepner, and Sandra Z. Keith. Letters to the Editor. The American Statistician, 43(4):289–292, November 1989.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685391.

Christensen:1981:BRB

[Chr81] Howard B. Christensen. Book review: [untitled]. The American Statistician, 35(4):259, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683304.

Chung:1988:WPF

[Chu88] Chang-Jo F. Chung. WIENER PACK: a FORTRAN 77 subroutine package for computing probabilities associated with Wiener (Brownian motion) and Brownian bridge processes. The American Statistician, 42(3):226–227, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685015.

Conover:1981:RTBa

[CI81a] W. J. Conover and Ronald L. Iman. Rank transformations as a bridge between parametric and nonparametric statistics. The American Statistician, 35(3):124–129, August 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683975.

Conover:1981:RTBb

[CI81b] W. J. Conover and Ronald L. Iman. [Rank transformations as a bridge between parametric and nonparametric statistics]: Rejoinder. The American Statistician, 35(3):132–133, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683978.

Campbell:1981:BRB

[CJ81] Cathy Campbell and Brian L. Joiner. Book review: [untitled]. The American Statistician, 35(3):159–160, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683989.

Chib:1987:ALS

[CJT87] Siddhartha Chib, S. Rao Jammalamadaka, and Ram C. Tiwari. Another look at some results on the recursive estimation in the general linear model. The American Statistician, 41(1):56–58, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684322.

Cleveland:1981:LPS

[Cle81] William S. Cleveland. LOWESS: a program for smoothing scatterplots by robust locally weighted regression. The American Statistician, 35(1):54, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683591.

Cleveland:1984:GMD

[Cle84a] William S. Cleveland. Graphical methods for data presentation: Full scale breaks, dot charts, and multibased logging. The American Statistician, 38(4):270–280, November 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683401.

Cleveland:1984:GSP

[Cle84b] William S. Cleveland. Graphs in scientific publications. The American Statistician, 38(4):261–269, November 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683400.

Cleveland:1983:CCO

[CM83] William S. Cleveland and Robert McGill. A color-caused optical illusion on a statistical graph. The American Statistician, 37(2):101– 105, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685868.

Cobb:1984:AAE

[Cob84] George W. Cobb. An algorithmic approach to elementary ANOVA. The American Statistician, 38(2):120–123, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683246.

Cohen:1982:EDC

[Coh82] Steven B. Cohen. Estimated data collection organization effect in the national medical care expenditure survey. The American Statistician, 36(4):337–341, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683081.

Cohen:1986:UPD

[Coh86] Joel E. Cohen. An uncertainty principle in demography and the unisex issue. The American Statistician, 40(1):32–39, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683114.

Cohen:1988:NCS

[Coh88] Jonathan D. Cohen. Noncentral chi-square: Some observations on recurrence. The American Statistician, 42(2):120–122, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684483.

Colton:1981:BCH

[Col81] Theodore Colton. Bill cochran: His contributions to medicine and public health and some personal recollections. The American Statistician, 35(3):167–170, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683997.

Collings:1983:CPD

[Col83] Bruce J. Collings. Characteristic polynomials by diagonal expansion. The American Statistician, 37(3):233–235, August 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683379.

Conlon:1987:MSC

[Con87] Michael Conlon. MLIB: a statistical computing library for the IBM-PC. The American Statistician, 41(4):329, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684759.

Conlon:1989:BRB

[Con89] Michael Conlon. Book review: [untitled]. The American Statistician, 43(3):171–174, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685068.

Cooper:1983:FAO

[Coo83] John C. B. Cooper. Factor analysis: An overview. The American Statistician, 37(2):141–147, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685875.

Cook:1984:DCD

[Coo84] R. D. Cook. [Demeaning conditioning diagnostics through centering]: Comment. The American Statistician, 38(2):78–79, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683237.

Cook:1987:BRB

[Coo87] R. Dennis Cook. Book review: [untitled]. The American Statistician, 41(3):233–236, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685115.

Corporation:1987:IIS

[Cor87] Lincoln Systems Corporation. ISP: Interactive statistical programs. The American Statistician, 41(3):238–239, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685118.

Cover:1989:DLG

[Cov89] Thomas M. Cover. Do longer games favor the stronger player? The American Statistician, 43(4):277–278, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685388.

Cox:1982:AWC

[Cox82] C. Philip Cox. An alternative way of calculating the X^{2} independence or association test statistic for a 2 × k contingency table. The American Statistician, 36(2):133, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684027.

Cox:1984:EIM

[Cox84] Christopher Cox. An elementary introduction to maximum likelihood estimation for multinomial models: Birch’s theorem and the delta method. The American Statistician, 38(4):283–287, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683403.

Charalambides:1984:LE

[CPE^{+}84] Ch. A. Charalambides, H. Papageorgiou, A. S. C. Ehrenberg, S. R. Searle, and Bruce Jay Collings. Letters to the Editor. The American Statistician, 38(1):71–72, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683569.

Cressie:1989:G

[Cre89] Noel Cressie. Geostatistics. The American Statistician, 43(4): 197–202, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685361.

Chen:1983:SSR

[CS83] Robert Chen and Larry A. Shepp. On the sum of symmetric random variables. The American Statistician, 37(3):237, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683381.

Chilton:1986:CRS

[CS86] Roland Chilton and Gordon F. Sutton. Classification by race and Spanish origin in the 1980 Census and its impact on white and nonwhite rates. The American Statistician, 40(3):197–201, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684535.

Carter:1986:RCU

[CSM86] Randy L. Carter, Richard L. Scheaffer, and Ronald G. Marks. The role of consulting units in statistics departments. The American Statistician, 40(4):260–264, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684596.

Chatterjee:1989:BRB

[CT89] Samprit Chatterjee and Aaron Tenenbein. Book review: [untitled]. The American Statistician, 43(4):244, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685371.

Carroll:1988:NAR

[CW88] Raymond J. Carroll and A. H. Welsh. A note on asymmetry and robustness in linear regression. The American Statistician, 42(4):285–287, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685149.

Dahiya:1981:IME

[Dah81] Ram C. Dahiya. An improved method of estimating an integerparameter by maximum likelihood. The American Statistician, 35 (1):34–37, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683579.

Dallal:1986:PSP

[Dal86a] Gerard E. Dallal. PC-SIZE: a program for sample-size determinations. The American Statistician, 40(1):52, February 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683121.

Dallal:1986:SSU

[Dal86b] Gerard E. Dallal. STATOOLS: Statistical utility programs. The American Statistician, 40(3):236, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684555.

Dallal:1987:OPL

[Dal87a] Gerard E. Dallal. OCTA: a program for the log-linear analysis of multidimensional contingency tables. The American Statistician, 41(4):328, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684757.

Dallal:1987:WWA

[Dal87b] Gerard E. Dallal. WUBA-WUBA: ASCII files to MYSTAT/ SYSTAT files. The American Statistician, 41(4):328, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684758.

Dallal:1988:GTM

[Dal88a] Gerard E. Dallal. GENERIC: a template for MYSTAT add-ons. The American Statistician, 42(3):227, August 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685016.

Dallal:1988:LLR

[Dal88b] Gerard E. Dallal. LOGISTIC: a logistic regression program for the IBM PC. The American Statistician, 42(4):272, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685137.

Dallal:1988:SML

[Dal88c] Gerard E. Dallal. Statistical microcomputing — like it is. The American Statistician, 42(3):212–216, August 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685009.

Dallal:1989:CCL

[Dal89a] Gerard E. Dallal. cLOGISTIC: a conditional logistic regression program for the IBM PC. The American Statistician, 43(2):125, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684520.

Dallal:1989:OCM

[Dal89b] Gerard E. Dallal. ODDJOB: a collection of miscellaneous statistical techniques. The American Statistician, 43(4):270, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685377.

Dallal:1989:PP

[Dal89c] Gerard E. Dallal. PCT: Percentiles. The American Statistician, 43(2):125–126, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684522.

Dallal:1989:TFP

[Dal89d] Gerard E. Dallal. Track: a FORTRAN program for calculating the Foulkes–Davis tracking index. The American Statistician, 43(2): 126, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684523.

Datta:1982:ETE

[Dat82] Lois-Ellin Datta. [Experimental tests in education: Recommendations from the holtzman report]: Comment. The American Statistician, 36(1):10–13, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684084.

Davis:1980:ULR

[Dav80] Anne S. Davis. Use of the likelihood ratio test on the inverse Gaussian distribution. The American Statistician, 34(2):108–110, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684116.

David:1981:MEP

[Dav81] Herbert A. David. In memoriam: Egon S. Pearson, 1895–1980. The American Statistician, 35(2):94–95, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683147.

David:1985:BUD

[Dav85a] H. A. David. Bias of S^{2} under dependence. The American Statistician, 39(3):201, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683929.

Davis:1985:BRB

[Dav85b] Marsha Davis. Book review: [untitled]. The American Statistician, 39(2):126, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682813.

Davis:1986:ETC

[Dav86] Linda June Davis. Exact tests for 2 × 2 contingency tables. The American Statistician, 40(2):139–141, May 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684874.

Davis:1989:BRB

[Dav89] William W. Davis. Book review: [untitled]. The American Statistician, 43(1):57–59, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685175.

Dawkins:1989:MAN

[Daw89] Brian Dawkins. Multivariate analysis of national track records. The American Statistician, 43(2):110–115, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684514.

Dickey:1986:URT

[DBM86] David A. Dickey, William R. Bell, and Robert B. Miller. Unit roots in time series models: Tests and implications. The American Statistician, 40(1):12–26, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683112.

Dennis:1985:PMG

[DBS^{+}85] B. Dennis, B. E. Brown, A. R. Stage, H. E. Burkhart, and S. Clark. Problems of modeling growth and yield of renewable resources. The American Statistician, 39(4):374–383, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683017.

DAgostino:1988:ASC

[DCB88] Ralph B. D’Agostino, Warren Chase, and Albert Belanger. The appropriateness of some common procedures for testing the equality of two independent binomial populations. The American Statistician, 42(3):198–202, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685002.

deCani:1986:NDI

[dCS86] John S. de Cani and Robert A. Stine. A note on deriving the information matrix for a logistic distribution. The American Statistician, 40(3):220–222, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684541.

Delfel:1986:CAL

[Del86] Norman E. Delfel. Complete automation of the layout and analysis of factorial experiments, including preparation of interaction tables. The American Statistician, 40(1):54–55, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683126.

DeMuth:1983:FSC

[DeM83] Christopher DeMuth. [Federal statistical coordination today: a disaster or a disgrace?]: Comment. The American Statistician, 37(3):198, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683369.

Deming:1986:SVI

[Dem86] Stanley N. Deming. SIMPLEX-V: An interactive computer program for experimental optimization. The American Statistician, 40(3):235, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684554.

Driscoll:1986:HDC

[DG86] Michael F. Driscoll and William R. Gundberg, Jr. A history of the development of Craig’s theorem. The American Statistician, 40(1):65–70, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683135.

Dallal:1989:MMD

[DGS^{+}89] Gerard E. Dallal, Shumei Guo, R. Simon, R. L. Klabansky, and J. E. Talmadge. MUDIFT: Multivariate distribution-free comparison of growth curves. The American Statistician, 43(2):125, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684521.

Darr:1985:FRF

[DH85] David R. Darr and Richard C. Hennemuth. Forecasting for resources in forestry and marine fisheries in the year 2000. The American Statistician, 39(4):384–392, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683018.

Draper:1987:RRS

[DH87] Norman R. Draper and Agnes M. Herzberg. A ridge-regression sidelight. The American Statistician, 41(4):282–283, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684750.

Davies:1988:MTT

[DHPR88] Kelly L. Davies, Dwight D. Harley, Trevor D. Petruk, and W. Todd Rogers. MEDPOL: Two- and three-dimensional Median Polish. The American Statistician, 42(3):227–228, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685017.

Dickey:1981:HPM

[Dic81] David A. Dickey. Histograms, percentiles, and moments. The American Statistician, 35(3):164–165, August 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683995.

Diderrich:1985:KFP

[Did85] George T. Diderrich. The Kalman filter from the perspective of Goldberger–Theil estimators. The American Statistician, 39 (3):193–198, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683927.

Dielman:1983:PCS

[Die83a] Terry E. Dielman. Pooled cross-sectional and time series data: a survey of current statistical methodology. The American Statistician, 37(2):111–122, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685870.

Diem:1983:SAS

[Die83b] J. E. Diem. A statistical assessment of the scientific evidence relating cotton dust exposure to chronic lung disease. The American Statistician, 37(4):395–403, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683502.

Dietz:1989:TRN

[Die89] E. Jacquelin Dietz. Teaching regression in a nonparametric statistics course. The American Statistician, 43(1):35–40, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685167.

Dixon:1988:BRB

[Dix88] Philip Dixon. Book review: [untitled]. The American Statistician, 42(2):158–159, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684493.

deJong:1984:GMC

[dJ84a] Piet de Jong. Gauss — a matrix calculator program. The American Statistician, 38(2):154, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683257.

Draper:1984:ROD

[DJ84b] Norman R. Draper and Brian L. Joiner. Residuals with one degree of freedom. The American Statistician, 38(1):55–57, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683561.

Dharmadhikari:1985:ENM

[DJD85] Sudhakar Dharmadhikari and Kumar Joag-Dev. Examples of nonunique maximum likelihood estimators. The American Statistician, 39(3):199–200, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683928.

Donner:1982:CPA

[DK82] Allan Donner and John J. Koval. CORREL: a program for the analysis of family data. The American Statistician, 36(4):384, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683094.

Dawes:1987:PRR

[DK87] Robyn M. Dawes and Joseph B. Kadane. Partial round-robin comparisons with perfect rankings. The American Statistician, 41 (3):204–205, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685104.

Dudewicz:1988:TVF

[DK88] Edward J. Dudewicz and Zaven A. Karian. TESTRAND for the VAX-11 family of computers: a random-number generation and testing library. The American Statistician, 42(3):228, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685018.

Dawson:1980:CS

[DKY80] Ree Dawson, John C. Klensin, and Douwe B. Yntema. The consistent system. The American Statistician, 34(3):169–176, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683876.

Doan:1982:RRA

[DL82] Thomas A. Doan and Robert B. Litterman. RATS-regression analysis of time series. The American Statistician, 36(1):63, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684101.

Devlin:1986:CAM

[DL86] Susan J. Devlin and Myrto Lefkopoulou. A characterization of 1979 ASA members. The American Statistician, 40(2):89–98, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684863.

Dempster:1981:MWG

[DM81] Arthur P. Dempster and Frederick Mosteller. In memoriam: William Gemmell Cochran, 1909–1980. The American Statistician, 35(1):38, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683580.

Dayton:1987:MMM

[DM87] C. Mitchell Dayton and George B. Macready. MODMULT and MODBIN: Microcomputer programs for concomitant-variable latent-class models. The American Statistician, 41(3):238, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685117.

Donner:1982:REP

[Don82] Allan Donner. The relative effectiveness of procedures commonly used in multiple regression analysis for dealing with missing values. The American Statistician, 36(4):378–381, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683092.

Dower:1985:APM

[Dow85] Roger C. Dower. Air pollution materials damage: Economics and the law. The American Statistician, 39(4):416–422, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683023.

Dupont:1985:CCD

[DP85] William D. Dupont and W. D. Plummer, Jr. CDM: Clinical data management software for interactive data entry and data quality assurance. The American Statistician, 39(3):211, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683937.

Denby:1987:EUG

[DP87] Lorraine Denby and Daryl Pregibon. An example of the use of graphics in regression. The American Statistician, 41(1):33–38, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684315.

Dudewicz:1983:TRN

[DR83] Edward J. Dudewicz and Thomas G. Ralley. TESTRAND: a random number generation and testing library. The American Statistician, 37(2):169–170, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685882.

Driscoll:1980:IUA

[Dri80] Michael F. Driscoll. Instructional uses of approximate convolutions and their graphs. The American Statistician, 34(3):150–154, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683872.

Driscoll:1986:NWK

[Dri86] Michael F. Driscoll. The Neter–Wasserman–Kutner models data. The American Statistician, 40(3):228, August 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684544.

Desu:1981:MSA

[DS81] M. M. Desu and Norman C. Severo. In memoriam: Sidney Addelman, 1932–1979. The American Statistician, 35(1):39, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683581.

Deming:1989:BTU

[DS89] W. Edwards Deming and Nozer D. Singpurwalla. A bear trap in using the Iowa-curve methodology for property retirements and depreciation charges. The American Statistician, 43(1):12–16, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685159.

Dawson-Saunders:1987:IBM

[DSAGR87] Beth Dawson-Saunders, Stanley Azen, Raymond S. Greenberg, and Allen H. Reed. The instruction of biostatistics in medical schools. The American Statistician, 41(4):263–266, November 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684743.

Dufour:1986:BLR

[Duf86] Jean-Marie Dufour. Bias of S^{2} in linear regressions with dependent errors. The American Statistician, 40(4):284–285, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684604.

Duncan:1981:IFS

[Dun81] Joseph W. Duncan. [Improving the federal statistical system: Report of the President’s reorganization project for the federal statistical system]: Comment. The American Statistician, 35(4):209, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683291.

Dunn:1987:VWF

[Dun87] Richard Dunn. Variable-width framed rectangle charts for statistical mapping. The American Statistician, 41(2):153–156, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684235.

Duncan:1988:SPS

[Dun88a] Joseph W. Duncan. [Statistical policy for state and local governments]: Comment. The American Statistician, 42(1):21–22, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685256.

Dunn:1988:FRC

[Dun88b] Richard Dunn. Framed rectangle charts or statistical maps with shading: An experiment in graphical perception. The American Statistician, 42(2):123–129, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684484.

Dunn:1989:DAT

[Dun89] Richard Dunn. A dynamic approach to two-variable color mapping. The American Statistician, 43(4):245–252, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685372.

Dusoir:1986:TVS

[Dus86] Tony Dusoir. TDDA V1.2: Statistics software for the Apple. The American Statistician, 40(1):55, February 1986. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683127.

Dusoir:1989:SSC

[Dus89] Tony Dusoir. The SC statistical calculator. The American Statistician, 43(4):270–271, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685378.

Dallal:1986:AAD

[DW86] Gerard E. Dallal and Leland Wilkinson. An analytic approximation to the distribution of Lilliefors’s test statistic for normality. The American Statistician, 40(4):294–296, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684607.

Engeman:1984:SCG

[ED84] Richard M. Engeman and William E. Dusenberry. Statistical consulting in a geographically dispersed organization. The American Statistician, 38(4):281–283, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683402.

Edgeman:1984:PDM

[Edg84] Rick L. Edgeman. P-DIST: a microcomputer program for obtaining probabilities from univariate distributions. The American Statistician, 38(4):321, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683426.

Edmonston:1985:MCC

[Edm85] Barry Edmonston. MICRO-CLUSTER: Cluster analysis software for microcomputers. The American Statistician, 39(3):209, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683934.

Edwards:1984:RRS

[Edw84] Howard P. Edwards. RANKSEL: a ranking and selection package. The American Statistician, 38(2):158–159, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683263.

Edwards:1986:RHB

[Edw86] A. W. F. Edwards. Is the reference in Hartley (1749) to Bayesian inference. The American Statistician, 40(2):109–110, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684867.

Elder:1989:LE

[EFJ^{+}89] Robert S. Elder, Neil S. Fleming, M. C. Jones, R. W. Farebrother, and Nils Lid Hjort. Letters to the Editor. The American Statistician, 43(1):68–69, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685180.

Efron:1986:WIEb

[Efr86a] B. Efron. [Why isn’t everyone a Bayesian?]: Reply. The American Statistician, 40(1):11, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683111.

Efron:1986:WIEa

[Efr86b] B. Efron. Why isn’t everyone a Bayesian? With discussion and a reply by the author. The American Statistician, 40(1):1–5, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683105.

Efron:1983:LLB

[EG83] Bradley Efron and Gail Gong. A leisurely look at the bootstrap, the jackknife, and cross-validation. The American Statistician, 37 (1):36–48, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685844.

Eisenberg:1987:IED

[EG87] Bennett Eisenberg and B. K. Ghosh. Independent events in a discrete uniform probability space. The American Statistician, 41 (1):52–56, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684321.

Ehrenberg:1981:PN

[Ehr81] A. S. C. Ehrenberg. The problem of numeracy. The American Statistician, 35(2):67–71, May 1981. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683143.

Ehrenberg:1982:WTP

[Ehr82] A. S. C. Ehrenberg. Writing technical papers or reports. The American Statistician, 36(4):326–329, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683079.

Ehrenberg:1983:DLS

[Ehr83a] A. S. C. Ehrenberg. Deriving the least squares regression equation. The American Statistician, 37(3):232, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683378.

Ehrenberg:1983:WMP

[Ehr83b] A. S. C. Ehrenberg. We must preach what is practiced. The American Statistician, 37(3):248–250, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683387.

Ehrenberg:1984:WMP

[Ehr84] A. S. C. Ehrenberg. [We must preach what should be practiced]: Reply. The American Statistician, 38(2):152–153, May 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683256.

Eddy:1987:GEC

[EJKS87] William F. Eddy, Albyn C. Jones, Robert E. Kass, and Mark J. Schervish. Graduate education in computational statistics. The American Statistician, 41(1):60–64, February 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684324.

Engeman:1988:EMS

[EL88] Richard M. Engeman and Amber D. Lusk. Expected mean squares on MS-DOS. The American Statistician, 42(3):229–230, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685021.

Elashoff:1984:CSP

[Ela84] Janet D. Elashoff. [Contrasting split plot and repeated measures experiments and analyses]: Comment. The American Statistician, 38(1):27–29, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683553.

Ellenberg:1983:EGS

[Ell83] Jonas H. Ellenberg. Ethical guidelines for statistical practice: a historical perspective. The American Statistician, 37(1):1–4, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685819.

Elliott:1987:BRB

[Ell87] J. Richard Elliott. Book review: [untitled]. The American Statistician, 41(3):220–221, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685110.

Elliott:1988:KLC

[Ell88] Alan C. Elliott. KWIKSTAT: a low-cost interactive statistical analysis package for IBM-PC-compatible computers. The American Statistician, 42(3):228–229, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685019.

Elliott:1989:KV

[Ell89] Alan C. Elliott. Kwikstat (Version 2.00). The American Statistician, 43(4):271, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685379.

Enas:1988:PQE

[Ena88] Nathan Enas. PASS: a quick and easy power and sample-size calculation program. The American Statistician, 42(3):229, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685020.

Easterling:1980:LE

[ENH^{+}80] Robert G. Easterling, Horace W. Norton, Richard M. Heiberger, Larry L. Laster, F. M. Speed, Charles J. Monlezun, Richard C. Taeuber, and Robert P. Nott. Letters to the Editor. The American Statistician, 34(1):61–63, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683007.

Ernster:1983:MAB

[Ern83] Virginia L. Ernster. The measurement of associations between environmental exposures and cancer. The American Statistician, 37(4):420–426, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683505.

Edgeman:1986:GPS

[ES86] Rick L. Edgeman and Robert C. Scott. GENERATOR: a program for simulating nonuniform random variates. The American Statistician, 40(1):54, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683125.

Esty:1985:ALD

[Est85] Warren W. Esty. Another look at daily lotteries. The American Statistician, 39(2):118–119, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682809.

Eubank:1985:SVD

[EW85] R. L. Eubank and J. T. Webster. The singular-value decomposition as a tool for solving estimability problems. The American Statistician, 39(1):64–66, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683912.

Eldridge:1982:PSC

[EWW^{+}82] Marie D. Eldridge, Katherine K. Wallman, Rolf M. Wulfsberg, Barbara A. Bailar, Yvonne M. Bishop, William E. Kibler, Beatrice S. Orleans, Dorothy P. Rice, Wesley Schaible, Seymour M. Selig, and Monroe G. Sirken. Preparing statisticians for careers in the federal government: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Government. The American Statistician, 36(2):69–81, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684004.

Fairley:1986:CTC

[Fai86] David Fairley. Cherry trees with cones? The American Statistician, 40(2):138–139, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684873.

Farlow:1981:GAI

[Far81] Stanley J. Farlow. The GMDH algorithm of ivakhnenko. The American Statistician, 35(4):210–215, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683292.

Farnum:1989:APS

[Far89] Nicholas R. Farnum. An alternate proof of Samuelson’s inequality and its extensions. The American Statistician, 43(1):46–47, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685171.

Friedman:1983:ARC

[FBC^{+}83] Robert M. Friedman, Rosina M. Bierbaum, Patricia A. Catherwood, Stuart C. Diamond, George G. Hoberg, Jr., and Valerie Ann Lee. The acid rain controversy: The limits of confidence. The American Statistician, 37(4):385–394, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683501.

Fenwick:1981:PPP

[FC81] Ian Fenwick and Sangit Chatterjee. Perception, preference, and patriotism: An exploratory analysis of the 1980 winter olympics. The American Statistician, 35(3):170–173, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683998.

Flack:1987:FSN

[FC87] Virginia F. Flack and Potter C. Chang. Frequency of selecting noise variables in subset regression analysis: a simulation study. The American Statistician, 41(1):84–86, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684336.

Ferson:1988:MSS

[Fer88] Scott Ferson. Microcomputer software for stochastic demography and ecological risk analysis. The American Statistician, 42(4):273, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685140.

Feldman:1980:TSI

[FF80] Dorian Feldman and Martin Fox. Three simple inductive proofs in probability and statistics. The American Statistician, 34(1):50–51, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682998.

Freeman:1983:PSP

[FGHK83] Daniel H. Freeman, Jr., María Elena González, David C. Hoaglin, and Beth A. Kilss. Presenting statistical papers. The American Statistician, 37(2):106–110, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685869.

Franck:1988:BAS

[FHI^{+}88] Wallace E. Franck, John E. Hewett, Mohammed Z. Islam, Sue Jane Wang, and Michael Cooperstock. A Bayesian analysis suitable for classroom presentation. The American Statistician, 42(1):75–77, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685267.

Frigge:1989:SIB

[FHI89] Michael Frigge, David C. Hoaglin, and Boris Iglewicz. Some implementations of the boxplot. The American Statistician, 43(1):50–54, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685173.

Fienberg:1983:FSC

[Fie83] Stephen E. Fienberg. [Federal statistical coordination today: a disaster or a disgrace?]: Comment. The American Statistician, 37 (3):192–196, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683367.

Field:1984:ROE

[Fie84] Christopher Field. A “Reverse order” elementary statistics course. The American Statistician, 38(2):117–119, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683245.

Firester:1982:ANP

[Fir82] Lynne Firester. ANOVA — a new P-STAT command. The American Statistician, 36(4):385–386, November 1982. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683096.

Feingold:1986:CDB

[FK86] Marcia Feingold and P. E. Korsog. The correlation and dependence between two F statistics with the same denominator. The American Statistician, 40(3):218–220, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684540.

Fuller:1987:MTA

[FK87] Wayne A. Fuller and Oscar Kempthorne. In memoriam: Theodore Alfonso Bancroft, 1907–1986. The American Statistician, 41(3): 167–168, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685097.

Fleiss:1982:SCL

[Fle82] Joseph L. Fleiss. A simplification of the classic large-sample standard error of a function of multinomial proportions. The American Statistician, 36(4):377–378, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683091.

Fligner:1981:RTB

[Fli81] Michael A. Fligner. [Rank transformations as a bridge between parametric and nonparametric statistics]: Comment. The American Statistician, 35(3):131–132, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683977.

Flury:1986:SRV

[Flu86] Bernhard K. Flury. On sums of random variables and independence. The American Statistician, 40(3):214–215, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684538.

Flury:1989:UPS

[Flu89] Bernhard W. Flury. Understanding partial statistics and redundancy of variables in regression and discriminant analysis. The American Statistician, 43(1):27–31, February 1989. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685165.

Fishman:1984:SDD

[FM84] George S. Fishman and Louis R. Moore III. Sampling from a discrete distribution while preserving monotonicity. The American Statistician, 38(3):219–223, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683661.

Frome:1989:EGF

[FM89] Edward L. Frome and Max D. Morris. Evaluating goodness of fit of Poisson regression models in cohort studies. The American Statistician, 43(3):144–147, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685060.

Folks:1982:RFU

[Fol82] J. Leroy Folks. Review of four UMAP modules. The American Statistician, 36(2):135, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684029.

Fowles:1988:ME

[Fow88] Richard Fowles. MICRO-EBA. The American Statistician, 42 (4):274, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685141.

Foxwell:1983:STS

[Fox83] Harry J. Foxwell. SURVTAB: Tabulation software for survey questionnaire data. The American Statistician, 37(2):175, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685890.

Fox:1989:CSS

[Fox89] David R. Fox. Computer selection of size-biased samples. The American Statistician, 43(3):168–171, August 1989. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685067.

Fahmy:1981:BDO

[FP81] Salwa Fahmy and Frank Proschan. Bounds on differences of order statistics. The American Statistician, 35(1):46–47, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683585.

Freund:1987:NLQ

[FP87] John E. Freund and Benjamin M. Perles. A new look at quartiles of ungrouped data. The American Statistician, 41(3):200–203, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685103.

Freedman:1989:RNS

[FP89] Laurence S. Freedman and David Pee. Return to a note on screening regression equations. The American Statistician, 43(4):279– 282, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685389.

Flury:1985:TLT

[FR85] Bernhard Flury and Hans Riedwyl. T 2 tests, the linear two-group discriminant function, and their computation by linear regression. The American Statistician, 39(1):20–25, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683901.

Flury:1986:SDU

[FR86] Bernhard K. Flury and Hans Riedwyl. Standard distance in univariate and multivariate analysis. The American Statistician, 40 (3):249–251, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684560.

Frane:1981:VB

[Fra81a] James W. Frane. The 1981 version of BMDP. The American Statistician, 35(4):261–262, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683306.

Frane:1981:SED

[Fra81b] James W. Frane. [The specification of experimental designs to ANOVA programs]: Comment. The American Statistician, 35(2): 105–106, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2683151.

Freund:1980:CMC

[Fre80] Rudolf J. Freund. The case of the missing cell. The American Statistician, 34(2):94–98, May 1980. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684112.

Free:1983:EGS

[Fre83a] Spencer M. Free. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):17, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685832.

Freedman:1983:NSR

[Fre83b] David A. Freedman. A note on screening regression equations. The American Statistician, 37(2):152–155, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685877.

Frome:1981:PRA

[Fro81] Edward L. Frome. Poisson regression analysis. The American Statistician, 35(4):262–263, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683307.

Fawcett:1987:DSC

[FS87] Richard F. Fawcett and Kathleen C. Salter. Distributional studies and the computer: An analysis of Durbin’s rank test. The American Statistician, 41(1):81–83, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684335.

Foster:1989:MPE

[FS89] David A. Foster and Kevin M. Sullivan. Microcomputer programs for exact calculation or maximum likelihood estimation of selected epidemiologic parameters. The American Statistician, 43(2):126, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684524.

Fox:1989:LE

[FSJ^{+}89] John Fox, Shayle Searle, Ian Jolliffe, John Mandel, B. J. Mandel, Peter C. Kiessler, Ralph L. Disney, Charles J. Mode, Roderick J. A. Little, Leigh W. Murray, John I. Marden, and Kevin P. Balanda. Letters to the Editor. The American Statistician, 43(2):129–132, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684528.

Fischhoff:1982:LFE

[FSL82] Baruch Fischhoff, Paul Slovic, and Sarah Lichtenstein. Lay foibles and expert fables in judgments about risk. The American Statistician, 36(3):240–255, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683835.

Fuchs:1980:LE

[FSM^{+}80] Camil Fuchs, George A. F. Seber, Marvin S. Margolis, Thomas A. Ryan, and Thomas A. Ryan, Jr. Letters to the Editor. The American Statistician, 34(2):122–123, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684126.

Farebrother:1987:LE

[FSW^{+}87] R. W. Farebrother, D. J. Saville, G. R. Wood, Edward L. Melnick, Aaron Tenenbein, James D. Broffitt, Hubert Lilliefors, Gerard E. Dallal, Leland Wilkinson, David J. Finney, James D. Knoke, Kent R. Bailey, David Judkins, Bob Wheeler, James R. Knaub, Jr., Dale Zimmerman, E. Carlstein, and H. S. Konijn. Letters to the Editor. The American Statistician, 41(3):242–248, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685123.

Freni-Titulaer:1984:CSG

[FTL84] Lambertina W. J. Freni-Titulaer and William C. Louv. Comparisons of some graphical methods for exploratory multivariate data analysis. The American Statistician, 38(3):184–188, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683649.

Fu:1984:MDD

[Fu84] James C. Fu. The moments do not determine a distribution. The American Statistician, 38(4):294–295, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683407.

Fuller:1980:BVS

[Ful80] Wayne A. Fuller. Balkrishna V. Sukhatme (1924–1979). The American Statistician, 34(1):49, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682997.

Galbraith:1984:SMS

[Gal84] Richard C. Galbraith. SIGSTAT: Mainframe statistics for the personal computer. The American Statistician, 38(4):317, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683419.

Games:1981:ASA

[Gam81] Paul A. Games. The ANOVR series: Analyses for repeated measure designs with optional robust procedures when assumptions are violated. The American Statistician, 35(2):109–110, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683155.

Gastwirth:1989:CRP

[Gas89] Joseph L. Gastwirth. [A category representation paradox]: Comment. The American Statistician, 43(2):95–96, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684508.

Gates:1989:EIC

[Gat89] Charles E. Gates. Editor’s invited column. The American Statistician, 43(1):1, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685156.

Ghosh:1987:LE

[GBL^{+}87] Subir Ghosh, Kenneth P. Burnham, Nico F. Laubscher, Gerard E. Dallal, Leland Wilkinson, Donald F. Morrison, Milton W. Loyer, Bennett Eisenberg, Solomon Kullback, Ian T. Jolliffe, and Jeffrey S. Simonoff. Letters to the Editor. The American Statistician, 41(4):338–341, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684769.

Grimson:1981:CPP

[GD81] Roger C. Grimson and Daniel DalCorso. CASCNTRL — a PL/I program for the analysis of case-control studies. The American Statistician, 35(3):163–164, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683993.

Ghosh:1989:UMV

[GD89] Malay Ghosh and Gauri Sankar Datta. Uniformly minimum variance unbiased estimation for discrete distributions with support depending on the parameter. The American Statistician, 43(1):17–19, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685160.

Gebert:1981:CCS

[Geb81] James R. Gebert. COMPSTAT: a comprehensive statistical package for microcomputers. The American Statistician, 35(2):111, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683157.

Gebski:1989:SVS

[Geb89] Val Gebski. SPIDA (Version 5.25): Statistical Package for Interactive Data Analysis. The American Statistician, 43(4):272, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685380.

Gehan:1983:EGS

[Geh83] Edmund A. Gehan. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):8–9, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685822.

Geisser:1984:PDBb

[Gei84a] Seymour Geisser. [On prior distributions for binary trials]: Reply. The American Statistician, 38(4):250–251, November 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683397.

Geisser:1984:PDBa

[Gei84b] Seymour Geisser. On prior distributions for binary trials. with comments by Jose M. Bernardo, Melvin R. Novick and Arnold Zellner and a reply by the author. The American Statistician, 38(4):244–247, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2683393. See correction [Zel85].

Gentle:1985:MIS

[Gen85] James E. Gentle. [More intelligent statistical software and statistical expert systems: Future directions]: Comment. The American Statistician, 39(1):8–9, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683896.

Gentleman:1986:NSE

[Gen86] Jane F. Gentleman. Note from the section editor. The American Statistician, 40(3):237, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684557.

Gibbons:1980:OSC

[GF80] Jean D. Gibbons and Rudolf J. Freund. Organizations for statistical consulting at colleges and universities. The American Statistician, 34(3):140–145, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683870.

Godambe:1985:LE

[GFL^{+}85] V. P. Godambe, Bernhard N. Flury, William C. Louv, Lambertina W. J. Freni-Titulaer, Josue Guzman, Agustin Ayuso, Shey Shiung Sheu, and Leon Katz. Letters to the Editor. The American Statistician, 39(2):156–158, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682831.

Garrett:1980:CPU

[GG80a] Robert G. Garrett and Thomas I. Goss. A computer program for unbalanced nested analysis of variance. The American Statistician, 34(1):60, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683006.

Gentle:1980:REI

[GG80b] James E. Gentle and Walton C. Gregory. Recent enhancements to the IMSL subroutine library. The American Statistician, 34(4):246, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684075.

Greenhouse:1980:JC

[GH80] Samuel W. Greenhouse and Max Halperin. Jerome Cornfield, 1912–1979. The American Statistician, 34(2):106–107, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684115.

Gale:1982:CBG

[GH82] Nathan Gale and William C. Halperin. A case for better graphics: The unclassed choropleth map. The American Statistician, 36(4):330–336, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683080.

Galpin:1984:URR

[GH84] Jacqueline S. Galpin and Douglas M. Hawkins. The use of recursive residuals in checking model fit in linear regression. The American Statistician, 38(2):94–105, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683242.

Gibberd:1988:RSA

[GH88] R. W. Gibberd and T. Hakulinen. A relative survival analysis package. The American Statistician, 42(3):230, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685022.

Gilbreath:1980:BRB

[Gil80] Glenn H. Gilbreath. Book review: [untitled]. The American Statistician, 34(4):237, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684071.

Gilula:1981:PCS

[Gil81] Zvi Gilula. On Pearson’s chi-squared statistic for independence in two-way contingency tables. The American Statistician, 35(2):96– 97, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683148.

Gill:1984:CSP

[Gil84] J. L. Gill. [Contrasting split plot and repeated measures experiments and analyses]: Comment. The American Statistician, 38 (1):29–30, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683554.

Gilpin:1986:AAH

[Gil86] Andrew R. Gilpin. APCLUST: Agglomerative hierarchical clustering analysis program for microcomputers. The American Statistician, 40(1):53, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683124.

Gaver:1984:RRI

[GK84] Donald P. Gaver and Karen Kafadar. A retrievable recipe for inverse t. The American Statistician, 38(4):308–311, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683413.

Gardner:1983:AIS

[GL83] John W. Gardner and Joseph L. Lyon. [Alternative interpretations of statistics on health effects of low-level radiation]: Comment. The American Statistician, 37(4):452–453, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683509.

Gleason:1986:TLF

[Gle86a] John R. Gleason. TURBO RAND: a library for fast Monte Carlo sampling and simulation on microcomputers. The American Statistician, 40(3):233, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684551.

Gleser:1986:SNR

[Gle86b] Leon J. Gleser. Some notes on refereeing. The American Statistician, 40(4):310–312, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684615.

Gleason:1988:ABB

[Gle88] John R. Gleason. Algorithms for balanced bootstrap simulations. The American Statistician, 42(4):263–266, November 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685134.

Gleser:1989:GDM

[Gle89] Leon Jay Gleser. The gamma distribution as a mixture of exponential distributions. The American Statistician, 43(2):115–117, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684515.

Goodisman:1981:LE

[GLM^{+}81] Leonard D. Goodisman, Fred C. Leone, Morton A. Meyer, Clark C. Abt, and Howard Wainer. Letters to the Editor. The American Statistician, 35(1):56–58, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683594.

Genest:1986:JCB

[GM86] Christian Genest and Jock MacKay. The joy of copulas: bivariate distributions with uniform marginals. The American Statistician, 40(4):280–283, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684602.

Gunter:1984:LE

[GMB^{+}84] Bert Gunter, Paul D. Minton, James R. Boen, Carl J. Johnson, William M. Bolstad, D. J. Finney, C. R. Henderson, and Nathan Mantel. Letters to the Editor. The American Statistician, 38 (3):230–233, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683669.

Good:1985:LE

[GN85] I. J. Good and Gottfried E. Noether. Letters to the Editor. The American Statistician, 39(1):81, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683919.

Gill:1982:IIC

[GNS82] D. S. Gill, Kendall E. Nygard, and Muhammad Syamsun. INTEL-REG: An interactive computer package for linear regression analyses. The American Statistician, 36(1):60, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684097.

Godfrey:1985:KRS

[God85] A. Blanton Godfrey. [The key role of statisticians in the transformation of North American industry]: Comment. The American Statistician, 39(3):231–232, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www. jstor.org/stable/2683946. See [Joi85a, Wac85, Hah85c, Boa85].

Goldstein:1982:PIM

[Gol82] Michael Goldstein. Preliminary inspection of multivariate data. The American Statistician, 36(4):358–362, November 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683085.

Golbeck:1986:PAC

[Gol86] Amanda L. Golbeck. Probabilistic approaches to current life table estimation. The American Statistician, 40(3):185–190, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684533.

Goldstein:1988:BRB

[Gol88] Richard Goldstein. Book review: [untitled]. The American Statistician, 42(3):217–219, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685010.

Goldstein:1989:PSS

[Gol89] Richard Goldstein. Power and sample size via MS/PC-DOS computers. The American Statistician, 43(4):253–260, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685373.

Gonzalez:1980:CFD

[Gon80] Maria Elena Gonzalez. Characteristics of formulas and data used in the allocation of federal funds. The American Statistician, 34 (4):200–211, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684061.

Gumpertz:1989:SAI

[GP89] Marcia Gumpertz and Sastry G. Pantula. A simple approach to inference in random coefficient models. The American Statistician, 43(4):203–210, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685362.

Greene:1983:LPE

[Gre83a] William H. Greene. LIMDEP: a program for estimating the parameters of qualitative and limited dependent variable models. The American Statistician, 37(2):170, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685883.

Greenhouse:1983:EGS

[Gre83b] Samuel W. Greenhouse. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):15–16, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685830.

Greene:1985:LEM

[Gre85] William H. Greene. LIMDEP: An econometric modeling program for the IBM PC. The American Statistician, 39(3):210, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683936.

Greenacre:1986:SPP

[Gre86a] Michael Greenacre. SIMCA: a program to perform simple correspondence analysis. The American Statistician, 40(3):230–231, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684547.

Greene:1986:ASF

[Gre86b] William Greene. Analysis of survival and failure time data with LIMDEP. The American Statistician, 40(3):234–235, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684553.

Gross:1982:MML

[Gro82] Alan L. Gross. MLIN: a multivariate linear model program. The American Statistician, 36(1):61, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684098.

Groggel:1986:DFT

[GS86] David J. Groggel and John H. Skillings. Distribution-free tests for main effects in multifactor designs. The American Statistician, 40(2):99–102, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684864.

Griffin:1982:DNE

[GSW82] Malcolm P. Griffin, J. Terry Smith, and Donald G. Watts. Deriving the normal and exponential densities using EDA techniques. The American Statistician, 36(4):373–377, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683090.

Guenther:1981:SSF

[Gue81] William C. Guenther. Sample size formulas for normal theory T tests. The American Statistician, 35(4):243–244, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683297.

Guenther:1987:SIT

[Gue87] William C. Guenther. Statistical inferences for two one-parameter families. The American Statistician, 41(1):50–52, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684320.

Gunst:1984:CTB

[Gun84] Richard F. Gunst. Comment: Toward a balanced assessment of collinearity diagnostics. The American Statistician, 38(2):79–82, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683238.

Gupta:1982:QSE

[Gup82] Shanti S. Gupta. [Quality of statistical education: Should ASA assist or assess?]: Comment. The American Statistician, 36(2):94, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684010.

Gussio:1986:DSS

[Gus86] Rick P. Gussio. DESCRIP-STAT SELECTOR: a microcomputer program using proportional reduction of error coefficients for selecting a measure of central location. The American Statistician, 40(2):

171, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684885.

Gelfand:1980:RVC

[GW80] Alan E. Gelfand and Crayton C. Walker. Review of the video cassette set: Statistics for Problem Solving and Decision Making. The American Statistician, 34(1):56, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683001.

Goldberg:1981:EMS

[GW81] Judith D. Goldberg and Janet T. Wittes. The evaluation of medical screening procedures. The American Statistician, 35(1):4–11, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683576.

Gan:1989:EPO

[GY89] Gaoxiong Gan and Shie-Shien Yang. An elementary proof of the order-statistics characterization of the Poisson process. The American Statistician, 43(1):45–46, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685170. See correction [GY90].

Gan:1990:CEP

[GY90] Gaoxiong Gan and Shie-Shien Yang. Correction: “An elementary proof of the order-statistics characterization of the Poisson process” [Amer. Statist. 43 (1989), no. 1, 45–46]. The American Statistician, 44(1):65, ???? 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [GY89].

Hauck:1984:SRR

[HA84] Walter W. Hauck and Sharon Anderson. A survey regarding the reporting of simulation studies. The American Statistician, 38 (3):214–216, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683659.

Hauck:1986:CLS

[HA86] Walter W. Hauck and Sharon Anderson. A comparison of largesample confidence interval methods for the difference of two binomial probabilities. The American Statistician, 40(4):318–322, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684618.

Holland:1989:FRD

[HA89] Burt Holland and Mohammad Ahsanullah. Further results on a distribution of meinhold and singpurwalla. The American Statistician, 43(4):216–219, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685364.

Hacker:1987:MSM

[Hac87] Carl S. Hacker. Medical statistics at the Medical School of the University of Texas Health Science Center at Houston. The American Statistician, 41(4):266–267, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684744.

Hahn:1985:MISa

[Hah85a] Gerald J. Hahn. More intelligent statistical software and statistical expert systems: Future directions. The American Statistician, 39(1):1–8, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683895.

Hahn:1985:MISb

[Hah85b] Gerald J. Hahn. [More intelligent statistical software and statistical expert systems: Future directions]: Reply. The American Statistician, 39(1):15–16, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683899.

Hahn:1985:KRS

[Hah85c] Gerald J. Hahn. [The key role of statisticians in the transformation of North American industry]: Comment. The American Statistician, 39(3):229–231, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2683945. See [Joi85a, Wac85, God85, Boa85].

Hahn:1989:SAM

[Hah89] Gerald J. Hahn. Statistics-aided manufacturing: a look into the future. The American Statistician, 43(2):74–79, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684502.

Hall:1988:RDT

[Hal88] Bronwyn H. Hall. Recent developments for the TSP program. The American Statistician, 42(1):85–86, February 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685273.

Hallahan:1989:BRB

[Hal89] Charles Hallahan. Book review: [untitled]. The American Statistician, 43(4):263–266, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685375.

Hamer:1981:PEC

[Ham81] Robert M. Hamer. Papers that evaluate computer programs. The American Statistician, 35(4):264, November 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683308.

Hamilton:1983:AISa

[Ham83a] L. D. Hamilton. Alternative interpretations of statistics on health effects of low-level radiation. The American Statistician, 37(4): 442–451, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683507.

Hamilton:1983:AISb

[Ham83b] L. D. Hamilton. [Alternative interpretations of statistics on health effects of low-level radiation]: Rejoinder. The American Statistician, 37(4):457–458, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683512.

Hamedani:1984:NLC

[Ham84] G. G. Hamedani. Nonnormality of linear combinations of normal random variables. The American Statistician, 38(4):295–296, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683408.

Hampton:1985:SIS

[Ham85] Angela Hampton. STAT80: An interactive statistical package. The American Statistician, 39(2):143, May 1985. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682822.

Hamburg:1987:MSK

[Ham87a] Morris Hamburg. In memoriam: Simon Kuznets, 1901–1985. The American Statistician, 41(2):93–94, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684218.

Hamilton:1987:SCV

[Ham87b] David Hamilton. Sometimes R^{2} > r^{2}yx_{1} + r^{2}yx_{2}: Correlated variables are not always redundant. The American Statistician, 41(2):129–132, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684224.

Han:1983:NNC

[Han83] Chien-Pai Han. NCSD-noncentral chi-squared distribution. The American Statistician, 37(2):173, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685887.

Han:1985:NNF

[Han85] Chien-Pai Han. NFD: Noncentral F distribution. The American Statistician, 39(3):211, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683938.

Han:1989:CTC

[Han89] Chien-Pai Han. Combining tests for correlation coefficients. The American Statistician, 43(4):211–215, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685363.

Harter:1980:MAF

[Har80] H. Leon Harter. Modified asymptotic formulas for critical values of the Kolmogorov test statistic. The American Statistician, 34(2):110–111, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684117.

Harley:1986:EPT

[Har86a] Dwight Harley. EQUORM: a program to test the homogeneity of a set of correlation matrices. The American Statistician, 40(1):51, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683120.

Harley:1986:HHF

[Har86b] Dwight Harley. HIFACT: Hierarchical factor analysis. The American Statistician, 40(4):307–308, November 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684613.

Harville:1986:UOL

[Har86c] David A. Harville. Using ordinary least squares software to compute combined intra-interblock estimates of treatment contrasts. The American Statistician, 40(2):153–157, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684879.

Harris:1988:ER

[Har88] Bernard Harris. Election recounting. The American Statistician, 42(1):66–68, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685264.

Hasselblad:1981:DPF

[Has81] Victor Hasselblad. DISFIT: a program for fitting discrete distributions to data. The American Statistician, 35(4):260, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683305.

Hastie:1987:CLD

[Has87] Trevor Hastie. A closer look at the deviance. The American Statistician, 41(1):16–20, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684312.

Hauck:1983:NCB

[Hau83] Walter W. Hauck. A note on confidence bands for the logistic response curve. The American Statistician, 37(2):158–160, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685879.

Hawkins:1980:NFR

[Haw80] Douglas M. Hawkins. A note on fitting a regression without an intercept term. The American Statistician, 34(4):233, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684067.

Hawkins:1989:USD

[Haw89] D. L. Hawkins. Using U statistics to derive the asymptotic distribution of Fisher’s Z statistic. The American Statistician, 43 (4):235–237, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685369.

Henry:1986:BRB

[HB86] Neil W. Henry and David F. Bauer. Book review: [untitled]. The American Statistician, 40(1):50–51, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683119.

Heiny:1986:BBE

[HC86] Robert L. Heiny and Carl E. Crosswhite. Best-ball events in golf: An application of the multinomial distribution. The American Statistician, 40(4):316–317, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684617.

Hosmer:1984:ITD

[HE84] Trina Hosmer and Robert Edwards. Imaging: a tool for data reduction. The American Statistician, 38(4):322–327, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683427.

Heffernan:1988:NMS

[Hef88] Peter M. Heffernan. New measures of spread and a simpler formula for the normal distribution. The American Statistician, 42(2):100– 102, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684480.

Heiberger:1981:SEDa

[Hei81a] Richard M. Heiberger. The specification of experimental designs to ANOVA programs. The American Statistician, 35(2):98–104, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683149.

Heiberger:1981:SEDb

[Hei81b] Richard M. Heiberger. [The specification of experimental designs to ANOVA programs]: Reply. The American Statistician, 35(2):108, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683154.

Heiselbetz:1986:PEP

[Hei86] Carina Heiselbetz. PLANUNG: Experimental planning facility. The American Statistician, 40(1):58, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683132.

Helms:1988:MSH

[Hel88] Ronald W. Helms. Manipulating statistical hypotheses and tests as Boolean functions. The American Statistician, 42(4):253–256, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685132.

Henstridge:1983:TTS

[Hen83] John D. Henstridge. TSA — a time series package. The American Statistician, 37(2):171–172, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685885.

Herr:1980:HUG

[Her80] David G. Herr. On the history of the use of geometry in the general linear model. The American Statistician, 34(1):43–47, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682995.

Herr:1986:HAU

[Her86] David G. Herr. On the history of ANOVA in unbalanced, factorial designs: the first 30 years. The American Statistician, 40 (4):265–270, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684597.

Habbema:1983:ISS

[HG83] J. D. F. Habbema and G. J. Gelpke. INDEP-SELECT: Selection of variables in diagnostic and prognostic problems. The American Statistician, 37(2):171, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685884.

Hines:1987:QGP

[HH87] W. G. S. Hines and R. J. O’Hara Hines. Quick graphical power-law transformation selection. The American Statistician, 41(1):21–24, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684313.

Halperin:1982:DO

[HHF82] William E. Halperin, Patricia A. Honchar, and Marilyn A. Fingerhut. Dioxin: An overview. The American Statistician, 36 (3):285–289, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683845.

Hidiroglou:1986:CSR

[Hid86] M. A. Hidiroglou. The construction of a self-representing stratum of large units in survey design. The American Statistician, 40 (1):27–31, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683113.

Higgins:1982:QSE

[Hig82] James J. Higgins. [Quality of statistical education: Should ASA assist or assess?]: Comment. The American Statistician, 36(2):94– 95, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684011.

Hill:1983:WIC

[Hil83] William J. Hill. [What industry can do for statistics]: Comment. The American Statistician, 37(1):29–30, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685839.

Hildebrand:1984:LMP

[Hil84] David K. Hildebrand. Lord’s MAD paradox and Jensen’s inequality. The American Statistician, 38(4):296–297, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683409.

Hilfiger:1985:BRB

[Hil85] John Jay Hilfiger. Book review: [untitled]. The American Statistician, 39(3):213–215, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683940.

Hill:1987:VLP

[Hil87] Bruce M. Hill. The validity of the Likelihood Principle. The American Statistician, 41(2):95–100, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684219.

Hill:1988:OIP

[Hil88] Bruce M. Hill. [Optimal information processing and Bayes’s theorem]: Comment. The American Statistician, 42(4):281–282, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685145.

Hinkins:1980:RPC

[Hin80] Susan Hinkins. RFACTOR — a program to create Rubin’s factorization when there are incomplete multivariate data. The American Statistician, 34(3):182–183, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683882.

Hintze:1982:NCC

[Hin82] Jerry L. Hintze. NUMBER CRUNCHER: a complete statistical package for microcomputers. The American Statistician, 36(4):386, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683097.

Hintze:1988:NGA

[Hin88] Jerry Hintze. NCSS-graphics: An advanced statistical graphics package for MS-DOS microcomputers. The American Statistician, 42(3):230–231, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685023.

Hjort:1988:EPN

[Hjo88] Nils Lid Hjort. The eccentric part of the noncentral chi square. The American Statistician, 42(2):130–132, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684485.

Hartigan:1984:MTR

[HK84] J. A. Hartigan and Beat Kleiner. A mosaic of television ratings. The American Statistician, 38(1):32–35, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683556.

Harley:1987:HPT

[HL87] Dwight D. Harley and R. P. Bryce Larke. HOMPRO: a program to test the homogeneity of a set of proportions. The American Statistician, 41(4):333, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684765.

Harrell:1987:LE

[HLT^{+}87] Frank E. Harrell, Jr., Kerry L. Lee, James R. Thompson, Rick L. Edgeman, Lawrence M. Leemis, Nathan Mantel, Paul D. James, R. W. Farebrother, Jerome P. Keating, Robert L. Mason, Michael E. Welch, George T. Diderrich, I. J. Good, and George Casella. Letters to the Editor. The American Statistician, 41 (1):87–92, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684337.

Heisey:1986:SIS

[HN86] Dennis Heisey and Gerard Nimis. STATISTIX: An interactive statistical analysis program for microcomputers. The American Statistician, 40(3):236–237, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684556.

Hedges:1983:RMR

[HO83] Larry V. Hedges and Ingram Olkin. Regression models in research synthesis. The American Statistician, 37(2):137–140, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685874.

Heitmann:1985:ILS

[HO85] George Heitmann and Keith Ord. An interpretation of the least squares regression surface. The American Statistician, 39(2):120– 123, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682810.

Hoaglin:1980:PP

[Hoa80] David C. Hoaglin. A Poissonness plot. The American Statistician, 34(3):146–149, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683871.

Hodges:1984:EPA

[Hod84a] Bob S. Hodges III. EXACHI: a procedure for analyzing r-by-c contingency tables with small cell frequencies and a large number of degrees of freedom. The American Statistician, 38(2):157, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683261.

Hodges:1984:LTL

[Hod84b] Bob S. Hodges III. LINCHI: Testing for linear trends in r-by-c ordinal contingency tables. The American Statistician, 38(2):157, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683260.

Hodges:1987:QSN

[Hod87] Bob S. Hodges III. QNumQ: Software for numbering questions, data processing fields, and skip patterns on questionnaires. The American Statistician, 41(4):331, November 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684762.

Hoerl:1985:RAY

[Hoe85] Roger W. Hoerl. Ridge analysis 25 years later. The American Statistician, 39(3):186–192, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683926.

Hoffer:1984:RCP

[Hof84] Alan R. Hoffer. Review of a collection of papers on the teaching of probability and statistics in an elementary school curriculum. The American Statistician, 38(3):213, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683658.

Hogg:1985:SEE

[Hog85] Robert V. Hogg. Statistical education for engineers: An initial task force report. The American Statistician, 39(3):168–175, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683924.

Hooke:1980:GPU

[Hoo80] Robert Hooke. Getting people to use statistics properly. The American Statistician, 34(1):39–42, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682994.

Horn:1983:MP

[Hor83] Paul S. Horn. A measure for peakedness. The American Statistician, 37(1):55–56, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685847.

Huba:1982:OFP

[HPB82] G. J. Huba, A. L. Palisoc, and P. M. Bentler. ORSIM2: a FOR-TRAN program for symmetric and asymmetric orthogonal rotation of canonical variates and interbattery factors. The American Statistician, 36(1):62, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684100.

Harley:1989:NNS

[HPM89] Dwight D. Harley, Trevor D. Petruk, and Philomena M. Mortensen. NONPAR: Nonparametric statistics package. The American Statistician, 43(4):272, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685381.

Hombas:1986:LE

[HRP^{+}86] V. C. Hombas, Thomas P. Ryan, Simo Putanen, Robert Weiss, Michael Lavine, Thomas Mathew, R. W. Farebrother, Jerome P. Keating, Robert L. Mason, K. S. Banerjee, H. Rodney Sharp, Pentti Vartia, Yrjő Vartia, and J. J. Rutledge. Letters to the Editor. The American Statistician, 40(2):178–182, May 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684887.

Hsieh:1981:CTM

[Hsi81] Hui-Kuang Hsieh. Chinese tally mark. The American Statistician, 35(3):174, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683999.

Hochberg:1983:MCM

[HT83] Yosef Hochberg and Ajit C. Tamhane. Multiple comparisons in a mixed model. The American Statistician, 37(4):305–307, ???? 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hirji:1989:MUE

[HTM89] Karim F. Hirji, Anastasios A. Tsiatis, and Cyrus R. Mehta. Median unbiased estimation for binary data. The American Statistician, 43(1):7–11, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685158.

Hu:1988:SMA

[Hu88] Tien-Chung Hu. A statistical method of approach to Stirling’s formula. The American Statistician, 42(3):204–205, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685005.

Hudson:1987:PSP

[Hud87] Joseph C. Hudson. Probability and statistics programs for MS-DOS. The American Statistician, 41(3):239, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685119.

Hunter:1981:PSR

[Hun81] William G. Hunter. The practice of statistics: The real world is an idea whose time has come. The American Statistician, 35(2):72–76, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683144.

Hunter:1983:SSL

[Hun83a] William G. Hunter. Statistics, science, law, and the environment. The American Statistician, 37(4):360–361, November 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683494.

Hunter:1983:WIC

[Hun83b] William G. Hunter. [What industry can do for statistics]: Comment. The American Statistician, 37(1):25–27, February 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685837.

Hunter:1988:DP

[Hun88] J. Stuart Hunter. The digidot plot. The American Statistician, 42(1):54, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685261.

Hunka:1989:MAP

[Hun89] Steve Hunka. A Macintosh APL program to calculate expected mean squares. The American Statistician, 43(4):273, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685382.

Hutt:1982:RH

[Hut82] Peter Barton Hutt. Regulatory history of DES. The American Statistician, 36(3):267, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683840.

Hawkins:1986:NTC

[HW86] Douglas M. Hawkins and R. A. J. Wixley. A note on the transformation of chi-squared variables to normality. The American Statistician, 40(4):296–298, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684608.

Ireland:1983:EGS

[IHD^{+}83] C. Terrence Ireland, Lee-Ann Hayek, Tore Dalenius, Sidney Hollander, Charles R. Mann, Eli S. Marks, Janace S. Pierce, Frederick J. Scheuren, and William Seltzer. Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics. The American Statistician, 37(1):5–6, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685820.

Iman:1982:GUL

[Ima82] Ronald L. Iman. Graphs for use with the Lilliefors test for normal and exponential distributions. The American Statistician, 36(2):109–112, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684019.

Ihm:1987:PPA

[IMS87] Peter Ihm, Hans-Georg Müller, and Thomas Schmitt. PRODOS: Probit analysis of several qualitative dose-response curves. The American Statistician, 41(1):79, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684332.

International:1985:TSP

[Int85] TSP International. Time series processor version 4.0. The American Statistician, 39(2):141, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682819.

Iman:1985:LPG

[IS85] Ronald L. Iman and Michael J. Shortencarier. LHS: a program to generate input samples for multivariate simulations. The American Statistician, 39(3):212, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683939.

Iversen:1985:SLA

[Ive85] Gudmund R. Iversen. Statistics in liberal arts education. The American Statistician, 39(1):17–19, February 1985. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683900.

Jackson:1984:JID

[Jac84] Paul H. Jackson. Justifying an invalid derivation of the probability of a union of events. The American Statistician, 38(4):291–293, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683406.

Jaynes:1988:OIP

[Jay88] E. T. Jaynes. [Optimal information processing and Bayes’s theorem]: Comment. The American Statistician, 42(4):280–281, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685144.

Joag-Dev:1989:MPM

[JD89] Kumar Joag-Dev. MAD property of a median: a simple proof. The American Statistician, 43(1):26–27, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685164.

Jabine:1983:LE

[JEE^{+}83] Thomas B. Jabine, Robert M. Eissner, Marie D. Eldridge, Benjamin H. Eichhorn, Murray Aitkin, R. Kirk Steinhorst, I. J. Good, C. Philip Cox, Lynn Roy LaMotte, J. E. Mann, Stephen M. Stigler, V. M. Joshi, Ronaldo Iachan, and C. G. Khatri. Letters to the Editor. The American Statistician, 37(1):92–97, February 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685853.

Jennings:1987:HDW

[Jen87] Dennis E. Jennings. How do we judge confidence-interval adequacy? The American Statistician, 41(4):335–337, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684768.

Jerrell:1988:CPD

[Jer88] Max E. Jerrell. Computer programs to demonstrate some hypothesis-testing issues. The American Statistician, 42(1):80–81, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685270.

Joiner:1983:URE

[JH83] Brian L. Joiner and David L. Hall. The ubiquitous role of f′/f in efficient estimation of location. The American Statistician, 37(2):128–133, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685872.

Johnson:1981:LE

[JKL^{+}81] Norman L. Johnson, Samuel Kotz, Robert F. Link, George Casella, D. J. Finney, Ronald D. Snee, Thomas J. Boardman, Gerald J. Hahn, William J. Hill, William G. Hunter, Ronald R. Hocking, William H. Lawton, R. Lyman Ott, and William E. Strawderman. Letters to the Editor. The American Statistician, 35(4):268–270, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683310.

Jacobson:1985:CGM

[JM85] Marc E. Jacobson and Charles J. Mode. A computer-generated model of human survival functions. The American Statistician, 39 (2):145, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682826.

Joshi:1989:LEP

[JN89] Suneeta M. Joshi and Smita P. Nabar. Linear estimators for the parameter in the Problem of the Nile. The American Statistician, 43(1):40–41, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2685168. See correction [Ano89h].

Jobson:1982:TZI

[Job82] J. D. Jobson. Testing for zero intercept in multivariate normal regression using the univariate multiple-regression F test. The American Statistician, 36(4):368–371, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683088.

Johnson:1981:PAA

[Joh81] M. Clemens Johnson. Pattern analysis by an adaptive classifier. The American Statistician, 35(2):110, May 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683156.

Johnson:1982:QSE

[Joh82a] Dallas E. Johnson. [Quality of statistical education: Should ASA assist or assess?]: Comment. The American Statistician, 36(2):95– 96, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684012.

Johnson:1982:RAA

[Joh82b] Edwin L. Johnson. Risk assessment in an administrative agency. The American Statistician, 36(3):232–239, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683834.

Johnson:1983:AIS

[Joh83] Carl J. Johnson. [Alternative interpretations of statistics on health effects of low-level radiation]: Comment. The American Statistician, 37(4):453–455, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683510.

Joiner:1981:LVS

[Joi81] Brian L. Joiner. Lurking variables: Some examples. The American Statistician, 35(4):227–233, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683295.

Joiner:1985:KRSa

[Joi85a] Brian L. Joiner. The key role of statisticians in the transformation of North American industry. The American Statistician, 39 (3):224–227, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/ stable/2683943. See comments [Wac85, Hah85c, God85, Boa85] and rejoinder [Joi85b].

Joiner:1985:KRSb

[Joi85b] Brian L. Joiner. [The key role of statisticians in the transformation of North American industry]: Reply. The American Statistician, 39(3):233–234, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683948.

Joglekar:1989:LFT

[JSL89] G. Joglekar, J. H. Schuenemeyer, and V. LaRiccia. Lack-of-fit testing when replicates are not available. The American Statistician, 43(3):135–143, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685059.

Jukes:1982:QEE

[Juk82] Thomas H. Jukes. A quantitative evaluation of estrogens, including DES, in the diet. The American Statistician, 36(3):273–277, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683842.

Jennings:1982:HIC

[JW82] Earl Jennings and Joe H. Ward, Jr. Hypothesis identification in the case of the missing cell. The American Statistician, 36(1):25–27, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684087.

Kahn:1987:CNB

[Kah87] William D. Kahn. A cautionary note for Bayesian estimation of the binomial parameter n. The American Statistician, 41(1):38–40, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684316.

Kaigh:1989:CRPb

[Kai89a] W. D. Kaigh. [A category representation paradox]: Reply. The American Statistician, 43(2):96–97, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684509.

Kaigh:1989:CRPa

[Kai89b] W. D. Kaigh. A category representation paradox. The American Statistician, 43(2):92–94, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684506.

Kaplan:1987:MOS

[Kap87] Harold M. Kaplan. A method of one-sided nonparametric inference for the mean of a nonnegative population. The American Statistician, 41(2):157–158, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684236.

Karakostas:1985:MVU

[Kar85] K. X. Karakostas. On minimum variance unbiased estimators. The American Statistician, 39(4):303–305, ???? 1985. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See corrections [Kar86, Kar87].

Karakostas:1986:CMV

[Kar86] K. X. Karakostas. Correction to: “On minimum variance unbiased estimators” [Amer. Statist. 39 (1985), no. 4, part 1, 303–305, MR 87b:62031]. The American Statistician, 40(3):258, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Kar85, Kar87].

Karakostas:1987:CMV

[Kar87] K. X. Karakostas. Correction: “On minimum variance unbiased estimators” [Amer. Statist. 39 (1985), no. 4, part 1, 303–305, MR 87b:62031]. The American Statistician, 41(1):92, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Kar85, Kar86].

Kass:1982:CJN

[Kas82] Robert E. Kass. A comment on “Is Jeffreys a ‘Necessarist’?”. The American Statistician, 36(4):390–391, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683099. See [Zel82a, Zel82b].

Koenker:1984:FPE

[KB84a] Roger W. Koenker and Gilbert W. Bassett. Four (pathological) examples in asymptotic statistics. The American Statistician, 38 (3):209–212, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683657.

Krewski:1984:NIE

[KB84b] Daniel Krewski and Mikelis Bickis. A note on independent and exhaustive events. The American Statistician, 38(4):290–291, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683405.

Kish:1984:ARS

[KC84] Charles W. Kish, Jr. and Walter H. Carter, Jr. An application of response surface methodology to the Atari[Registered trademark] miniature golf video game. The American Statistician, 38 (4):327–329, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683428.

Kaigh:1987:NGD

[KD87] W. D. Kaigh and Michael F. Driscoll. Numerical and graphical data summary using O-statistics. The American Statistician, 41 (1):25–32, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684314.

Kirmani:1983:NMG

[KE83] S. N. U. A. Kirmani and E. Mirhakkak Esfahani. A note on the moment generating function. The American Statistician, 37(2):161, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685880.

Kempthorne:1980:TSC

[Kem80] Oscar Kempthorne. The teaching of statistics: Content versus form. The American Statistician, 34(1):17–21, February 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682989.

Kempthorne:1983:EGS

[Kem83] Oscar Kempthorne. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):14, February 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685827.

Kempthorne:1989:EIC

[Kem89a] Oscar Kempthorne. Editor’s invited column: The fate worse than death and other curiosities and stupidities. The American Statistician, 43(3):133–134, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685058.

Kempthorne:1989:EOL

[Kem89b] Oscar Kempthorne. [The equality of the ordinary least squares estimator and the best linear unbiased estimator]: Comment. The American Statistician, 43(3):161–162, August 1989. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685063.

Kennedy:1986:ANE

[Ken86] William J. Kennedy. Announcement of new editor. The American Statistician, 40(4):259, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684595.

Kriska:1989:MUP

[KF89] S. David Kriska and Mark C. Fulcomer. MADMANager utility programs (version B-02): a powerful and portable tool kit for the management and manipulation of ASCII data files on microcomputers. The American Statistician, 43(4):273–274, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685384.

Kanal:1980:IIS

[KG80a] Laveen N. Kanal and Edzard S. Gelsema. ISPAHAN: An interactive system for pattern analysis and classification. The American Statistician, 34(3):185–186, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683886.

Knapp:1980:BRB

[KG80b] Thomas R. Knapp and Norman G. Gunderson. Book review: [untitled]. The American Statistician, 34(1):57, February 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683002.

Kodell:1984:CST

[KGHS84] Ralph L. Kodell, David W. Gaylor, M. Gary Haskin, and Gary W. Shaw. CHRONIC: An SAS[registered trademark] procedure for statistical analysis of carcinogenesis studies. The American Statistician, 38(2):158, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683262.

Kupper:1989:HAP

[KH89] Lawrence L. Kupper and Kerry B. Hafner. How appropriate are popular sample size formulas? The American Statistician, 43(2): 101–105, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684511.

Kott:1986:LE

[KHH^{+}86] Phillip S. Kott, M. A. Hidiroglou, Roger W. Hoerl, James C. Fu, Arnold Zellner, Herbert I. Weisberg, and Bradley Efron. Letters to the Editor. The American Statistician, 40(4):329–331, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684621.

Kepner:1989:NEC

[KHK89] James L. Kepner, James D. Harper, and Sandra Z. Keith. A note on evaluating a certain orthant probability. The American Statistician, 43(1):48–49, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685172.

Kinnison:1989:CCG

[Kin89] Robert Kinnison. Correlation coefficient goodness-of-fit test for the extreme-value distribution. The American Statistician, 43(2):98– 100, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684510.

Kish:1982:MRL

[Kis82] Leslie Kish. In memoriam: Rensis Likert, 1903–1981. The American Statistician, 36(2):124–125, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684023.

Kish:1983:EGS

[Kis83] Leslie Kish. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):17, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685831.

Kobialka:1989:SV

[KJ89] Edward J. Kobialka and Paul D. James. STATGRAPHICS (version 3.0). The American Statistician, 43(2):127, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684525.

Karson:1984:DUC

[KK84] Marvin J. Karson and Badrig M. Kurkjian. A derivation of a useful consequence of the Gauss–Markov theorem. The American Statistician, 38(3):202–204, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683653.

Kempthorne:1980:LE

[KKN^{+}80] Oscar Kempthorne, D. G. Kabe, K. R. Nair, Henry Wolkowicz, George P. H. Styan, James R. Lackritz, Samuel Leinhardt, Stanley Wasserman, Libert Ehrman, Robert J. K. Jacob, William C. Shelton, Stephen E. Fienberg, Ruma Falk, Gary G. Koch, and John Neter. Letters to the Editor. The American Statistician, 34(4):249–254, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684078.

Kruskal:1983:IPB

[KL83] J. B. Kruskal and J. M. Landwehr. Icicle plots: Better displays for hierarchical clustering. The American Statistician, 37(2):162–168, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685881.

Krotki:1988:LE

[KLC^{+}88] Karol J. Krotki, Roderick J. A. Little, Ronald Christensen, W. A. Thompson, Jr., I. J. Good, Martin Rutsch, H. N. Nagaraja, Irwin D. Bross, Fred L. Ramsey, Hanspeter Thoni, Ludo M. J. Gypen, and Joseph Lee Rodgers. Letters to the Editor. The American Statistician, 42(4):288–291, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685150.

Kademan:1989:IVF

[KLV89] Edmund Kademan, Wei-Yin Loh, and Nunta Vanichsetakul. An improved version of FACT with S front end. The American Statistician, 43(4):273, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685383.

Keating:1985:PRA

[KM85] Jerome P. Keating and Robert L. Mason. Practical relevance of an alternative criterion in estimation. The American Statistician, 39(3):203–205, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683931.

Katz:1987:QVR

[KM87] Richard W. Katz and Allan H. Murphy. Quality/value relationship for imperfect information in the umbrella problem. The American Statistician, 41(3):187–189, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685099.

Keating:1988:JSE

[KM88] Jerome P. Keating and Robert L. Mason. James–Stein estimation from an alternative perspective. The American Statistician, 42(2):160–164, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684494.

Keller-McNulty:1989:BRB

[KM89a] Sallie Keller-McNulty. Book review: [untitled]. The American Statistician, 43(2):123–124, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684518.

Kruskal:1989:CRI

[KM89b] William Kruskal and Ruth Majors. Concepts of relative importance in recent scientific literature. The American Statistician, 43(1):2–6, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685157.

Keller-McNulty:1987:IPA

[KMM87] Sallie Keller-McNulty and Mark McNulty. The independent pairs assumption in hypothesis tests based on rank correlation coefficients. The American Statistician, 41(1):40–41, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684317.

Knapp:1982:QSE

[Kna82] Thomas R. Knapp. [Quality of statistical education: Should ASA assist or assess?]: Comment. The American Statistician, 36(2):97, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684013.

Koch:1985:BDR

[Koc85] Gary G. Koch. A basic demonstration of the [−1, 1] range for the correlation coefficient. The American Statistician, 39(3):201–202, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683930.

Koehn:1981:RDE

[Koe81] Uwe Koehn. Review of the design of experiments video cassette course. The American Statistician, 35(1):48, February 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683586.

Koenker:1986:BRB

[Koe86] Roger Koenker. Book review: [untitled]. The American Statistician, 40(3):223–225, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684542.

Koehn:1987:BRB

[Koe87] Uwe Koehn. Book review: [untitled]. The American Statistician, 41(3):221, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685111.

Konijn:1987:DFO

[Kon87] Hendrik S. Konijn. Distribution-free and other prediction intervals. The American Statistician, 41(1):11–15, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684311.

Korn:1984:RLV

[Kor84] Edward L. Korn. The ranges of limiting values of some partial correlations under conditional independence. The American Statistician, 38(1):61–62, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683563.

Kott:1986:WMR

[Kot86] Phillip S. Kott. When a mean-of-ratios is the best linear unbiased estimator under a model. The American Statistician, 40 (3):202–204, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684536.

Kenyon:1987:BRB

[KP87] J. R. Kenyon and H. O. Posten. Book review: [untitled]. The American Statistician, 41(1):58–59, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684323.

Kraemer:1988:AAG

[Kra88] Helena Chmura Kraemer. Assessment of 2×2 associations: Generalization of signal-detection methodology. The American Statistician, 42(1):37–49, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685259.

Kruskal:1987:RIA

[Kru87] William Kruskal. Relative importance by averaging over orderings. The American Statistician, 41(1):6–10, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684310.

Kabak:1983:LDL

[KS83] Irwin W. Kabak and Jeffrey S. Simonoff. A look at daily lotteries. The American Statistician, 37(1):49–52, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685845.

Kotz:1988:HSA

[KS88] Samuel Kotz and Kathleen Smith. The Hausdorff space and applied statistics: a view from the U.S.S.R. The American Statistician, 42(4):241–244, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685129.

Knapp:1981:BRB

[KSB^{+}81] Thomas R. Knapp, David B. Swanson, Austin J. Bonis, Albert D. Rickmers, Mason E. Wescott, and Mitchell J. Mergenthaler. Book review: [untitled]. The American Statistician, 35(1):50–51, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683588.

Kullback:1988:OIP

[Kul88] Solomon Kullback. [Optimal information processing and Bayes’s theorem]: Comment. The American Statistician, 42(4):282–283, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685147.

Lackritz:1984:EPV

[Lac84] James R. Lackritz. Exact p values for F and t tests. The American Statistician, 38(4):312–314, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683414.

Lachenbruch:1985:BRB

[Lac85] Peter A. Lachenbruch. Book review: [untitled]. The American Statistician, 39(2):146–148, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682828.

Lentner:1989:EBH

[LAH89] Marvin Lentner, Jesse C. Arnold, and Klaus Hinkelmann. The efficiency of blocking: How to use MS(Blocks)/MS(Error) correctly. The American Statistician, 43(2):106–108, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684512.

Lamm:1980:CPL

[Lam80] R. McFall Lamm, Jr. A computer program for linear systems analysis: LSAP. The American Statistician, 34(2):116–117, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684121.

Lam:1983:PMP

[Lam83] K. Lam. Probabilistic modeling-from physics to computing science. The American Statistician, 37(3):213–215, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683372.

Landau:1985:P

[Lan85] Emanuel Landau. Preface. The American Statistician, 39(4):403, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683020.

Lave:1982:LB

[Lav82] Lester B. Lave. Lessons from benzene. The American Statistician, 36(3):260–261, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683837.

Lenth:1988:LE

[LBW^{+}88] Russell V. Lenth, Sean Becketti, Peter Westfall, Thomas M. Davis, Donald A. Berry, S. Mitra, Rudolf J. Freund, David C. Hamilton, George Casella, and William M. Alley. Letters to the Editor. The American Statistician, 42(1):87–91, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685275.

Leemis:1986:RAC

[Lee86] Lawrence M. Leemis. Relationships among common univariate distributions. The American Statistician, 40(2):143–146, May 1986.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684876.

Lehmann:1980:ELE

[Leh80] E. L. Lehmann. Efficient likelihood estimators. The American Statistician, 34(4):233–235, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684068.

Lehnen:1988:SPS

[Leh88] Robert G. Lehnen. Statistical policy for state and local governments. The American Statistician, 42(1):10–16, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685253.

Lenth:1983:SPS

[Len83] Russell V. Lenth. Some properties of U statistics. The American Statistician, 37(4):311–313, ???? 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:1987:PMP

[Len87] Russell V. Lenth. PowerPack: a microcomputer package for power and sample-size problems. The American Statistician, 41(3):239– 240, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685120.

Leone:1982:QSE

[Leo82] Fred C. Leone. [Quality of statistical education: Should ASA assist or assess?]: Comment. The American Statistician, 36(2):100–101, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684016.

Leone:1983:WIC

[Leo83] Fred C. Leone. [What industry can do for statistics]: Comment. The American Statistician, 37(1):30–31, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685840.

LePage:1987:BRB

[LeP87] Raoul LePage. Book review: [untitled]. The American Statistician, 41(1):64–67, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684325.

Levy:1980:NAS

[Lev80] Kenneth J. Levy. Nonparametric applications of Shaffer’s extension of Dunnett’s procedure. The American Statistician, 34(2):99–102, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684113.

Levy:1985:NNM

[Lev85] Martin S. Levy. A note on nonunique MLE’s and sufficient statistics. The American Statistician, 39(1):66, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683913.

Levenbach:1987:TSF

[Lev87a] Hans Levenbach. 4CaST/2: a time series forecasting and data analysis system. The American Statistician, 41(1):77, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684328.

Levine:1987:BRB

[Lev87b] David M. Levine. Book review: [untitled]. The American Statistician, 41(2):139–142, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684228.

Leviton:1987:PCM

[Lev87c] Laura C. Leviton. PREVIEW: a computer model to project cost of lifestyle-related illnesses and savings from worksite health promotion. The American Statistician, 41(4):332–333, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684764.

Lewis:1980:AEA

[Lew80] Stephen D. Lewis. APLECON-econometric analysis with APL. The American Statistician, 34(3):184, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683884.

Looney:1985:UCC

[LG85] Stephen W. Looney and Thomas R. Gulledge, Jr. Use of the correlation coefficient with normal probability plots. The American Statistician, 39(1):75–79, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683917.

Lunde:1981:IAA

[LH81] Anders S. Lunde and William G. Hunter. International activities and the American Statistical Association. The American Statistician, 35(3):117–120, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683973.

Liem:1980:CPB

[Lie80] Tran Cong Liem. A computer program for Box–Cox transformations in regression models with heteroscedastic and autoregressive residuals. The American Statistician, 34(2):121, May 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684125.

Liggett:1985:BRB

[Lig85] Walter Liggett. Book review: [untitled]. The American Statistician, 39(1):70–72, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683915.

Liggett:1988:BRB

[Lig88] Walter Liggett. Book review: [untitled]. The American Statistician, 42(2):155–157, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684492.

Lindstrom:1980:SOF

[Lin80] F. Tom Lindstrom. SUREXPL: An objective function surface exploration algorithm. The American Statistician, 34(3):183, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683883.

Link:1981:MDP

[Lin81] Robert F. Link. Moments of discrete probability distributions derived using finite difference operators. The American Statistician, 35(1):44–46, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683584.

Lin:1985:MNC

[Lin85a] Shang P. Lin. MORPOL: Numerical construction of multidimensional orthogonal polynomials. The American Statistician, 39(2): 142, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682820.

Ling:1985:BRB

[Lin85b] Robert F. Ling. Book review: [untitled]. The American Statistician, 39(3):215–219, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683941.

Lin:1986:SCP

[Lin86a] Shang P. Lin. SPCTL: Computing percentiles of individual survival function in Cox proportional hazard model. The American Statistician, 40(2):170, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684883.

Lindley:1986:WIE

[Lin86b] D. V. Lindley. [Why isn’t everyone a bayesian?]: Comment. The American Statistician, 40(1):6–7, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683107.

Ling:1987:BRB

[Lin87] Robert F. Ling. Book review: [untitled]. The American Statistician, 41(1):71–76, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684327.

Ling:1988:BRB

[Lin88] Robert F. Ling. Book review: [untitled]. The American Statistician, 42(3):220–223, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685011.

Little:1982:DST

[Lit82] Roderick J. A. Little. Direct standardization: a tool for teaching linear models for unbalanced data. The American Statistician, 36(1):38–43, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684091.

Little:1988:SSA

[Lit88] Roderick J. A. Little. Some statistical analysis issues at the world fertility survey. The American Statistician, 42(1):31–36, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685258.

Little:1989:TET

[Lit89] Roderick J. A. Little. Testing the equality of two independent binomial proportions. The American Statistician, 43(4):283–288, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685390.

Los:1986:EMP

[LK86] Cornelis A. Los and Christopher McC. Kell. EMEPS: Microcomputer programs for the estimation and prediction of time-varying parameters and other latent variables. The American Statistician, 40(4):306–307, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684612.

Long:1986:PSC

[LM86] Susan B. Long and Hubert B. Mueller. Programming SAS for causal analysis. The American Statistician, 40(3):233–234, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684552.

Lee:1984:SSP

[LMM84] Alan J. Lee, Peter McInerney, and Peter R. Mullins. STATCALC: a statistical package for the Apple IIe and II Plus computers. The American Statistician, 38(4):321, November 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683425.

Libby:1981:CAD

[LNC^{+}81] David L. Libby, Melvin R. Novick, James J. Chen, George G. Woodworth, and Robert M. Hamer. The computer-assisted data analysis (CADA) monitor (1980). The American Statistician, 35 (3):165–166, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683996.

Lodge:1985:MDE

[Lod85] James P. Lodge. Materials damage by environmental pollutants: Data requirements. The American Statistician, 39(4):412–415, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683022.

Louis:1981:CIB

[Lou81] Thomas A. Louis. Confidence intervals for a binomial parameter after observing no successes. The American Statistician, 35(3):154, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683985.

Love:1983:BAA

[Lov83] Richard R. Love. Biological aspects of associations between environmental exposures and cancer. The American Statistician, 37(4):413–419, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683504.

Lovell:1984:LDM

[Lov84] Michael C. Lovell. Lovell’s data Miner. The American Statistician, 38(3):225, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683664.

Lowerre:1982:IMM

[Low82] James M. Lowerre. An introduction to modern matrix methods and statistics. The American Statistician, 36(2):113–115, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684020.

Lowerre:1983:IBN

[Low83] James M. Lowerre. An integral of the bivariate normal and an application. The American Statistician, 37(3):235–236, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683380.

Loyer:1983:BPG

[Loy83] Milton W. Loyer. Bad probability, good statistics, and group testing for binomial estimation. The American Statistician, 37 (1):57–59, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685848.

Lehmann:1982:JN

[LR82] E. L. Lehmann and Constance Reid. Jerzy Neyman, 1894–1981. The American Statistician, 36(3):161–162, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Little:1983:JEP

[LR83] Roderick J. A. Little and Donald B. Rubin. On jointly estimating parameters and missing data by maximizing the complete-data likelihood. The American Statistician, 37(3):218–220, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683374.

Lehmann:1988:ID

[LS88] E. L. Lehmann and Juliet Popper Shaffer. Inverted distributions. The American Statistician, 42(3):191–194, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684999.

Looney:1989:ERM

[LS89] Stephen W. Looney and William B. Stanley. Exploratory repeated measures analysis for two or more groups: Review and update. The American Statistician, 43(4):220–225, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685365.

Levin:1989:AVT

[LSWB89] Joel R. Levin, Ronald C. Serlin, and Lisa Webne-Behrman. Analysis of variance through simple correlation. The American Statistician, 43(1):32–34, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685166.

Lund:1980:MIS

[Lun80] Richard E. Lund. MSUSTAT-an interactive statistical analysis package. The American Statistician, 34(1):58–59, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683003.

Lund:1984:MIS

[Lun84] Richard E. Lund. MSUSTAT: An interactive statistical analysis package. The American Statistician, 38(4):317, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683418.

Lunneborg:1988:BAB

[Lun88] C. E. Lunneborg. Bootstrap applications for the behavioral sciences. The American Statistician, 42(1):86, February 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685274.

Lin:1989:END

[LV89] Lawrence I-Kuei Lin and Edward F. Vonesh. An empirical nonlinear data-fitting approach for transforming data to normality. The American Statistician, 43(4):237–243, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685370.

Lusk:1982:DPD

[LW82] Edward J. Lusk and Haviland Wright. Deriving the probability density for sums of uniform random variables. The American Statistician, 36(2):128–130, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684025.

Milliken:1984:NSC

[MA84] George A. Milliken and Mohammed Albohali. On necessary and sufficient conditions for ordinary least squares estimators to be best linear unbiased estimators. The American Statistician, 38 (4):298–299, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683411.

Macdonald:1986:MIP

[Mac86] Peter D. M. Macdonald. MIX: An interactive program for fitting mixtures of distributions. The American Statistician, 40(1):53, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683123.

Mage:1982:OGM

[Mag82] David T. Mage. An objective graphical method for testing normal distributional assumptions using probability plots. The American Statistician, 36(2):116–120, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684021.

Makridakis:1981:STT

[Mak81] Spyros Makridakis. SIBYL/TFS: The total forecasting system. The American Statistician, 35(2):113, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683160.

Mantel:1981:CSW

[Man81] Nathan Mantel. Calculation of scores for a Wilcoxon generalization applicable to data subject to arbitrary right censorship. The American Statistician, 35(4):244–247, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683298.

Mandel:1982:USV

[Man82] John Mandel. Use of the singular value decomposition in regression analysis. The American Statistician, 36(1):15–24, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684086.

Mantel:1983:OAT

[Man83] Nathan Mantel. Ordered alternatives and the 1 1/2-tail test. The American Statistician, 37(3):225–228, August 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683376.

Mantel:1987:UWT

[Man87] Nathan Mantel. Understanding Wald’s test for exponential families. The American Statistician, 41(2):147–148, May 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684232.

Mann:1988:BRB

[Man88] Barbara L. Mann. Book review: [untitled]. The American Statistician, 42(4):270–271, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685136.

Marquardt:1981:CEP

[Mar81a] Donald W. Marquardt. Criteria for evaluating the performance of statistical consultants in industry. The American Statistician, 35(4):216–219, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683293.

Martin:1981:IFS

[Mar81b] Margaret E. Martin. [Improving the federal statistical system: Report of the President’s reorganization project for the federal statistical system]: Comment. The American Statistician, 35(4): 197–199, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683287.

Markey:1982:LIR

[Mar82] Howard T. Markey. The legal implications of risk. The American Statistician, 36(3):256–259, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683836.

Margolis:1983:RVS

[Mar83a] Marvin S. Margolis. Review of videotape short course for business statistics. The American Statistician, 37(3):240–241, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683385.

Martin:1983:SDAa

[Mar83b] James A. Martin. Science and democracy in the age of technology: Separating fact from value. The American Statistician, 37(4):367–371, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683496.

Martin:1983:SDAb

[Mar83c] James A. Martin. [Science and democracy in the age of technology: Separating fact from value]: Response. The American Statistician, 37(4):373, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683498.

Martin:1983:EGS

[Mar83d] Margaret E. Martin. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):7–8, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685821.

Marquardt:1984:NTE

[Mar84] Donald W. Marquardt. New technical and educational directions for managing product quality. The American Statistician, 38(1):8–14, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683550.

Marasinghe:1985:BRB

[Mar85] Mervyn G. Marasinghe. Book review: [untitled]. The American Statistician, 39(1):72–74, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683916.

Masse:1989:PSS

[Mas89] Jean-Yves Masse. PONC: Software on statistical analysis concerning a spatial point process. The American Statistician, 43 (4):274–275, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685385.

Musser-Brauner:1984:BBW

[MBD84] Melonia L. Musser-Brauner and W. J. Dixon. BMD or BMDP? What’s the difference? The American Statistician, 38(4):318–319, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683421.

Monlezun:1984:CSPa

[MBM84a] Charles J. Monlezun, David C. Blouin, and Linda C. Malone. Contrasting split plot and repeated measures experiments and analyses. The American Statistician, 38(1):21–27, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683552.

Monlezun:1984:CSPb

[MBM84b] Charles J. Monlezun, David C. Blouin, and Linda C. Malone. [Contrasting split plot and repeated measures experiments and analyses]: Reply. The American Statistician, 38(1):30–31, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683555.

McCulloch:1985:EAE

[MBM^{+}85] Charles E. McCulloch, Daniel R. Boroto, Duane Meeter, Ronald Polland, and Douglas A. Zahn. An expanded approach to educating statistical consultants. The American Statistician, 39(3):159–167, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683923.

Mansfield:1987:DVR

[MC87] Edward R. Mansfield and Michael D. Conerly. Diagnostic value of residual and partial residual plots. The American Statistician, 41(2):107–116, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684221.

McBride:1980:BRB

[McB80] James McBride. Book review: [untitled]. The American Statistician, 34(2):113–114, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684119.

McCabe:1980:IRA

[McC80] George P. McCabe, Jr. The interpretation of regression analysis results in sex and race discrimination problems. The American Statistician, 34(4):212–215, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684062.

McCulloch:1988:ILF

[McC88] Robert E. McCulloch. Information and the likelihood function in exponential families. The American Statistician, 42(1):73–75, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685266.

McDonald:1983:WIC

[McD83] Gary C. McDonald. [What industry can do for statistics]: Comment. The American Statistician, 37(1):32–33, February 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685842.

Mode:1987:EEC

[MCJ87] Charles J. Mode, Marc D. Cain, and Marc E. Jacobson. ECPSES: Estimating critical population size in the presence of environmental stochasticity. The American Statistician, 41(4):334, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684767.

Meeker:1981:CUO

[MD81] William Q. Meeker, Jr. and Steven D. Duke. CENSOR — a useroriented computer program for life data analysis. The American Statistician, 35(2):112, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683159.

Meeks:1983:NUC

[MD83] S. L. Meeks and R. B. D’Agostino. A note on the use of confidence limits following rejection of a null hypothesis. The American Statistician, 37(2):134–136, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685873.

Mahmoud:1985:ISP

[MD85] Essam Mahmoud and Timothy A. Davidson. Interactive statistical programs: a comprehensive tool for teaching statistics. The American Statistician, 39(2):140, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682817.

Morris:1984:IPS

[ME84] Max D. Morris and Sherwood F. Ebey. An interesting property of the sample mean under a first-order autoregressive model. The American Statistician, 38(2):127–129, May 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683248.

Meeden:1986:SPC

[Mee86] Glen Meeden. Sufficiency and partitions of the class of all possible discrete distributions. The American Statistician, 40(1):42–44, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683116.

Meeden:1987:EWU

[Mee87] Glen Meeden. Estimation when using a statistic that is not sufficient. The American Statistician, 41(2):135–136, May 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684226.

Marsh:1986:OPU

[MEP^{+}86] Gary M. Marsh, Jeffrey Ehland, Myunghee Paik, Mary Preininger, and Richard Caplan. OCMAP/PC: a user-oriented occupational cohort mortality analysis program for the IBM PC. The American Statistician, 40(4):308–309, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684614.

Mergerson:1988:APM

[Mer88] James W. Mergerson. ALLOC.P: a multivariate allocation program. The American Statistician, 42(1):85, February 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685272.

Mergerson:1989:GUO

[Mer89] James W. Mergerson. A generalized univariate optimal allocation program. The American Statistician, 43(2):128, May 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684527.

Martin:1984:RAT

[MFG^{+}84] Margaret E. Martin, Martin R. Frankel, Noreen Goldman, Daniel G. Horvitz, Joseph B. Kadane, Graham Kalton, Samuel H. Preston, Bruce D. Spencer, Michael A. Stoto, and Franklin D. Wilson. Report of the ASA technical panel on the census undercount. The American Statistician, 38(4):252–256, November 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683398.

Martin:1982:CNS

[MGS82] Margaret E. Martin, Edwin D. Goldfield, and Miron L. Straf. The Committee on National Statistics: 10 years later. The American Statistician, 36(2):103–108, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684018.

Mohammad:1985:AEM

[MGS85] W. A. Mohammad, M. Grossman, and R. D. Shanks. Algebraic equivalence of matrix inversion, elimination, and absorption for use in animal breeding. The American Statistician, 39(2):124–125, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682812.

Mantel:1987:LE

[MHP^{+}87] Nathan Mantel, P. S. Hewlett, R. L. Plackett, T. T. Nguyen, S. H. Chou, Bernhard Flury, Robert J. Myers, Samuel M. Slowinski, Alfred C. Hexter, M. Adrian, Calvin F. Schmid, Irwin D. Bross, Steven K. Seilkop, Gerald Akland, Richard J. Cook, and Michael R. Stoline. Letters to the Editor. The American Statistician, 41(2):159–165, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684237.

Miller:1984:RTB

[Mil84] Don M. Miller. Reducing transformation bias in curve fitting. The American Statistician, 38(2):124–126, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683247.

Millard:1987:EMS

[Mil87] Steven P. Millard. Environmental monitoring, statistics, and the law: Room for improvement. The American Statistician, 41(4): 249–253, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684739.

Milliken:1989:BRB

[Mil89] George A. Milliken. Book review: [untitled]. The American Statistician, 43(3):174–175, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685069.

Minton:1980:PSC

[Min80] Paul D. Minton. [Preparing statisticians for careers in industry: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Industry]: Comment. The American Statistician, 34(2):77–78, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684108.

McLeod:1984:NDS

[MJ84] A. Ian McLeod and Carlos Jimenez. Nonnegative definiteness of the sample autocovariance function. The American Statistician, 38(4):297–298, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683410.

Moulton:1986:LE

[MKF^{+}86] Lawrence H. Moulton, Leon Katz, R. W. Farebrother, Irwin D. Bross, D. V. Lindley, John T. Shelton, Victor W. Lowe, Jr., and George A. Milliken. Letters to the Editor. The American Statistician, 40(1):77–79, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683138.

Mazumdar:1980:CBL

[MLB80] Sati Mazumdar, Ching Chun Li, and G. Rex Bryce. Correspondence between a linear restriction and a generalized inverse in linear model analysis. The American Statistician, 34(2):103–105, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684114.

Monroe:1980:GMC

[MM80] Robert J. Monroe and Francis E. McVay. Gertrude Mary Cox (1900–1978). The American Statistician, 34(1):48, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682996.

Miller:1984:MUO

[MM84] Barbara Miller and Sati Mazumdar. MVSPEC: a user-oriented multivariate spectral analysis program. The American Statistician, 38(4):319, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683422.

McPeek:1982:MJP

[MMM82] Bucknam McPeek, Cornelia McPeek, and Frederick Mosteller. In memoriam: John Parker Gilbert, (1926-1980). The American Statistician, 36(1):37, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684090.

Mallows:1987:UUE

[MN87] Colin L. Mallows and Vijayan N. Nair. A unique unbiased estimator with an interesting property. The American Statistician, 41 (3):205–206, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685105.

Manoukian:1988:NHS

[MN88] Edward B. Manoukian and Pierre Nadeau. A note on the hyperbolic-secant distribution. The American Statistician, 42(1): 77–79, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685268.

Moore:1984:ASG

[MO84] David S. Moore and Ingram Olkin. Academic statistics: Growth, change, and federal support. The American Statistician, 38(1):1–7, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683549.

Moore:1981:RSV

[Moo81] David S. Moore. Review of selected volumes of Quantitative Applications in the Social Sciences. The American Statistician, 35 (3):160–161, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683990.

Moors:1986:MKD

[Moo86] J. J. A. Moors. The meaning of kurtosis: Darlington reexamined. The American Statistician, 40(4):283–284, November 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684603.

Morris:1986:WIE

[Mor86] C. N. Morris. [Why isn’t everyone a bayesian?]: Comment. The American Statistician, 40(1):7–8, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683108.

Morey:1987:BRB

[Mor87a] Mathew J. Morey. Book review: [untitled]. The American Statistician, 41(2):143–145, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684229.

Morganstein:1987:BRB

[Mor87b] David Morganstein. Book review: [untitled]. The American Statistician, 41(4):318–321, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684754.

Mosteller:1980:CPP

[Mos80] Frederick Mosteller. Classroom and platform performance. The American Statistician, 34(1):11–17, February 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682988.

Moses:1981:IFS

[Mos81] Lincoln E. Moses. [Improving the federal statistical system: Report of the President’s reorganization project for the federal statistical system]: Comment. The American Statistician, 35(4):203–204, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683289.

Moses:1982:PSC

[Mos82] Lincoln E. Moses. [Preparing statisticians for careers in the federal government: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Government]: Comment. The American Statistician, 36(2):82, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684005.

Mosteller:1983:EGS

[Mos83] Frederick Mosteller. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):10–11, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685824.

Mosteller:1988:BSS

[Mos88] Frederick Mosteller. Broadening the scope of statistics and statistical education. The American Statistician, 42(2):93–99, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684479.

Marsh:1980:OUO

[MP80] Gary M. Marsh and Mary Preininger. OCMAP: a user-oriented occupational cohort mortality analysis program. The American Statistician, 34(4):245–246, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684074.

Mode:1988:CMR

[MP88] Charles J. Mode and Gary T. Pickens. Computational methods for renewal theory and semi-Markov processes with illustrative examples. The American Statistician, 42(2):143–152, May 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684490.

Mixon:1988:PDP

[MPH88] Wilson Mixon, R. Prince, and J. Hoftyzer. P-DIST: a package to calculate binomial probabilities. The American Statistician, 42(3):231, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685024.

Moore:1989:SLA

[MR89] Thomas L. Moore and Rosemary A. Roberts. Statistics at liberal arts colleges. The American Statistician, 43(2):80–85, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684503.

Marsh:1989:OME

[MRP^{+}89] Gary M. Marsh, B. Raja Rao, Mary Preininger, Jeffrey Ehland, and Harvey T. CoChien. OCMAP (Module 6): Extended proportional mortality analysis with simultaneous inferential procedures. The American Statistician, 43(2):127–128, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684526.

Meinhold:1983:UKF

[MS83] Richard J. Meinhold and Nozer D. Singpurwalla. Understanding the Kalman filter. The American Statistician, 37(2):123–127, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685871.

Meinhold:1987:KFS

[MS87] Richard J. Meinhold and Nozer D. Singpurwalla. A Kalmanfilter smoothing approach for extrapolations in certain doseresponse, damage-assessment, and accelerated-life-testing studies. The American Statistician, 41(2):101–106, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684220.

Misra:1987:BTP

[MSGG87] S. Chandra Misra, Hardeo Sahai, Anil P. Gore, and Joseph K. Garrett. A bibliography on the teaching of probability and statistics. The American Statistician, 41(4):284–310, November 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684751.

Mosteller:1981:EFS

[MSTY81] Frederick Mosteller, Andrew F. Siegel, Edward Trapido, and Cleo Youtz. Eye fitting straight lines. The American Statistician, 35 (3):150–152, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683983.

Melnick:1982:MND

[MT82] Edward L. Melnick and Aaron Tenenbein. Misspecifications of the normal distribution. The American Statistician, 36(4):372–373, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683089.

Muller:1980:ASC

[Mul80] Mervin E. Muller. Aspects of statistical computing: What packages for the 1980’s ought to do. The American Statistician, 34 (3):159–168, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683875.

Muller:1982:UCC

[Mul82] Keith E. Muller. Understanding canonical correlation through the general linear model and principal components. The American Statistician, 36(4):342–354, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683082.

Mullooly:1988:VLT

[Mul88] John P. Mullooly. The variance of left-truncated continuous nonnegative distributions. The American Statistician, 42(3):208–210, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685007.

Murray:1986:EMC

[Mur86] Leigh W. Murray. Estimation of missing cells in randomized block and Latin square designs. The American Statistician, 40 (4):289–293, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684606.

Murray:1988:MCE

[Mur88] Leigh W. Murray. Missing-cell estimates in the randomized block design with random effects. The American Statistician, 42(2):134– 136, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684487.

Norton:1985:TM

[NA85] Robert M. Norton and Steven Arnold. A theorem on moments. The American Statistician, 39(2):106, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682806.

Nagel:1981:PEZ

[Nag81] P. J. A. Nagel. Programs for the evaluation of zonal polynomials. The American Statistician, 35(1):53, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683590.

Naiman:1987:CRS

[Nai87] Daniel Q. Naiman. Characterization of risk sets for simple versus simple hypothesis testing. The American Statistician, 41(3):218– 220, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685109.

Najarian:1983:AIS

[Naj83] Thomas Najarian. [Alternative interpretations of statistics on health effects of low-level radiation]: Comment. The American Statistician, 37(4):455–457, November 1983. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683511.

Nash:1986:BRBa

[Nas86a] John C. Nash. Book review: [untitled]. The American Statistician, 40(4):301–303, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684610.

Nash:1986:BRBb

[Nas86b] John C. Nash. Book review: [untitled]. The American Statistician, 40(4):303–306, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684611.

Noether:1982:LE

[NBA^{+}82] Gottfried E. Noether, George A. Barnard, J. Aitchison, J. H. Bacon-Shone, Nathan Mantel, James R. Lackritz, and Jerry A. Warren. Letters to the Editor. The American Statistician, 36 (4):394–396, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683101.

Neal:1982:IRP

[Nea82] Robert A. Neal. Introductory remarks to the panel on 2,4,5-T/dioxin. The American Statistician, 36(3):284, August 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683844.

Neffendorf:1983:SPM

[Nef83] Hugh Neffendorf. Statistical packages for microcomputers: a listing. The American Statistician, 37(1):83–86, February 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685851.

Nelson:1981:RRE

[Nel81] Lloyd S. Nelson. Review of the reliability engineering videotape course. The American Statistician, 35(1):49–50, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683587.

Nelder:1985:AIS

[Nel85] John A. Nelder. An alternative interpretation of the singularvalue decomposition in regression. The American Statistician, 39 (1):63–64, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683911.

Neumann:1985:RSM

[Neu85] Charles J. Neumann. The role of statistical models in the prediction of tropical cyclone motion. The American Statistician, 39(4):347–357, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683013.

Nash:1980:LE

[NMF^{+}80] John C. Nash, Nathan Mantel, Camil Fuchs, Robert J. Buehler, Philip Cox, and John Neter. Letters to the Editor. The American Statistician, 34(3):190–192, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683892.

Nie:1982:MECb

[NN82a] Norman H. Nie and Marija H. Norusis. [More on evaluating computer programs]: Reply. The American Statistician, 36(2):142, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684036.

Nie:1982:MECa

[NN82b] Norman H. Nie and Marija J. Norusis. More on evaluating computer programs. The American Statistician, 36(2):141, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684034.

NAG:1986:NRG

[NN86] Numerical Algorithms Group Ltd. and Numerical Algorithms Group Inc. The new release of GLIM: GLIM 3.77. The American Statistician, 40(3):232, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684550.

Noether:1980:RNI

[Noe80] Gottfried E. Noether. The role of nonparametrics in introductory statistics courses. The American Statistician, 34(1):22–23, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682990.

Noether:1981:RTB

[Noe81] Gottfried E. Noether. [Rank transformations as a bridge between parametric and nonparametric statistics]: Comment. The American Statistician, 35(3):129–130, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683976.

Noether:1984:NEY

[Noe84] Gottfried E. Noether. Nonparametrics: The early years — impressions and recollections. The American Statistician, 38(3):173–178, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683647.

Norton:1984:DED

[Nor84] Robert M. Norton. The double exponential distribution: Using calculus to find a maximum likelihood estimator. The American Statistician, 38(2):135–136, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683252.

North:1987:USU

[Nor87] Philip M. North. U-SP: a user-friendly survey analysis package. The American Statistician, 41(1):78, February 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684331.

Norton:1989:PCA

[Nor89] Robert M. Norton. Pocket-calculator approximation for areas under the standard normal curve. The American Statistician, 43 (1):24–26, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685163.

Novick:1984:MBI

[Nov84] Melvin R. Novick. Multiparameter Bayesian indifference procedures revisited: Comment. The American Statistician, 38(4): 248–249, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683395.

Nisbet:1982:SAT

[NP82] Ian C. T. Nisbet and Mary Burr Paxton. Statistical aspects of three-generation studies of the reproductive toxicity of TCDD and 2,4,5-T. The American Statistician, 36(3):290–298, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683846.

Norusis:1980:SCS

[NW80] Marija J. Norusis and Chih-Ming Wang. The SCSS conversational system. The American Statistician, 34(4):247–248, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684076.

Norusis:1981:SED

[NW81] Marija J. Norusis and Chih-Ming Wang. [The specification of experimental designs to ANOVA programs]: Comment. The American Statistician, 35(2):106–107, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683152.

Nedelman:1986:BTP

[NW86] Jerry Nedelman and Ted Wallenius. Bernoulli trials, Poisson trials, surprising variances, and Jensen’s inequality. The American Statistician, 40(4):286–289, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684605.

Nelson:1988:SIP

[NZ88] Dean E. Nelson and Gary O. Zerbe. A SAS/IML program to execute randomization analysis of response curves with multiple comparisons. The American Statistician, 42(3):231–232, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685025.

OBrien:1988:BRB

[O’B88] Ralph G. O’Brien. Book review: [untitled]. The American Statistician, 42(4):266–270, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685135.

Ott:1985:CMO

[ONC85] M. Gerald Ott, S. K. Norwood, and R. R. Cook. The collection and management of occupational exposure data. The American Statistician, 39(4):432–436, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683026.

Ord:1984:MMG

[Ord84] Keith Ord. In memoriam: Maurice George Kendall, 1907–1983. The American Statistician, 38(1):36–37, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683557.

Orris:1980:MIS

[Orr80] J. Burdeane Orris. MICROSTAT: An interactive statistical package for microcomputers. The American Statistician, 34(3):188–189, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683890.

Ott:1983:WIC

[Ott83] R. Lyman Ott. [What industry can do for statistics]: Comment. The American Statistician, 37(1):27–28, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685838.

Owen:1982:QSE

[Owe82] D. B. Owen. [Quality of statistical education: Should ASA assist or assess?]: Comment. The American Statistician, 36(2):97–99, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684014.

Paik:1985:GRT

[Pai85] Minja Paik. A graphic representation of a three-way contingency table: Simpson’s paradox and correlation. The American Statistician, 39(1):53–54, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683907.

Parzen:1980:PSC

[Par80] Emanuel Parzen. [Preparing statisticians for careers in industry: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Industry]: Comment. The American Statistician, 34(2):78–79, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684109.

Park:1982:RA

[Par82] Colin N. Park. Risk assessment for 2,4,5-T. The American Statistician, 36(3):299–300, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683847.

Parks:1989:BRB

[Par89] Robert B. Parks. Book review: [untitled]. The American Statistician, 43(1):60–63, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685176.

Patil:1985:P

[Pat85] G. P. Patil. Preface. The American Statistician, 39(4):361–362, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683015.

Paul:1981:BTU

[Pau81] Jerome L. Paul. A Bayes theorem urn model optimization problem. The American Statistician, 35(4):247–249, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683299.

Paul:1985:EPC

[Pau85] S. R. Paul. Estimation of the parameters of a clumped binomial model via the EM algorithm. The American Statistician, 39(2):136–139, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682816.

Pavur:1988:TER

[Pav88] Robert Pavur. Type I error rates for multiple comparison procedures with dependent data. The American Statistician, 42(3):171– 173, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684994.

Payne:1981:SED

[Pay81] R. W. Payne. [The specification of experimental designs to ANOVA programs]: Comment. The American Statistician, 35(2):107, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683153.

Pickens:1984:MIP

[PBM84] Gary T. Pickens, Robert C. Busby, and Charles J. Mode. MATH-IST: An interactive program for modeling maternity histories. The American Statistician, 38(3):224–225, August 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683663.

Preston:1983:SPI

[PC83] Dale L. Preston and Douglas B. Clarkson. SURVREG: a program for the interactive analysis of survival regression models. The American Statistician, 37(2):174, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685889.

Piegorsch:1985:EFN

[PC85] Walter W. Piegorsch and George Casella. The existence of the first negative moment. The American Statistician, 39(1):60–62, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683910.

Peck:1980:PPE

[Pec80] Jon K. Peck. PEC: Program for econometric computation. The American Statistician, 34(3):189, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683891.

Peddada:1985:SNP

[Ped85] Shyamal Das Peddada. A short note on Pitman’s measure of nearness. The American Statistician, 39(4):298–299, ???? 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Peixoto:1987:HVS

[Pei87] Julio L. Peixoto. Hierarchical variable selection in polynomial regression models. The American Statistician, 41(4):311–313, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684752.

Pentico:1981:DUO

[Pen81] David W. Pentico. On the determination and use of optimal sample sizes for estimating the difference in means. The American Statistician, 35(1):40–42, February 1981. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683582.

Pentico:1988:BBN

[Pen88] David W. Pentico. A box of balls: New lessons from an old model. The American Statistician, 42(3):197–198, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685001.

Perlman:1984:UCD

[Per84] Gary Perlman. UNIX—STAT: Compact data analysis package. The American Statistician, 38(2):155, May 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683258.

Pugh:1989:BRB

[PF89] Marian Pugh and David Findley. Book review: [untitled]. The American Statistician, 43(1):63–66, February 1989. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685177.

Plane:1982:SPN

[PG82] Donald R. Plane and Kenneth R. Gordon. A simple proof of the nonapplicability of the central limit theorem to finite populations. The American Statistician, 36(3):175–176, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pelz:1983:CKS

[PG83] Wolfgang Pelz and I. J. Good. Calculating the Kolmogorov– Smirnov one-sample statistic. The American Statistician, 37(2): 173, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685888.

Perlman:1986:UDA

[PH86] Gary Perlman and Frederick L. Horan. UNIX—STAT 5.2: Data analysis programs for MS–DOS and UNIX. The American Statistician, 40(3):229–230, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684546.

Pollack:1983:P

[PHA83] Alison K. Pollack, William G. Hunter, and Gordon J. Apple. Preface. The American Statistician, 37(4):359, November 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683493.

Pollack:1984:HAU

[PHA84] Alison K. Pollack, William G. Hunter, and Gordon J. Apple. [How to avoid undue influence of federal agencies on the editorial content of ASA journals]: Reply. The American Statistician, 38(3):229, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683667.

Pierce:1980:SRD

[Pie80] David A. Pierce. A survey of recent developments in seasonal adjustment. The American Statistician, 34(3):125–134, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683868.

Pike:1985:UMD

[Pik85] Arthur C. Pike. The use of meteorological data in the prediction of tropical cyclones. The American Statistician, 39(4):335–346, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683012.

Pinault:1988:ASR

[Pin88] Steven C. Pinault. An analysis of subset regression for orthogonal designs. The American Statistician, 42(4):275–277, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685142.

Pittenger:1980:SPF

[Pit80] Donald B. Pittenger. Some problems in forecasting population for government planning purposes. The American Statistician, 34 (3):135–139, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683869.

Peterson:1982:MMC

[PK82] Arthur V. Peterson, Jr. and Richard A. Kronmal. On mixture methods for the computer generation of random variables. The American Statistician, 36(3):184–191, August 1982. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pinsker:1986:RFP

[PKG86] I. Sh. Pinsker, Victor Kipnis, and Eugene Grechanovsky. A recursive formula for the probability of occurrence of at least m out of N events. The American Statistician, 40(4):275–276, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684599.

Platt:1986:BRB

[Pla86] W. Gerald Platt. Book review: [untitled]. The American Statistician, 40(2):167–169, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684882.

Paik:1987:OPR

[PN87] Minja Paik and Elliott Nebenzahl. The overall percentile rank versus the individual percentile ranks. The American Statistician, 41(2):136–138, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684227.

Polhemus:1980:AAF

[Pol80] N. W. Polhemus. ACFPLOT: An APL function for graphical display of cross-correlation matrices. The American Statistician, 34 (3):180–181, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683879.

Polhemus:1982:STI

[Pol82] Neil W. Polhemus. STATGRAPHICS TM: An interactive statistical graphics system in APL. The American Statistician, 36(4):385, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683095.

Poole:1980:ETR

[Poo80] W. Kenneth Poole. The equivalence of two rules of classification for two populations. The American Statistician, 34(1):55, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683000.

Posten:1980:BRB

[Pos80] Harry O. Posten. Book review: [untitled]. The American Statistician, 34(4):236, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684070.

Posten:1981:RST

[Pos81] Harry O. Posten. Review of statistical teaching materials for 11– 16-year-olds. The American Statistician, 35(4):258–259, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683303.

Posten:1982:BRB

[Pos82] Harry O. Posten. Book review: Probability Activity Books for Elementary and Junior High School Students. The American Statistician, 36(2):135–136, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684030.

Posten:1983:BRB

[Pos83] Harry O. Posten. Book review: Transparency Set for Basic Statistics. The American Statistician, 37(3):239, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683383.

Posten:1989:EAN

[Pos89] Harry O. Posten. An effective algorithm for the noncentral chisquared distribution function. The American Statistician, 43(4): 261–263, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685374.

Press:1986:WIE

[Pre86] S. James Press. [Why isn’t everyone a bayesian?]: Comment. The American Statistician, 40(1):9–10, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683109.

Price:1987:EMS

[Pri87] Bertram Price. [Environmental monitoring, statistics, and the law: Room for improvement]: Comment. The American Statistician, 41(4):257–259, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684741.

Potter:1981:PJT

[PS81] R. W. Potter and G. W. Sturm. The power of Jonckheere’s test. The American Statistician, 35(4):249–250, November 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683300.

Pearl:1982:MRF

[PS82] Robert Pearl and Seymour Sudman. In memoriam: Robert Ferber, 1922–1981. The American Statistician, 36(4):355, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683083.

Park:1983:QRA

[PS83] Colin N. Park and Ronald D. Snee. Quantitative risk assessment: State-of-the-art for carcinogenesis. The American Statistician, 37 (4):427–441, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683506.

Puntanen:1989:EOLa

[PS89a] Simo Puntanen and George P. H. Styan. The equality of the ordinary least squares estimator and the best linear unbiased estimator. With comments by Oscar Kempthorne and Shayle R. Searle and a reply by the authors. The American Statistician, 43(3):153–161, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685062.

Puntanen:1989:EOLb

[PS89b] Simo Puntanen and George P. H. Styan. [The equality of the ordinary least squares estimator and the best linear unbiased estimator]: Reply. The American Statistician, 43(3):164, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685065.

Potthoff:1983:LE

[PSD^{+}83] Richard F. Potthoff, Michael Stuart, Larry Dion, Donald Fridshal, Edward Melnick, Aaron Tenenbein, L. Pirktl, Michael Goldstein, Michael F. Fuller, Gerald S. Rogers, Laurence Gillick, J. B. Garner, Hans Zeisel, Nathan Gale, and Bernard W. Lindgren. Letters to the Editor. The American Statistician, 37(3):251–254, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683388.

Pohl:1980:CAC

[PT80] Norval F. Pohl and San-Yun W. Tsai. A computerized aid for calculating the power of chi-squared tests. The American Statistician, 34(3):184–185, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683885.

Raghavachari:1986:LPP

[Rag86] M. Raghavachari. A linear programming proof of Kantorovich’s inequality. The American Statistician, 40(2):136–137, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684872.

Rahlfs:1981:DIP

[Rah81] Volker W. Rahlfs. DIALOG: An interactive program package with EDA techniques for microcomputers. The American Statistician, 35(3):162–163, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683991.

Raiffa:1982:SPT

[Rai82] Howard Raiffa. Science and policy: Their separation and integration in risk analysis. The American Statistician, 36(3):225–231, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683833.

Ramsey:1986:FP

[Ram86] Fred L. Ramsey. A fable of PCA. The American Statistician, 40(4):323–324, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684619.

Raveh:1985:UIC

[Rav85] Adi Raveh. On the use of the inverse of the correlation matrix in multivariate data analysis. The American Statistician, 39(1):39–42, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683904.

Raveh:1986:MMA

[Rav86] Adi Raveh. On measures of monotone association. The American Statistician, 40(2):117–123, May 1986. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684869.

Rayner:1985:MLE

[Ray85] J. C. W. Rayner. Maximum likelihood estimation of μ and Σ from a multivariate normal distribution. The American Statistician, 39(2):123–124, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682811.

Rouanet:1986:NSI

[RBL86] Henry Rouanet, Jean-Marc Bernard, and Bruno Lecoutre. Nonprobabilistic statistical inference: a set-theoretic approach. The American Statistician, 40(1):60–65, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683134.

Rode:1988:UBC

[RC88] Richard A. Rode and Vernon M. Chinchilli. The use of Box–Cox transformations in the development of multivariate tolerance regions with applications to clinical chemistry. The American Statistician, 42(1):23–30, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685257.

Reid:1988:APC

[RD88] John G. Reid and Michael F. Driscoll. An accessible proof of Craig’s theorem in the noncentral case. The American Statistician, 42(2):139–142, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684489.

Read:1985:APS

[Rea85] K. L. Q. Read. ANOVA problems with simple numbers. The American Statistician, 39(2):107–111, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682807.

Regal:1986:BRB

[Reg86] Ronald R. Regal. Book review: [untitled]. The American Statistician, 40(2):164–167, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684881.

Reilly:1980:PGC

[Rei80] David P. Reilly. PACK: a generalized computer program for singleequation Box–Jenkins models. The American Statistician, 34(3): 181, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2683880.

Reilly:1984:ABA

[Rei84] David P. Reilly. AUTOBJ, BOXX, AUTOBOX, and SIMULA-TOR: Smart software for time series data analysis. The American Statistician, 38(4):315, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683415.

Reilly:1987:ABA

[Rei87] David P. Reilly. AUTOBOX, BOXX, AUTOBJ, MTS, and SIM-BOXJ: Smart software for time series analysis and forecasting. The American Statistician, 41(2):145–146, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684230.

Relles:1985:DDA

[Rel85] Daniel A. Relles. dasy: Data analysis system. The American Statistician, 39(3):210, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683935.

Remington:1983:EGS

[Rem83] Richard D. Remington. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):14–15, February 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685828.

Reynolds:1988:SPS

[Rey88] Reid T. Reynolds. [Statistical policy for state and local governments]: Comment. The American Statistician, 42(1):18–20, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685255.

Rickard:1988:NCE

[REZ88] Russel S. Rickard, Richard M. Engeman, and Gary O. Zerbe. Nonparametric comparison of exponential growth curves. The American Statistician, 42(3):232, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685026.

Rice:1983:EGS

[Ric83] Dorothy P. Rice. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):9, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685823.

Rivest:1981:SCN

[Riv81] Louis-Paul Rivest. On the sum of contaminated normals. The American Statistician, 35(3):155–156, August 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683986.

Rao:1982:MMD

[RJ82] B. Raja Rao and K. G. Janardan. On the moments of multivariate discrete distributions using finite difference operators. The American Statistician, 36(4):381–383, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683093.

Romesburg:1980:CCC

[RM80] H. Charles Romesburg and Kim Marshall. CLUSTAR and CLUSTID: Computer programs for hierarchical cluster analysis. The American Statistician, 34(3):186, August 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683887.

Romesburg:1981:LCP

[RM81] H. Charles Romesburg and Kim Marshall. LIFE: a computer program for stochastic life table analysis. The American Statistician, 35(2):114, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683161.

Ruberg:1988:IPA

[RM88] Stephen J. Ruberg and Robert L. Mason. Increasing public awareness of statistics as a science and a profession-starting in the high schools. The American Statistician, 42(3):167–170, August 1988.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684993.

Rodgers:1988:TWL

[RN88] Joseph Lee Rodgers and W. Alan Nicewander. Thirteen ways to look at the correlation coefficient. The American Statistician, 42 (1):59–66, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685263.

Rodgers:1984:LIO

[RNT84] Joseph Lee Rodgers, W. Alan Nicewander, and Larry Toothaker. Linearly independent, orthogonal, and uncorrelated variables. The American Statistician, 38(2):133–134, May 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683250.

Roberts:1983:EGS

[Rob83a] Harry V. Roberts. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):18, February 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685833.

Robinson:1983:RLB

[Rob83b] Raymond M. Robinson. The rule of law between nations: An acid test. The American Statistician, 37(4):377–384, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683500.

Roberts:1987:DAM

[Rob87] Harry V. Roberts. Data analysis for managers. The American Statistician, 41(4):270–278, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684747.

Rockower:1988:IIR

[Roc88] Edward B. Rockower. Integral identities for random variables. The American Statistician, 42(1):68–72, February 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685265.

Rogers:1984:KPA

[Rog84] Gerald S. Rogers. Kronecker products in ANOVA — a first step. The American Statistician, 38(3):197–202, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683652.

Rohlf:1984:BPS

[Roh84] F. James Rohlf. BIOM: a package of statistical computer programs. The American Statistician, 38(4):318, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683420.

Rohlf:1987:NPM

[Roh87] F. James Rohlf. NTSYS-pc: Microcomputer programs for numerical taxonomy and multivariate analysis. The American Statistician, 41(4):330, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684761.

Romesburg:1987:CPI

[Rom87] H. Charles Romesburg. CLUSTAR-PC: Interactive program for hierarchical cluster analysis. The American Statistician, 41(1):79, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684333.

Rosenbaum:1981:TSP

[Ros81a] Paul R. Rosenbaum. The two-scale plot: An exploratory display of data with heterogeneous variances. The American Statistician, 35(4):265–266, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683309.

Rosenberg:1981:RCI

[Ros81b] Harry M. Rosenberg. A report and comments on improving the federal statistical system: Overview. The American Statistician, 35(4):183, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683285.

Ross:1987:EMS

[Ros87] N. Phillip Ross. [Environmental monitoring, statistics, and the law: Room for improvement]: Comment. The American Statistician, 41(4):254–256, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684740.

Rosenbaum:1989:EPP

[Ros89] Paul R. Rosenbaum. Exploratory plots for paired data. The American Statistician, 43(2):108–109, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684513.

Royall:1986:ESS

[Roy86] Richard M. Royall. The effect of sample size on the meaning of significance tests. The American Statistician, 40(4):313–315, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684616.

Read:1983:SPS

[RR83] K. L. Q. Read and I. S. Riley. Statistics problems with simple numbers. The American Statistician, 37(3):229–231, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683377.

Rosenbaum:1984:SBI

[RR84] Paul R. Rosenbaum and Donald B. Rubin. Sensitivity of Bayes inference with data-dependent stopping rules. The American Statistician, 38(2):106–109, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683243.

Rosenbaum:1985:CCG

[RR85] Paul R. Rosenbaum and Donald B. Rubin. Constructing a control group using multivariate matched sampling methods that incorporate the propensity score. The American Statistician, 39(1):33–38, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683903.

Rich:1984:LE

[RSI^{+}84] Herbert Rich, Marvin A. Schneiderman, C. Terrence Ireland, Calvin Berry, Colin P. Hargreaves, A. S. C. Ehrenberg, Peter M. Neely, Nathan Mantel, Philippe Kent, Joseph B. Kruskal, and James M. Landwehr. Letters to the Editor. The American Statistician, 38(2):160–163, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683265.

Ranney:1981:SCD

[RT81] G. B. Ranney and C. C. Thigpen. The sample coefficient of determination in simple linear regression. The American Statistician, 35(3):152–153, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683984.

Rueter:1985:CSP

[Rue85] Frederick H. Rueter. Conceptual and statistical problems in identifying thresholds of effect for materials damage and soiling. The American Statistician, 39(4):404–411, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683021.

Ruppert:1987:WKI

[Rup87] David Ruppert. What is kurtosis?: An influence function approach. The American Statistician, 41(1):1–5, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684309.

Russell:1985:Pa

[Rus85a] Clifford S. Russell. Preface. The American Statistician, 39(4):333, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683010.

Russell:1985:Pb

[Rus85b] Clifford S. Russell. Preface. The American Statistician, 39(4):334, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683011.

Ryan:1982:MWR

[Rya82] Walter F. Ryan. In memoriam: William Ramsdell Leonard, 1904– 1981. The American Statistician, 36(2):122–123, May 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684022.

Suissa:1987:GLR

[SA87] Samy Suissa and Jeanne Adam. Generalized linear regression for discrete data using SAS. The American Statistician, 41(3):241, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685122.

Salsburg:1985:RSP

[Sal85a] David S. Salsburg. The religion of statistics as practiced in medical journals. The American Statistician, 39(3):220–223, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683942.

Salvia:1985:SQC

[Sal85b] Anthony A. Salvia. Statistical quality control in the first statistics course. The American Statistician, 39(3):205–207, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683932.

Sall:1988:SII

[Sal88] John Sall. SAS/IML: An interactive matrix language. The American Statistician, 42(3):232–233, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685027.

Saunders:1984:SPI

[Sau84a] Lucy Saunders. SPSS/PC: The information analysis system for the IBM PC. The American Statistician, 38(4):320, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683424.

Saunders:1984:SPP

[Sau84b] Lucy Saunders. SPSS/Pro: a package for the DEC Professional 350. The American Statistician, 38(4):320, November 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683423.

Savage:1982:LLA

[Sav82] L. James Savage. Log-linear analysis: An APL workspace enhanced to compute formal variables and expected odds ratios. The American Statistician, 36(1):59–60, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684096.

Sawyer:1984:FPD

[Saw84] John W. Sawyer, Jr. First partial differentiation by computer with an application to categorical data analysis. The American Statistician, 38(4):300–308, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683412.

Seber:1983:LE

[SB83] G. A. F. Seber and A. Buse. Letters to the Editor. The American Statistician, 37(2):176, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685891.

Sims:1985:NHR

[SB85] John H. Sims and Duane D. Baumann. Natural hazard research and policy: Time for a gadfly. The American Statistician, 39 (4):358–360, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683014.

Scott:1985:OTM

[SBA85] Del T. Scott, G. Rex Bryce, and David M. Allen. Orthogonalizationtriangularization methods in statistical computations. The American Statistician, 39(2):128–135, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682815.

Stegman:1982:LE

[SBE^{+}82] Charles E. Stegman, David Bellhouse, A. S. C. Ehrenberg, Nathan Mantel, Frank Proschan, Daniel Gianola, S. R. Searle, F. M. Speed, and G. A. Milliken. Letters to the Editor. The American Statistician, 36(1):64–67, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684102.

Samuel-Cahn:1981:CTS

[SC81] Ester Samuel-Cahn. Comparison of three sampling designs. The American Statistician, 35(3):156–157, August 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683987.

Stoline:1986:SSA

[SC86] Michael R. Stoline and Richard J. Cook. A study of statistical aspects of the Love Canal Environmental Monitoring Study. The American Statistician, 40(2):172–177, May 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684886.

Schilling:1984:LE

[SCC^{+}84] Mark F. Schilling, Guy Chatillon, C. Philip Cox, H. J. Keselman, Hanspeter Thoni, Charles J. Monlezun, David C. Blouin, and Linda C. Malone. Letters to the Editor. The American Statistician, 38(4):330–332, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683429.

Schervish:1981:STC

[Sch81a] Mark J. Schervish. Simultaneous tests and confidence intervals. The American Statistician, 35(3):175, August 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684000.

Schmitt:1981:EHS

[Sch81b] Robert C. Schmitt. Early Hawaiian statistics. The American Statistician, 35(1):1–3, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683575.

Schmoyer:1984:EAC

[Sch84a] Richard L. Schmoyer. Everyday application of the cell means model. The American Statistician, 38(1):49–52, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683559.

Schultz:1984:SSM

[Sch84b] Mark Schultz. SSI/SSF: Multivariate statistical forecasting made easy. The American Statistician, 38(4):316, November 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683416.

Schwager:1984:BSL

[Sch84c] Steven J. Schwager. Bonferroni sometimes loses. The American Statistician, 38(3):192–197, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683651.

Schervish:1985:BRB

[Sch85a] Mark J. Schervish. Book review: [untitled]. The American Statistician, 39(1):67–70, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683914.

Schey:1985:GDO

[Sch85b] Harry M. Schey. A geometric description of orthogonal contrasts in one-way analysis of variance. The American Statistician, 39(2):104–106, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682805.

Schmid:1986:WHS

[Sch86] Calvin F. Schmid. Whatever has happened to the semilogarithmic chart? The American Statistician, 40(3):238–244, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684558.

Schoenberg:1988:LLC

[Sch88a] Ronald Schoenberg. LINCS: Linear covariance structure analysis for the microcomputer. The American Statistician, 42(3):233, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685028.

Schwarz:1988:ABP

[Sch88b] W. Schwarz. Approximating the birthday problem. The American Statistician, 42(3):195–196, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685000.

Schmittlein:1989:SIU

[Sch89a] David C. Schmittlein. Surprising inferences from unsurprising observations: Do conditional expectations really regress to the mean? The American Statistician, 43(3):176–183, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685070.

Schucany:1989:CRP

[Sch89b] William R. Schucany. [A category representation paradox]: Comment. The American Statistician, 43(2):94–95, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684507.

Street:1988:NCR

[SCR88] James O. Street, Raymond J. Carroll, and David Ruppert. A note on computing robust regression estimates via iteratively reweighted least squares. The American Statistician, 42(2):152–154, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684491. See cor-

rection [Ano89e].

Scheers:1986:RCP

[SD86] N. J. Scheers and C. Mitchell Dayton. RRCOV: Computer program for covariate randomized response models. The American Statistician, 40(3):229, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684545.

Scariano:1987:EVI

[SD87] Stephen M. Scariano and James M. Davenport. The effects of violations of independence assumptions in the one-way ANOVA. The American Statistician, 41(2):123–129, May 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684223.

Strijbosch:1989:CPE

[SD89] Leo Strijbosch and Ronald J. M. M. Does. Computer programs for the experimental design and statistical evaluation of limiting dilution assays. The American Statistician, 43(4):275–276, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685387.

Searle:1981:ACO

[Sea81] S. R. Searle. Annotated computer output for variance components. The American Statistician, 35(2):111–112, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683158.

Searle:1984:RGI

[Sea84] S. R. Searle. Restrictions and generalized inverses in linear models. The American Statistician, 38(1):53–54, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683560.

Searle:1988:ACO

[Sea88a] S. R. Searle. Annotated computer output, second edition: ACO_{2}. The American Statistician, 42(4):272, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685138.

Searle:1988:PLR

[Sea88b] Shayle R. Searle. Parallel lines in residual plots. The American Statistician, 42(3):211, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685008.

Searle:1989:EIC

[Sea89a] Shayle R. Searle. Editor’s invited column: Statistical computing packages: Some words of caution. The American Statistician, 43(4):189–190, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685359.

Searle:1989:EOL

[Sea89b] Shayle R. Searle. [The equality of the ordinary least squares estimator and the best linear unbiased estimator]: Comment. The American Statistician, 43(3):162–163, August 1989. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685064.

Seigel:1989:SDC

[Sei89] Daniel Seigel. Should data be collected on a dependent variable late in a clinical trial if not measured at baseline? The American Statistician, 43(3):184–185, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685071.

Searle:1980:CGD

[SF80] S. R. Searle and P. A. Firey. Computer generation of data sets for homework exercises in simple regression. The American Statistician, 34(1):51–54, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2682999.

Shapiro:1982:CCT

[Sha82] Stanley H. Shapiro. Collapsing contingency tables — a geometric approach. The American Statistician, 36(1):43–46, February 1982.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684092.

Shah:1984:SSD

[Sha84] B. V. Shah. Software for survey data analysis. The American Statistician, 38(1):68–69, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683566.

Shah:1985:SAA

[Sha85] Arvind K. Shah. A simpler approximation for areas under the standard normal curve. The American Statistician, 39(1):80, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683918.

Shahabuddin:1986:ENE

[Sha86a] Syed Shahabuddin. EXPOSMOO: a new exponential smoothing computer package. The American Statistician, 40(2):170–171, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684884.

Shapiro:1986:IAH

[Sha86b] Arthur Shapiro. Interactive ad hoc segmented straightening transformations. The American Statistician, 40(1):47–49, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683118.

Sheu:1984:PAB

[She84] Shey Shiung Sheu. The Poisson approximation to the binomial distribution. The American Statistician, 38(3):206–207, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683655.

Searle:1983:ACO

[SHF83] S. R. Searle, G. F. S. Hudson, and W. T. Federer. Annotated computer output for analysis of covariance. The American Statistician, 37(2):172–173, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685886.

Shier:1988:MPM

[Shi88a] D. R. Shier. The monotonicity of power means using entropy. The American Statistician, 42(3):203–204, August 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685004.

Shiffler:1988:MSO

[Shi88b] Ronald E. Shiffler. Maximum Z scores and outliers. The American Statistician, 42(1):79–80, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685269.

Shrout:1987:BRB

[Shr87] Patrick E. Shrout. Book review: [untitled]. The American Statistician, 41(4):314–317, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684753.

Shuster:1982:NOS

[Shu82] Jonathan J. Shuster. Nonparametric optimality of the sample mean and sample variance. The American Statistician, 36(3):176–178, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shulte:1983:RMR

[Shu83] Albert P. Shulte. Review of mathematics resources project material in statistics. The American Statistician, 37(3):238–239, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683382.

Shumway:1986:BRB

[Shu86] R. H. Shumway. Book review: [untitled]. The American Statistician, 40(4):299–301, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684609.

Shuster:1988:ECE

[Shu88] Jonathan J. Shuster. EXACTB and CONF: Exact unconditional procedures for binomial data. The American Statistician, 42(3): 234, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685029.

Shvyrkov:1984:WMP

[Shv84] Vladislav V. Shvyrkov. We must preach what should be practiced. The American Statistician, 38(2):151, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683255.

Smith:1982:ECT

[SI82] T. M. Smith and B. Iglewicz. An effective classroom technique for comparison of robust estimators. The American Statistician, 36(3):163–168, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schafer:1986:MSP

[SJ86] William D. Schafer and Charles E. Johnson. A microcomputer statistics package for instructional support. The American Statistician, 40(3):231, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684548.

Stanek:1985:EPE

[SK85] Edward J. Stanek III and Gary G. Koch. The equivalence of parameter estimates from growth curve models and seemingly unrelated regression models. The American Statistician, 39(2):149–152, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682829.

Slater:1983:FSC

[Sla83] Courtenay Slater. [Federal statistical coordination today: a disaster or a disgrace?]: Comment. The American Statistician, 37 (3):197–198, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683368.

Smith:1983:STE

[SM83] David W. Smith and L. W. Murray. A simplified treatment of the estimation of parameters and tests of hypotheses in constrained design models with unbalanced data. The American Statistician, 37(2):156–158, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685878.

Snee:1984:CCD

[SM84] Ronald D. Snee and Donald W. Marquardt. Comment: Collinearity diagnostics depend on the domain of prediction, the model, and the data. The American Statistician, 38(2):83–87, May 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683239.

Smith:1981:MHO

[Smi81] W. B. Smith. In memoriam: Herman Otto Hartley, (1912-1980). The American Statistician, 35(3):142–143, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683980.

Smith:1984:LTF

[Smi84] Charles W. Smith. LOKATE TM: a file for determining longitudelatitude map location using only ZIP code and alpha place name. The American Statistician, 38(2):159, May 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683264.

Smith:1985:PFD

[Smi85] William P. Smith. PC FIT: a distribution fitting program. The American Statistician, 39(2):141, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682818.

Smith:1986:WIE

[Smi86] Adrian F. M. Smith. [Why isn’t everyone a bayesian?]: Comment. The American Statistician, 40(1):10, February 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683110.

Snee:1982:PSC

[Sne82] Ronald D. Snee. [Preparing statisticians for careers in the federal government: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Government]: Comment. The American Statistician, 36(2):82–87, May 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684006.

Snee:1984:CBU

[Sne84] Ronald D. Snee. Cooperation between university and industry statisticians. The American Statistician, 38(1):15–20, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683551.

Solomon:1983:EGS

[Sol83] Herbert Solomon. [Ethical guidelines for statistical practice: Report of the Ad Hoc Committee on Professional Ethics]: Comment. The American Statistician, 37(1):15, February 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685829.

Somes:1986:GMH

[Som86] Grant W. Somes. The generalized Mantel–Haenszel statistic. The American Statistician, 40(2):106–108, May 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684866.

Southward:1984:NLA

[SP84] G. Morris Southward and Harry O. Posten. A new listing of audiovisual materials for statistical education. The American Statistician, 38(2):110–116, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683244.

Searle:1985:EPS

[SP85a] Shayle R. Searle and Friedrich Pukelsheim. Establishing χ^{2} properties of sums of squares using induction. The American Statistician, 39(4):301–303, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Southward:1985:ANL

[SP85b] G. Morris Southward and Harry O. Posten. Addendum to a new listing of audio-visual materials for statistical education. The American Statistician, 39(3):208, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683933.

Searle:1986:EIC

[SP86] Shayle R. Searle and Friedrich Pukelsheim. Effects of intraclass correlation on weighted averages. The American Statistician, 40(2): 103–105, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2684865.

Sahrmann:1987:SOH

[SPC87] Herman F. Sahrmann, Gregory F. Piepel, and John A. Cornell. In search of the optimum Harvey Wallbanger recipe via mixture experiment techniques. The American Statistician, 41(3):190–194, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2685100.

Spiegelman:1985:JIG

[Spi85] C. H. Spiegelman. Jensen’s inequality for general location parameter. The American Statistician, 39(1):54, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683908.

Spiegelman:1986:TPU

[Spi86] Clifford H. Spiegelman. Two pitfalls of using standard regression diagnostics when both X and Y have measurement error. The American Statistician, 40(3):245–248, August 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684559. See correction [Spi87].

Spiegelman:1987:CTP

[Spi87] Clifford H. Spiegelman. Correction: “Two pitfalls of using standard regression diagnostics when both X and Y have measurement error” [amer. statist. 40 (1986), no. 3, 245–248]. The American Statistician, 41(1):92, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Spi86].

Silva:1983:EPU

[SR83] Luis Carlos Silva and Antonio Rubi. On the estimation of proportions using complete cluster sampling. The American Statistician, 37(1):53–54, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685846.

Suissa:1984:UMP

[SS84] Samy Suissa and Jonathan J. Shuster. Are uniformly most powerful unbiased tests really best. The American Statistician, 38 (3):204–206, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683654.

Schader:1989:TRT

[SS89] Martin Schader and Friedrich Schmid. Two rules of thumb for the approximation of the binomial distribution by the normal distribution. The American Statistician, 43(1):23–24, February 1989.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685162.

Searle:1981:SCM

[SSH81] S. R. Searle, F. M. Speed, and H. V. Henderson. Some computational and model equivalences in analyses of variance of unequalsubclass-numbers data. The American Statistician, 35(1):16–33, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683578.

Searle:1980:PMM

[SSM80] S. R. Searle, F. M. Speed, and G. A. Milliken. Population marginal means in the linear model: an alternative to least squares means. The American Statistician, 34(4):216–221, November 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684063.

Stanish:1983:EIC

[ST83] William M. Stanish and Noel Taylor. Estimation of the intraclass correlation coefficient for the analysis of covariance model. The American Statistician, 37(3):221–224, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683375.

SCI:1986:PNS

[Sta86] Statistical Consultants, Inc. PCNONLIN and NONLIN84: Software for the statistical analysis of nonlinear models. The American Statistician, 40(1):52, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683122.

Stanek:1988:CPP

[Sta88] Edward J. Stanek III. Choosing a pretest-posttest analysis. The American Statistician, 42(3):178–183, August 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684996.

Steinhorst:1982:RCC

[Ste82] R. K. Steinhorst. Resolving current controversies in analysis of variance. The American Statistician, 36(2):138–139, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684032.

Stevens:1985:SPM

[Ste85] Robert D. Stevens. Statistics programs for microcomputers: Directory, descriptions, and evaluations. The American Statistician, 39(2):142–143, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682821.

Stephenson:1986:BRB

[Ste86] W. R. Stephenson. Book review: [untitled]. The American Statistician, 40(1):45–46, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683117.

Stenson:1987:TPA

[Ste87] Herbert H. Stenson. TESTAT: Psychometric analysis of test data. The American Statistician, 41(2):146, May 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684231.

Steinberg:1988:PLT

[Ste88a] Dan Steinberg. PROBIT, LOGIT, TOBIT, and 2SLS: Supplements for SYSTAT. The American Statistician, 42(3):234–235, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685030.

Steinhorst:1988:BRB

[Ste88b] R. Kirk Steinhorst. Book review: [untitled]. The American Statistician, 42(1):82–84, February 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685271.

Steinberg:1989:SSM

[Ste89a] Dan Steinberg. Survival: a supplementary module for SYSTAT. The American Statistician, 43(1):67, February 1989. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2685179.

Stenson:1989:SSM

[Ste89b] Herbert H. Stenson. SIGNAL: a supplementary module for SYS-TAT. The American Statistician, 43(4):275, November 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685386.

Stine:1980:EDA

[Sti80] Robert A. Stine. An exploratory data analysis package. The American Statistician, 34(3):187–188, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683889.

Stigler:1982:MPN

[Sti82] Stephen M. Stigler. A modest proposal: a new standard for the normal. The American Statistician, 36(2):137–138, May 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684031.

Stigler:1984:KPJ

[Sti84] Stephen M. Stigler. Kruskal’s proof of the joint distribution of X̄ and s^{2}. The American Statistician, 38(2):134–135, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683251.

Stigler:1986:ESC

[Sti86] Stephen M. Stigler. Estimating serial correlation by visual inspection of diagnostic plots. The American Statistician, 40(2):111–116, May 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684868.

Stotland:1980:CTA

[Sto80] Victor G. Stotland. Create table: Automatic composition of statistical tables. The American Statistician, 34(4):244–245, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684073.

Stoline:1981:SMC

[Sto81] Michael R. Stoline. The status of multiple comparisons: Simultaneous estimation of all pairwise comparisons in one-way ANOVA designs. The American Statistician, 35(3):134–141, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683979.

Stolarski:1982:FSO

[Sto82] R. S. Stolarski. Fluorocarbons and stratospheric ozone: a review of current knowledge. The American Statistician, 36(3):303–311, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683849.

Stoto:1988:DUS

[Sto88] Michael A. Stoto. Dealing with uncertainty: Statistics for an aging population. The American Statistician, 42(2):103–110, May 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684481.

Strauss:1980:SGP

[Str80] Stanley H. Strauss. STATPAC: a general purpose package for data analysis and fitting statistical models to data. The American Statistician, 34(1):59–60, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683005.

Straf:1982:P

[Str82a] Miron L. Straf. Preface. The American Statistician, 36(3):221, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683831.

Straf:1982:RAE

[Str82b] Miron L. Straf. Risk assessment in environmental decision making. The American Statistician, 36(3):222–224, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683832.

Stuart:1982:GAP

[Stu82] Michael Stuart. A geometric approach to principal components analysis. The American Statistician, 36(4):365–367, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683087.

SMSC:1986:SST

[Sug86] Sugar Mill Software Corp. Stat 1 — a statistical toolbox. The American Statistician, 40(1):57, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683131.

Saville:1986:MTS

[SW86] D. J. Saville and G. R. Wood. A method for teaching statistics using N-dimensional geometry. The American Statistician, 40 (3):205–214, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684537.

Swanson:1982:QRA

[Swa82] Steven M. Swanson. Quantitative risk assessment in light of the benzene decision. The American Statistician, 36(3):262–263, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683838.

Swanson:1984:STS

[Swa84] David A. Swanson. SUREX: Testing survival data for exponentiality on the VIC-20 microcomputer. The American Statistician, 38 (1):69–70, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683568.

Swindel:1981:GRR

[Swi81] Benee F. Swindel. Geometry of ridge regression illustrated. The American Statistician, 35(1):12–15, February 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683577.

Swift:1982:QSE

[Swi82] Jim Swift. [Quality of statistical education: Should ASA assist or assess?]: Comment. The American Statistician, 36(2):99, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684015.

Saw:1988:CCI

[SYC88] John G. Saw, Mark C. K. Yang, and Tse Chin. Correction to: “Chebyshev inequality with estimated mean Mo and variance” [Amer. Statist. 38 (1984), no. 2, 130–132, MR 85b:62019]. The American Statistician, 42(2):166, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [SYM84].

Saw:1984:CIE

[SYM84] John G. Saw, Mark C. K. Yang, and Tse Chin Mo. Chebyshev inequality with estimated mean and variance. The American Statistician, 38(2):130–132, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2683249. See correction [SYC88].

Seth:1987:PEP

[SZM87] Anand K. Seth, Benny Zee, and Sati Mazumdar. POLYGM: Estimation of parameters of a second-degree polynomial model under intraclass correlation structure and arbitrarily missing observations. The American Statistician, 41(4):331–332, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684763.

Taeuber:1981:IFS

[Tae81] Conrad Taeuber. [Improving the federal statistical system: Report of the President’s reorganization project for the federal statistical system]: Comment. The American Statistician, 35(4):199–202, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683288.

Tam:1984:COS

[Tam84] S. M. Tam. On covariances from overlapping samples. The American Statistician, 38(4):288–289, November 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683404.

Tanur:1982:QSEa

[Tan82a] Judith M. Tanur. Quality of statistical education: Should ASA assist or assess? The American Statistician, 36(2):90–93, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684009.

Tanur:1982:QSEb

[Tan82b] Judith M. Tanur. [Quality of statistical education: Should ASA assist or assess?]: Reply. The American Statistician, 36(2):101–102, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684017.

Tanis:1989:BRB

[Tan89] Elliot A. Tanis. Book review: [untitled]. The American Statistician, 43(2):121–122, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684517.

Tardiff:1981:LRC

[Tar81] Robert M. Tardiff. L’Hospital’s rule and the Central Limit Theorem. The American Statistician, 35(1):43, February 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683583.

Theil:1988:ITM

[TC88] Henri Theil and Ching-Fan Chung. Information-theoretic measures of fit for univariate and multivariate linear regressions. The American Statistician, 42(4):249–252, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685131.

Taeuber:1989:MWH

[TG89] Conrad Taeuber and Edwin D. Goldfield. In memoriam: William H. Shaw, 1909–1988. The American Statistician, 43(2):73, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684501.

Taylor:1981:SDS

[TH81] Warren H. Taylor, Jr. and H. Gill Hilton. A structure diagram symbolization for analysis of variance. The American Statistician, 35(2):85–93, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683146.

Thigpen:1987:SSP

[Thi87] Charles C. Thigpen. A sample-size problem in simple linear regression. The American Statistician, 41(3):214–215, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685107.

Thomas:1982:SCR

[Tho82] William A. Thomas. Supreme Court review of the OSHA Benzene Standard. The American Statistician, 36(3):264–266, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683839.

Thompson:1987:SSE

[Tho87] Steven K. Thompson. Sample size for estimating multinomial proportions. The American Statistician, 41(1):42–46, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684318.

Thurman:1989:UAR

[Thu89] Walter N. Thurman. Unconditional asymptotic results for simple linear regression. The American Statistician, 43(3):148–152, August 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685061.

Tiao:1983:USM

[Tia83] G. C. Tiao. Use of statistical methods in the analysis of environmental data. The American Statistician, 37(4):459–470, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683513.

Tomkins:1984:APB

[Tom84] R. James Tomkins. Another proof of Borel’s strong law of large numbers. The American Statistician, 38(3):208–209, August 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683656.

Torres:1986:NTI

[Tor86] Victor Aguirre Torres. A note on teaching infinite moving averages. The American Statistician, 40(1):40–41, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683115.

Trojanowski:1986:MSS

[Tro86] David Trojanowski. MICROSCALE: Statistical software system for testing human personality. The American Statistician, 40 (3):231–232, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684549.

Trumbo:1981:TCB

[Tru81] Bruce E. Trumbo. A theory for coloring bivariate statistical maps. The American Statistician, 35(4):220–226, November 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683294.

Turner:1980:FFP

[TT80] Danny W. Turner and F. Eugene Tidmore. FACES — a FOR-TRAN program for generating Chernoff-type faces on a line printer. The American Statistician, 34(3):187, August 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683888.

Tukey:1980:WNB

[Tuk80] John W. Tukey. We need both exploratory and confirmatory. The American Statistician, 34(1):23–25, February 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682991.

Tukey:1985:CMI

[Tuk85a] John W. Tukey. Comments on More intelligent statistical software and statistical expert systems: Future directions. The American Statistician, 39(??):12–14, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1985:DPP

[Tuk85b] John W. Tukey. Discussion: Projection pursuit. The American Statistician, 39(2):517–518, June 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349533;   http://www.
jstor.org/stable/2241189.

Tukey:1985:MIS

[Tuk85c] John W. Tukey. [More intelligent statistical software and statistical expert systems: Future directions]: Comment. The American Statistician, 39(1):12–14, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683898.

Tukey:1986:SS

[Tuk86] John W. Tukey. Sunset salvo. The American Statistician, 40 (1):72–76, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683137.

Turner:1987:BRB

[Tur87a] David L. Turner. Book review: [untitled]. The American Statistician, 41(1):68–70, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684326.

Turner:1987:MCS

[Tur87b] David L. Turner. Matrix calculator and stepwise interactive regression programs. The American Statistician, 41(4):329–330, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684760.

Tornqvist:1985:HSR

[TVV85] Leo Tornqvist, Pentti Vartia, and Yrjo O. Vartia. How should relative changes be measured? The American Statistician, 39 (1):43–46, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683905.

Thissen:1986:XMG

[TW86] David Thissen and Howard Wainer. XTREE: a multivariate graphical icon applicable in the evaluation of statistical estimators. The American Statistician, 40(2):149–153, May 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684878.

VanEck:1980:SAD

[Van80] Neal A. Van Eck. Statistical analysis and data management highlights of OSIRIS IV. The American Statistician, 34(2):119–121, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684124.

Vardeman:1986:LRI

[Var86] Stephen B. Vardeman. The legitimate role of inspection in modern SQC. The American Statistician, 40(4):325–328, November 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684620.

Vardeman:1984:SQP

[VD84] Stephen Vardeman and Herbert T. David. Statistics for quality and productivity: a new graduate-level statistics course. The American Statistician, 38(4):235–243, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683392.

vanderLaan:1988:LE

[vdLFS^{+}88] Paul van der Laan, Richard F. Fawcett, Kathleen C. Salter, J. Wanzer Drane, and Maria E. Gonzalez. Letters to the Editor. The American Statistician, 42(2):165–166, May 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684495.

Velleman:1985:MIS

[Vel85] Paul F. Velleman. [More intelligent statistical software and statistical expert systems: Future directions]: Comment. The American Statistician, 39(1):10–11, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683897.

Vinod:1989:BRB

[Vin89] H. D. Vinod. Book review: [untitled]. The American Statistician, 43(4):266–269, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685376.

Venables:1989:CBM

[VM89] W. N. Venables and K. W. Morris. A connection between minimum variance unbiased estimation and maximum likelihood. The American Statistician, 43(2):118–121, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684516.

Verner:1985:ECS

[VPO^{+}85] Jared Verner, Robert Pastorok, Joel O’Connor, William Severinghaus, Norman Glass, and Benee Swindel. Ecological community structure analyses in the formulation, implementation, and enforcement of law and policy. The American Statistician, 39(4):393–402, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683019.

Vaughan:1983:MPS

[VR83] William J. Vaughan and Clifford S. Russell. Monitoring point sources of pollution: Answers and more questions from statistical quality control. The American Statistician, 37(4):476–487, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683515.

Velleman:1981:ECR

[VW81] Paul F. Velleman and Roy E. Welsch. Efficient computing of regression diagnostics. The American Statistician, 35(4):234–242, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683296.

Velleman:1982:MEC

[VW82] Paul F. Velleman and Roy E. Welsch. [More on evaluating computer programs]: Comment. The American Statistician, 36(2):141– 142, May 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684035.

Vaupel:1985:HRS

[VY85] James W. Vaupel and Anatoli I. Yashin. Heterogeneity’s ruses: some surprising effects of selection on population dynamics. The American Statistician, 39(3):176–185, August 1985. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683925.

Wacker:1985:KRS

[Wac85] John G. Wacker. [The key role of statisticians in the transformation of North American industry]: Comment. The American Statistician, 39(3):228–229, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2683944. See [Joi85a, Hah85c, God85, Boa85].

Wagner:1982:SPR

[Wag82] Clifford H. Wagner. Simpson’s paradox in real life. The American Statistician, 36(1):46–48, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684093.

Wainio:1980:PSI

[Wai80] Marguerite Wainio. P-STAT 78 — an interactive computing system for file maintenance, cross-tabulation, and statistical analysis. The American Statistician, 34(3):179–180, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683878.

Wainer:1983:PPS

[Wai83] Howard Wainer. Pyramid power: Searching for an error in test scoring with 830,000 helpers. The American Statistician, 37(1): 87–91, February 1983. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2685852.

Wainer:1984:HDD

[Wai84] Howard Wainer. How to display data badly. The American Statistician, 38(2):137–147, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683253.

Wallman:1982:ETE

[Wal82] Katherine K. Wallman. [Experimental tests in education: Recommendations from the holtzman report]: Comment. The American Statistician, 36(1):9–10, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684083.

Wallman:1983:FSC

[Wal83] Katherine K. Wallman. Federal statistical coordination today: An epilogue as prologue. The American Statistician, 37(3):177–178, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683365.

Walsh:1987:PUF

[Wal87] David Walsh. PCPROBIT: a user-friendly probit analysis program for microcomputers. The American Statistician, 41(1):78, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684329.

Ware:1985:LMA

[War85] James H. Ware. Linear models for the analysis of longitudinal studies. The American Statistician, 39(2):95–101, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682803.

Watson:1980:PSC

[Wat80] Geoffrey S. Watson. [Preparing statisticians for careers in industry: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Industry]: Comment. The American Statistician, 34(2):79–80, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684110.

Watts:1981:TAA

[Wat81] Donald G. Watts. A task-analysis approach to designing a regression analysis course. The American Statistician, 35(2):77–84, May 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683145.

Watson:1982:SOO

[Wat82] Geoffrey S. Watson. Stratospheric ozone-observations and data analysis. The American Statistician, 36(3):312–316, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683850.

Watts:1983:RVS

[Wat83] Donald G. Watts. Review of videotape short course Probability. The American Statistician, 37(3):239–240, August 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683384.

Woolson:1981:ECC

[WB81] Robert F. Woolson and Stephen S. Brier. Equivalence of certain chi-squared test statistics. The American Statistician, 35 (4):250–253, November 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683301.

Wallman:1988:EOS

[WBH^{+}88] Katherine K. Wallman, Thomas Boardman, William Hill, Jerome Sacks, and Robert Tortora. Establishing an office of scientific and public affairs. The American Statistician, 42(4):245–248, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685130.

Waters:1987:OAM

[WC87] James R. Waters and Alexander J. Chester. Optimal allocation in multivariate, two-stage sampling designs. The American Statistician, 41(1):46–50, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684319.

Williams:1984:RVC

[WDS84] Robert Williams, Damon Disch, and Steven Sidik. Review of videotape course: Engineering statistics. The American Statistician, 38(1):66–67, February 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683565.

Weber:1987:CAD

[Web87] Ernst Weber. Computer analysis of data from limiting dilution experiments of cells of the immune system. The American Statistician, 41(3):240, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685121.

Weinstein:1982:IRA

[Wei82a] Milton C. Weinstein. Issues in risk assessment and analysis for the chlorofluorocarbon problem. The American Statistician, 36 (3):317–320, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683851.

Weiss:1982:MJW

[Wei82b] Lionel Weiss. In memoriam: Jacob Wolfowitz, (1910-1981). The American Statistician, 36(2):126–127, May 1982. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684024.

Weissman:1983:FGD

[Wei83] Ishay Weissman. On francis Galton’s difference problem. The American Statistician, 37(3):216–217, August 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683373.

Weisberg:1988:BRB

[Wei88] Sanford Weisberg. Book review: [untitled]. The American Statistician, 42(3):224–225, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685012.

Welch:1985:AAC

[Wel85] William J. Welch. ACED: Algorithms for the construction of experimental designs. The American Statistician, 39(2):146, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682827.

Wainer:1980:EIC

[WF80] Howard Wainer and Carl M. Francolini. An empirical inquiry concerning human understanding of two-variable color maps. The American Statistician, 34(2):81–93, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684111.

Wheeler:1985:RTKa

[Whe85] Bob Wheeler. Regression tool kit: Software for the algorithmic design of experiments. The American Statistician, 39(2):144, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682823;
http://www.jstor.org/stable/2682824.

Whittemore:1985:ACM

[Whi85] Alice S. Whittemore. Analyzing cohort mortality data. The American Statistician, 39(4):437–441, November 1985. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683027.

Whitmore:1986:PLU

[Whi86] G. A. Whitmore. Prediction limits for a univariate normal observation. The American Statistician, 40(2):141–143, May 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684875.

Whittemore:1989:EVR

[Whi89] Alice S. Whittemore. Errors-in-variables regression using Stein estimates. The American Statistician, 43(4):226–228, November 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2685366.

Wilkinson:1980:RMG

[Wil80] Leland Wilkinson. REGM: a multivariate general linear hypothesis program. The American Statistician, 34(3):182, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683881.

Williams:1981:BSO

[Wil81] James G. Williams. BKWPAC: Single observation regression diagnostics with plots and reweighted least squares. The American Statistician, 35(3):163, August 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683992.

Williams:1982:SDR

[Wil82] Byron K. Williams. A simple demonstration of the relationship between classification and canonical variates analysis. The American Statistician, 36(4):363–365, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683086.

Willemain:1986:S

[Wil86] Thomas R. Willemain. SmartForecasts. The American Statistician, 40(1):57, February 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683130.

Wilkinson:1987:MPV

[Wil87] Leland Wilkinson. MYSTAT: a personal version of SYSTAT. The American Statistician, 41(4):334, November 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684766.

Wyzga:1985:FMD

[WL85] Ronald E. Wyzga and Fred W. Lipfert. Forecasting materials damage from air pollution. The American Statistician, 39(4):423–430, November 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683024.

Wolf:1985:BRB

[WM85] Fredric M. Wolf and Bruce R. Maxim. Book review: [untitled]. The American Statistician, 39(2):127, May 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682814.

White:1987:SGC

[WM87] Kenneth J. White and Robert N. McRae. SHAZAM: a general computer program for econometric methods (version 5). The American Statistician, 41(1):80, February 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684334.

Warren:1982:BIS

[WMB82] J. A. Warren, R. D. Michener, and A. Bugbee. A buffered introductory statistics course. The American Statistician, 36(1):31–36, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684089.

Womer:1985:ILS

[Wom85] Norman Keith Womer. Identification and least squares in simultaneous equations. The American Statistician, 39(4):295–297, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wood:1980:LNL

[Woo80] Fred S. Wood. LINWOOD and NONLINWOOD-linear and nonlinear least squares curve-fitting programs. The American Statistician, 34(3):177–179, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683877.

Wood:1984:CEC

[Woo84] Fred S. Wood. Comment: Effect of centering on collinearity and interpretation of the constant. The American Statistician, 38(2):88– 90, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683240.

Woodworth:1987:BRB

[Woo87] George Woodworth. Book review: [untitled]. The American Statistician, 41(3):231–233, August 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2685114.

Wolkowicz:1980:HSI

[WS80] H. Wolkowicz and G. P. H. Styan. A history of Samuelson’s inequality. The American Statistician, 34(??):250, ???? 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Whitmore:1987:HDI

[WS87] G. A. Whitmore and V. Seshadri. A heuristic derivation of the inverse Gaussian distribution. The American Statistician, 41(4): 280–281, November 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684749.

Willett:1988:ACN

[WS88] John B. Willett and Judith D. Singer. Another cautionary note about R^{2}: Its use in weighted least-squares regression analysis. The American Statistician, 42(3):236–238, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685031.

Worsley:1985:LE

[WSA^{+}85] Keith J. Worsley, Steven J. Schwager, J. Aleong, Ed Carlstein, Daniel Krewski, Mikelis Bickis, A. Arcese, H. Joseph Newton, A. Ian McLeod, Carlos Jimenez, Steven D. Stellman, William S.

Cleveland, A. I. Dale, Adi Raveh, Norman L. Johnson, Walter W. Piegorsch, George Casella, R. W. Farebrother, S. R. Searle, Spyros Missiakoulis, Robin Darton, Oscar Kempthorne, and Craig F. Ansley. Letters to the Editor. The American Statistician, 39 (3):235–241, August 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683949.

Wolde-Tsadik:1980:SSI

[WT80] Girma Wolde-Tsadik. Sample sizes in the interval estimation of the correlation coefficient. The American Statistician, 34(4):227–228, November 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684065.

YakiSoft:1988:Y

[Yak88] YakiSoft. YREG. The American Statistician, 42(4):273, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2685139.

Yang:1988:CLT

[Yan88] Shie-Shien Yang. A Central Limit Theorem for the bootstrap mean. The American Statistician, 42(3):202–203, August 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685003.

Yellin:1983:SDA

[Yel83a] Joel Yellin. [Science and democracy in the age of technology: Separating fact from value]: Rejoinder. The American Statistician, 37(4):372–373, November 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2683497.

Yellin:1983:WSM

[Yel83b] Joel Yellin. Who shall make environmental decisions? The American Statistician, 37(4):362–366, November 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683495.

Young:1980:AMS

[YTL80] Forrest W. Young, Yoshio Takane, and Rostylaw Lewyckyj. ALSCAL: a multidimensional scaling package with several individual differences options. The American Statistician, 34(2):117–118, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684122.

Zahn:1982:PSC

[Zah82] Douglas A. Zahn. [Preparing statisticians for careers in the federal government: Report of the ASA Section on Statistical Education Committee on Training of Statisticians for Government]: Comment. The American Statistician, 36(2):88–89, May 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684007.

Zarnoch:1980:NNS

[Zar80] Stanley J. Zarnoch. NONPAR: a nonparametric statistics computer package. The American Statistician, 34(2):115–116, May 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2684120.

Zellner:1982:JN

[Zel82a] Arnold Zellner. Is Jeffreys a ‘necessarist’? The American Statistician, 36(1):28–30, February 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2684088. See comment [Kas82] and reply [Zel82b].

Zellner:1982:RCJ

[Zel82b] Arnold Zellner. Reply to a comment on “Is Jeffreys a ‘Necessarist’?”. The American Statistician, 36(4):392–393, November 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683100. See

[Zel82a, Kas82].

Zellner:1984:PDB

[Zel84] Arnold Zellner. [On prior distributions for binary trials]: Comment. The American Statistician, 38(4):249–250, November 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683396.

Zellner:1985:CPD

[Zel85] Arnold Zellner. Correction: “On prior distributions for binary trials” [Amer. Statist. 38 (1984), no. 4, 244–251, MR 87c:62004a] by S. Geisser. The American Statistician, 39(2):158, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Gei84b].

Zellner:1988:OIPb

[Zel88a] Arnold Zellner. [Optimal information processing and Bayes’s Theorem]: Reply. The American Statistician, 42(4):283–284, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2685148.

Zellner:1988:OIPa

[Zel88b] Arnold Zellner. Optimal information processing and Bayes’s Theorem. With comments and a reply by the author. The American Statistician, 42(4):278–280, November 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685143.

Zellner:1989:PPB

[Zel89] Arnold Zellner. PCBRAP: a PC Bayesian regression package. The American Statistician, 43(2):124, May 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684519.

Zervos:1982:FBM

[ZR82] Constantine Zervos and Joseph V. Rodricks. FDA’s ban of DES in meat production. The American Statistician, 36(3):278–283, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683843.